### v1

Abstract.
We extend the classical theorems of Khintchine and Schmidt in metric Diophantine approximation to the context of self-similar measures on $\mathbb{R}^{d}$. For this, we establish effective equidistribution of associated random walks on $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$.
Our result strengthens that of [7] which requires $d=1$ and restricts Schmidt-type counting estimates to approximation functions which decay fast enough.
Novel techniques include a bootstrap scheme for the associated random walks despite algebraic obstructions, and a refined treatment of Dani’s correspondence. We also establish non-concentration properties of self-similar measures near algebraic subvarieties of $\mathbb{R}^{d}$.
2010 Mathematics Subject Classification:
Primary 37A99, 11J83; Secondary 22F30.
W.H. is supported by the National Key R&D Program of China (No. 2022YFA1007500) and the National Natural Science Foundation of China (No. 12288201).
H.Z. is supported by the startup grant of Soochow University.
1. Introduction
Let $d\geq 1$ be an integer. Denote by $\operatorname{Sim}(\mathbb{R}^{d})$ the group of similarities of $\mathbb{R}^{d}$, i.e. transformations $\phi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ of the form $\boldsymbol{s}\mapsto\mathtt{r}_{\phi}O_{\phi}\boldsymbol{s}+\mathtt{b}_{\phi}$ where $\mathtt{r}_{\phi}>0$ and $O_{\phi}\in O(d)$ is a linear orthogonal transformation of $\mathbb{R}^{d}$, and $\mathtt{b}_{\phi}\in\mathbb{R}^{d}$. A probability measure $\lambda$ on $\operatorname{Sim}(\mathbb{R}^{d})$ is called a randomized self-similar iterated function system (IFS). We say that $\lambda$ is strongly irreducible if $\mathbb{R}^{d}$ is the only finite union of affine subspaces of $\mathbb{R}^{d}$ which is invariant under $\lambda$-almost every $\phi$. We also say that $\lambda$ has a finite exponential moment if there exists $\varepsilon>0$ such that
$\int_{\operatorname{Sim}(\mathbb{R}^{d})}\mathtt{r}_{\phi}^{\varepsilon}+\mathtt{r}_{\phi}^{-\varepsilon}+\|\mathtt{b}_{\phi}\|^{\varepsilon}\,\mathrm{d}\lambda(\phi)\,<+\infty.$
Throughout this paper, we consider a probability measure $\sigma$ on $\mathbb{R}^{d}$ which is self-similar in the sense that it is stationary under some randomized self-similar IFS $\lambda$ which is strongly irreducible and has a finite exponential moment. Recall that stationarity means
$\sigma=\int_{\operatorname{Sim}(\mathbb{R}^{d})}\phi_{\star}\sigma\,\mathrm{d}\lambda(\phi).$
Although we do not impose the maps in the support of $\lambda$ to be contractive, it turns out that the existence of $\sigma$ implies that $\lambda$ is contractive in average, i.e.
$\int_{\operatorname{Sim}(\mathbb{R}^{d})}\log\mathtt{r}_{\phi}\,\mathrm{d}\lambda(\phi)<0,$
and vice versa, see [12, Theorem 2.5].
Example.
On $\mathbb{R}$, classical examples of self-similar measures are the Lebesgue measure, the normalized Hausdorff measure on a missing digit Cantor set, or Bernouilli convolutions. In higher dimension, one may consider powers of missing digit Cantor measures, the normalized Hausdorff measure on a Sierpiński triangle, Sierpiński carpet, etc. In general, every randomized self-similar IFS with finite exponential moment and which is contractive in average, has a unique stationary probability measure [38, Theorem 3.1].
The goal of the paper is to study the Diophantine properties of typical points chosen by a self-similar measure on $\mathbb{R}^{d}$. This topic originates from a question of Mahler [43], asking how well points in the middle-thirds Cantor set can be approximated by rationals. Mahler’s question is later recast by Kleinbock-Lindenstrauss-Weiss [33] who ask whether self-similar measures may satisfy a dichotomy in the spirit of Khintchine’s theorem. We first recall this theorem and present our main result, then we discuss how it connects to earlier works.
Given a non-increasing function $\psi:\mathbb{N}\rightarrow\mathbb{R}_{\geq 0}$, we write $W(\psi)$ the set of $\psi$-approximable vectors in $\mathbb{R}^{d}$. In other terms, a vector $\boldsymbol{s}\in\mathbb{R}^{d}$ belongs to $W(\psi)$ if for infinitely many $(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\mathbb{N}$, one has
(1.1) $\|q\boldsymbol{s}-\boldsymbol{p}\|_{\infty}<\psi(q),$
where $\lVert\cdot\rVert_{\infty}$ denotes the supremum norm on $\mathbb{R}^{d}$. The celebrated Khintchine theorem [30, 31] states that $W(\psi)$ has either null or full Lebesgue measure, and that each scenario can be read simply on the function $\psi$, as they respectively correspond to $\sum_{q\in\mathbb{N}}\psi(q)^{d}<\infty$ and $\sum_{q\in\mathbb{N}}\psi(q)^{d}=\infty$. In the divergent case, Khintchine’s theorem has been further refined by Schmidt [49] who provided an asymptotic estimate for the number of solutions to the Diophantine inequality (1.1) with bounded $q$.
In this paper, we extend the theorems of Khintchine and Schmidt to the context of self-similar measures. Our main theorem is the following.
Theorem 1.1 (Khintchine and Schmidt for self-similar measures).
Let $\lambda$ be a probability measure on $\operatorname{Sim}(\mathbb{R}^{d})$, and assume $\lambda$ is strongly irreducible with finite exponential moment. Let $\sigma$ be a $\lambda$-stationary probability measure on $\mathbb{R}^{d}$. Let $\psi:\mathbb{N}\rightarrow\mathbb{R}_{\geq 0}$ be a non-increasing function.
Then we have the dichotomy
(1.2) $\sigma(W(\psi))=\begin{cases}0&\text{ if }\sum_{q\in\mathbb{N}}\psi(q)^{d}<\infty,\\ 1&\text{ if }\sum_{q\in\mathbb{N}}\psi(q)^{d}=\infty.\end{cases}$
Moreover, in the divergent case $\sum_{q\in\mathbb{N}}\psi(q)^{d}=\infty$, we have the following asymptotic: for $\sigma$-a.e. $\boldsymbol{s}\in\mathbb{R}^{d}$, as $n\to+\infty$:
$\big{|}\{(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\llbracket 0,n\rrbracket\,:\,\|q\boldsymbol{s}-\boldsymbol{p}\|_{\infty}<\psi(q)\}\big{|}\sim_{\boldsymbol{s},\psi}2^{d}\sum_{q=0}^{n}\psi(q)^{d}.$
Here we use the notation $\llbracket 0,n\rrbracket=[0,n]\cap\mathbb{N}$. An asymptotic estimate of the number of primitive solutions of the Diophantine inequality is also provided, see (8.5).
Let us now explain how the topic has evolved from Mahler’s question in the 80’s to the above ˜1.1. In the early literature, the convergence and divergence aspects of ˜1.1 were addressed separately for specific approximation functions. The first significant result in this direction was obtained by Weiss [56] for $\psi(q)=1/q^{1+\varepsilon}$ and measures on the real line satisfying a certain decay condition, including the middle thirds Cantor measure. This work was later generalized by Kleinbock-Lindenstrauss-Weiss [33] to a broader class of measures on $\mathbb{R}^{d}$, known as friendly measures. Subsequent important developments, adopting similar terminology, include [45, 17, 18]. The study of the case $\psi(q)=\varepsilon/q$ was conducted by Einsiedler-Fishman-Shapira [21] for missing digit Cantor measures, and generalized significantly by Simmons-Weiss [52] to arbitary self-similar measures.
Without restriction on the non-increasing function $\psi$, Khalil-Luethi [28] obtained the Khintchine dichotomy (1.2) under the condition that the self-similar measure $\sigma$ has a large enough dimension, and $\lambda$ is rational, finitely supported, contractive, and satisfies the open set condition. For instance, the Cantor measure with single-missing digit in base $5$ is in this class, but not the middle thirds Cantor measure, see also the subsequent related works [58, 19] which all assume large dimension for $\sigma$ in some sense. In our previous paper [7], we proposed a new approach, based on projection theorems à la Bourgain, which led to the Khintchine dichotomy (1.2) in the case where $d=1$, and with no further restriction, i.e. dealing with any self-similar measure $\sigma$ and any non-increasing function $\psi$. The assumption $d=1$ is however needed at the most crucial dimension bootstrapping step in the proof. We also obtained Schmidt’s counting theorem for primitive solutions under the additional assumption $\psi(q)<1/q$. Pushing one step further, ˜1.1 establishes the Khintchine dichotomy for self-similar measures in arbitrary dimension, along with a full Schmidt counting theorem, i.e without any restriction on $\psi$ nor asking solutions of (1.1) to be primitive.
Other related topics. Mahler [43] also suggested the study of intrinsic Diophantine approximations on fractal sets, i.e. approximation by rationals sitting in the fractal itself. For works in this direction, see e.g. [54, 14]. Beside fractal sets, Diophantine approximation on embedded submanifolds has also attracted much attention over the past years, see e.g. [36, 55, 9, 10].
˜1.1 will be deduced from a dynamical statement which we now present. Let $G=\operatorname{SL}_{d+1}(\mathbb{R})$, let $\Lambda\subseteq G$ be a lattice. Consider the quotient space $X=G/\Lambda$ equipped with the standard Riemannian metric and Haar probability measure $m_{X}$. For $x\in X$, we write $\operatorname{inj}(x)$ the injectivity radius of $x$. For $l\in\mathbb{N}$, we denote by $B^{\infty}_{\infty,l}(X)$ the collection of smooth functions on $X$ whose derivatives up to order $l$ are bounded, and set ${\mathcal{S}}_{\infty,l}(\cdot)$ the associated norm (see Section˜2 for more details on these conventions). For $t>0$ and $\boldsymbol{s}=(s_{1},\cdots,s_{d})\in\mathbb{R}^{d}$, consider $a(t),u(\boldsymbol{s})\in G$ given by
$a(t)=\begin{pmatrix}t^{\frac{1}{d+1}}&&&\\ &\ddots&&\\ &&t^{\frac{1}{d+1}}&\\ &&&t^{-\frac{d}{d+1}}\end{pmatrix},\,\,\,\,\,\quad\quad u(\boldsymbol{s})=\begin{pmatrix}1&&&s_{1}\\ &\ddots&&\vdots\\ &&1&s_{d}\\ &&&1\end{pmatrix}$.
We show
Theorem 1.2 (Effective equidistribution of expanding fractals).
Let $\sigma$ be as in ˜$1.$ For every $x\in X$, $t>1$, $f\in B^{\infty}_{\infty,l}(X)$, we have
$\left\lvert\int_{\mathbb{R}^{d}}f\bigl{(}a(t)u(\boldsymbol{s})x\bigr{)}\,\mathrm{d}\sigma(\boldsymbol{s})-m_{X}(f)\right\rvert\leq C{\mathcal{S}}_{\infty,l}(f)\operatorname{inj}(x)^{-1}t^{-c}.$
where the constants $C,c>0$ only depend on $\Lambda,\sigma$, and $l=\lceil\frac{1}{2}\dim\operatorname{SO}(d+1)\rceil$.
˜1.2 establishes the exponential equidistribution of the measure $\sigma$, viewed along a unipotent orbit based at an arbitrary point $x$ and expanded under the associated diagonal flow. The term $\operatorname{inj}(x)^{-1}$ in the error term reflects that equidistribution takes longer when the basepoint $x$ is high in a cusp.
The qualitative convergence (i.e. without rate) implied by ˜1.2 resonates with Khalil-Luethi-Weiss [29], which establishes equidistribution under expansion by a broader class of diagonal flows, but at the cost of restricting the IFS to be “carpet”, i.e. rational, with equal contraction ratios, and no rotation part.
From a quantitative point of view, ˜1.2 can be seen as an effective version of Ratner’s theorem for the multiparameter unipotent flow $(u(\boldsymbol{s}))_{\boldsymbol{s}\in\mathbb{R}^{d}}$ in the context of fractal measures. The non-fractal case (i.e. $\sigma$ is absolutely continuous with respect to Lebesgue) is due to Kleinbock-Margulis, in the broader context of expanding translates of horospherical subgroups [35, 37]. Related works in this direction for the non-horospherical case include [53, 15, 32, 41, 57, 42]. In the context of fractal measures, Khalil-Luethi [28] obtained ˜1.2 under the additional assumption the point $x$ lies in a certain countable set determined by the IFS. They also required the IFS to be rational, contractive, finitely supported, to satisfy the open set condition and has large enough dimension. See also [19] for a different approach for $d=1$. Provided $d=1$, those constraints were eliminated in our previous work [7]. We now generalize [7] to arbitrary dimensions, achieving the theorem in full generality.
The connection between ˜1.2 and the Khintchine dichotomy in ˜1.1 comes from Dani’s philosophy [16]. Roughly speaking, it claims that the Diophantine properties of a vector $\boldsymbol{s}\in\mathbb{R}^{d}$ can be read in the dynamics of the trajectory $(a(t)u(\boldsymbol{s})x_{0})_{t>1}$ on $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$, where $x_{0}$ is the identity coset. The correspondence is rooted in the simple computation
$a(t)u(\boldsymbol{s})(-\boldsymbol{p},q)=(t^{\frac{1}{d+1}}(q\boldsymbol{s}-\boldsymbol{p}),\,t^{-\frac{d}{d+1}}q),$
which yields that for every $I\subseteq\mathbb{R},J\subseteq\mathbb{R}^{d}$, the statement
$\exists(\boldsymbol{p},q)\in\mathbb{Z}^{d+1}\,:\,q\in I\text{ and }q\boldsymbol{s}-\boldsymbol{p}\in J$
is equivalent to the lattice $a(t)u(\boldsymbol{s})\mathbb{Z}^{d+1}\subseteq\mathbb{R}^{d+1}$ intersecting the product set $t^{\frac{1}{d+1}}J\times t^{-\frac{d}{d+1}}I$. Now, identifying $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$ to the space of covolume $1$ lattices in $\mathbb{R}^{d+1}$ and provided the product set $t^{\frac{1}{d+1}}J\times t^{-\frac{d}{d+1}}I$ looks like a ball, the latter condition can be interpreted dynamically, at which point ˜1.2 can be used. Since Dani’s insight [16], many works have exploited this connection, e.g. [36, 34, 15, 28, 7]. In Section˜8, we will show how effective decorrelation of expanding translates (consequence of ˜1.2) can be utilized to obtain the divergent case of the Khintchine dichotomy, along with a rate as in Schmidt’s counting theorem. This extends [7] which assumed $d=1$ and restricted counting to primitive solutions, both assumptions being required to deal with bounded Siegel transforms. In this paper, we will tackle Siegel transforms which are not even $L^{2}$.
To prove ˜1.2, we exploit the self-similarity of $\sigma$ to see that the translate $a(t)u(\boldsymbol{s})x\,\mathrm{d}\sigma(s)$ is roughly given by the $\log t$-step of a random walk on $X$ associated to $\lambda$ (see Lemma˜7.2). This change of view point originates in the work of Simmons-Weiss [52], see also [46, 47, 28, 20, 1] for further developments in this direction. The random walk is defined as follows. Assume $\lambda$-a.e. $\phi$ is orientation preserving (i.e. $\det O_{\phi}=1$), write $k_{\phi}=\begin{pmatrix}O_{\phi}&\\ &1\end{pmatrix}$, and let $\mu$ be the probability measure on $\operatorname{SL}_{d+1}(\mathbb{R})$ given by
(1.3) $\mu=\int_{\operatorname{Sim}(\mathbb{R}^{d})}\delta_{k^{-1}_{\phi}a(\mathtt{r}^{-1}_{\phi})u(\mathtt{b}_{\phi})}\,\mathrm{d}\lambda(\phi).$
We show
Theorem 1.3 (Effective equidistribution of the $\mu$-random walk).
Let $\lambda$ be as in ˜1.1 and orientation preserving. Let $\mu$ be the associated probability measure on $\operatorname{SL}_{d+1}(\mathbb{R})$ as in (1.3). Then for every $x\in X$, $n\geq 1$ and $f\in B^{\infty}_{\infty,l}(X)$, we have
$\left\lvert\mu^{*n}*\delta_{x}(f)-m_{X}(f)\right\rvert\leq C{\mathcal{S}}_{\infty,l}(f)\operatorname{inj}(x)^{-1}e^{-cn}$
where the constants $C,c>0$ only depend on $\Lambda,\lambda$, and $l=\lceil\frac{1}{2}\dim\operatorname{SO}(d+1)\rceil$.
Let us explain the steps of the proof of ˜1.3. The overall strategy is similar to that in [7], which itself is inspired by [6]. It is done in three phases, each analyzing the dimension of the $\mu$-random walk. In contrast to the one-dimensional scenario in [7], the present higher dimensional setting imposes new difficulties in the second phase, as we explain below.
The first phase is to show that the random walk acquires some (small) positive dimension (Proposition˜5.1): there exists $A,\kappa>0$ depending on $\Lambda,\lambda$ such that for any $\rho>0$ small, $x,y\in X$, and any $n\geq\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}x\rvert$, we have
$\mu^{*n}*\delta_{x}(B_{\rho}y)\leq\rho^{\kappa}.$
The proof of this statement closely follows the argument in [7, Proposition 3.1], and is based on effective recurrence of the random walk (Proposition˜4.1).
In the second phase, we bootstrap the initial dimension $\kappa$ arbitrarily close to the dimension of the ambient space by convolving with $\mu$ suitably many times (Proposition˜6.2). However, unlike the one-dimensional case in [7], the multislicing method proposed in [6, Section 2] cannot be applied directly to implement this bootstrapping argument in the current higher dimensional setting. This limitation arises because the non-concentration hypothesis described in [6, Theorem 2.1] is never satisfied for $d\geq 2$ due to algebraic obstructions (see Lemma˜6.3). To resolve this limitation, we promote a mild non-concentration hypothesis (MNC) which also enables the dimensional bootstrapping (Definition˜6.9, Proposition˜6.14)  [^1] Such a strategy resonates with the concurrent and independent work of Zuo Lin [40], which manages algebraic obstructions to dimensional bootstrapping in the context of homogeneous spaces of $\operatorname{SL}_{4}(\mathbb{R})$ acted upon by certain two-parameter unipotent flows. . To validate (MNC) in our setting, we first need to rule out potential algebraic obstructions, which is done in Propositions 6.6, 6.7. We then upgrade the absence of obstruction to the non-concentration property (MNC). That part of the argument requires regularity properties of the self-similar measure $\sigma$, namely that $\sigma$ is Hölder regular with respect to proper algebraic subvarieties (˜3.3). We establish this property in Section˜3; it is of independent interest and generalizes some of the results in [33, Section 7].
In the concluding phase, as the dimension approaches that of the ambient space, we complete the proof by applying the spectral gap property of the convolution operator $f\mapsto\mu*f$ acting on $L^{2}(X)$. Finally, ˜1.2 follows from ˜1.3, via the connection between random walks and expanding translates of self-similar measures given in Lemma˜7.2.
To derive the Khintchine dichotomy and Schmidt’s counting theorem, we remark that the convergence part follows from the single effective equidistribution theorem in ˜1.2, as explained in [28, Theorem 9.1]. In order to handle the asymptotic counting in the divergence part, we need to truncate the associated Siegel’s transform appropriately, and then apply different counting strategies according to the values of $\psi$. More precisely, for the case where $\psi(q)$ is not too large, we adopt a refined version of the counting method in [7], which is based on effective double equidistribution and is inspired by the original papers of Schmidt [49, 50]. To handle the part where $\psi(q)$ is large, inspired by Huang-Saxcé [27] and Pfitscher [44], we make use of the fact that for $\sigma$-a.e. $\boldsymbol{s}$, the lattice $a(t)u(\boldsymbol{s})\mathbb{Z}^{d+1}$ is not too "distorted" as $t\to+\infty$, which guarantees the number of lattice points in $a(t)u(\boldsymbol{s})\mathbb{Z}^{d+1}$ intersecting a large ball is asymptotic to the volume of the ball. This analysis is conducted in Section˜8. The main challenge for this section, which is new compared to [7], is that we have to deal with Siegel transforms which are not bounded, and not even $L^{2}$ when $d=1$.
Structure of the paper. In Section˜2, we set up notations for the rest of the paper, and recall some basic facts on self-similar measures. In Section˜3, we prove the Hölder regularity of self-similar measures with respect to proper algebraic subvarieties of $\mathbb{R}^{d}$. In Section˜4, we recall the effective recurrence property of the $\mu$-random walk on $X$. In Section˜5, we show that the distribution $\mu^{*n}*\delta_{x}$ acquires small positive dimension at exponentially small scales as long as $n$ is large enough depending on $\operatorname{inj}(x)$. In Section˜6, we promote a mild non-concentration property and implement it in our context to bootstrap the dimension arbitrarily close to $\dim X$. In Section˜7, we deduce Theorems 1.2, 1.3 using the spectral gap of associated Markov operator, we also derive a double equidistribution property. Finally in Section˜8, we show ˜1.1.
Acknowledgements. The authors thank René Pfitscher and Shucheng Yu for insightful discussions concerning Schmidt’s counting theorem and moment estimates of Siegel transforms.
2. Notations
Throughout the paper, we fix the following notations.
We let $d\geq 1$ be an integer. We write $G=\operatorname{SL}_{d+1}(\mathbb{R})$, fix $\Lambda\subseteq G$ to be a lattice, and set $X=G/\Lambda$ the quotient space. We let $m_{G}$ denote the $G$-invariant Borel measure on $G$ normalised so that the induced finite measure $m_{X}$ on $X$ has tatal mass $1$. Both $m_{G}$ and $m_{X}$ are referred to as Haar measures.
Metric. Given integers $m,n\geq 1$, we write $\operatorname{M}_{m,n}(\mathbb{R})$ the space of real matrices with $m$ rows and $n$ columns. We equip $\operatorname{M}_{m,n}(\mathbb{R})$ with its standard Euclidean structure. More precisely, writing $E_{i,j}$ the matrix with coefficient $1$ at the position $(i,j)$ and null coefficients elsewhere, the collection $\{E_{i,j}\,:\,1\leq i\leq m,1\leq j\leq n\}$ is an orthonormal basis of $\operatorname{M}_{m\times n}(\mathbb{R})$. This Euclidean structure extends naturally to the exterior algebra $\bigwedge^{*}\operatorname{M}_{m,n}(\mathbb{R})$. We denote by $\|\cdot\|$ the associated Euclidean norm on $\operatorname{M}_{m,n}(\mathbb{R})$, and more generally on $\bigwedge^{*}\operatorname{M}_{m,n}(\mathbb{R})$.
Set $\mathfrak{g}:=\mathfrak{sl}_{d+1}(\mathbb{R})$ the Lie algebra of $G$. We equip $G$ with the unique right $G$-invariant Riemannian metric which coincides with $\|\cdot\|_{|\mathfrak{g}}$ at $\mathfrak{g}=T_{\operatorname{Id}}G$. We write $\operatorname{dist}(\cdot,\cdot)$ the induced distance on $G$, or the quotient distance on $X$.
Given $\rho>0$, we write $B_{\rho}$ the open ball of radius $\rho$ centered at the neutral element $\operatorname{Id}$ in $G$. In particular, given a point $x\in X$, the set $B_{\rho}x$ is the ball in $X$ of radius $\rho$ and center $x$.
We define the injectivity radius of $X$ at $x$ by
$\operatorname{inj}(x)=\sup\{\,\rho>0\,:\,\text{the map $B_{\rho}\rightarrow X,g\mapsto gx$ is injective}\,\}.$
In the Euclidean space $\mathbb{R}^{d}$, we set $B^{\mathbb{R}^{d}}_{\rho}$ the open ball of radius $\rho$ centered at the origin. On some rare occasions (Section˜3), we might use $B_{\rho}$ as a shorthand for $B^{\mathbb{R}^{d}}_{\rho}$. We will explicitely warn about this exception at the few places it occurs.
Sobolev norms. Write ${\mathcal{A}}=\{E_{i,j}\,:\,1\leq i,j\leq d+1,\,i\neq j\}\cup\{E_{i,i}-E_{i+1,i+1}\,:\,i=1,\dots d\}$ the standard basis of $\mathfrak{g}$. Given $l\in\mathbb{N}$, write $\Xi_{l}$ the set of words of length $l$ with letters in $\mathcal{A}$}. Each $D\in\Xi_{l}$ acts as a differential operator on the space of smooth functions $C^{\infty}(X)$. Given $f\in C^{\infty}(X)$, $p\in[1,\infty]$, we set
${\mathcal{S}}_{p,l}(f)=\sum_{D\in\Xi_{l}}\|Df\|_{L^{p}},$
where $\|\cdot\|_{L^{p}}$ refers to the $L^{p}$-norm for the Haar probability measure $m_{X}$ on $X$. We let $B^{\infty}_{p,l}(X)$ denote the space of smooth functions $f$ on $X$ such that ${\mathcal{S}}_{p,l}(f)<\infty$.
Driving measures $\lambda$ and $\mu$. Let $\operatorname{Sim}(\mathbb{R}^{d})^{+}$ be the set of orientation preserving similarities of $\mathbb{R}^{d}$. Every $\phi\in\operatorname{Sim}(\mathbb{R}^{d})^{+}$ can be written uniquely
$\phi(\boldsymbol{s})=r_{\phi}O_{\phi}\boldsymbol{s}+\mathtt{b}_{\phi},\quad\boldsymbol{s}\in\mathbb{R}^{d},$
for some $O_{\phi}\in\operatorname{SO}_{d}(\mathbb{R})$, $r_{\phi}>0$ and $\mathtt{b}_{\phi}\in\mathbb{R}^{d}$.
We set
$K^{\prime}=\begin{pmatrix}\operatorname{SO}_{d}(\mathbb{R})&\\ &1\end{pmatrix}$,
$A^{\prime}=\{\,a(t):=\operatorname{diag}(t^{\frac{1}{d+1}},\cdots,t^{\frac{1}{d+1}},t^{-\frac{d}{d+1}}):t\in\mathbb{R}_{>0}\,\}$
$U=\left\{\,u(\boldsymbol{s}):=\begin{pmatrix}1&&&s_{1}\\ &\ddots&&\vdots\\ &&1&s_{d}\\ &&&1\end{pmatrix}:\boldsymbol{s}\in\mathbb{R}^{d}\,\right\}.$
Here, we use the implicit convention the $s_{i}$’s are the coordinates of $\boldsymbol{s}$, more precisely $\boldsymbol{s}=(s_{1},\dots,s_{d})$. Throughout the paper, this convention applies.
Consider the subgroup $P^{\prime}=K^{\prime}A^{\prime}U\subseteq G$. Every $g\in P^{\prime}$ can be uniquely written as
$g=k_{g}^{-1}a(\mathtt{r}_{g}^{-1})u(\mathtt{b}_{g})$
where $k_{g}=\begin{pmatrix}O_{g}&\\ &1\end{pmatrix}\in K^{\prime}$, $\mathtt{r}_{g}>0$, and $\mathtt{b}_{g}\in\mathbb{R}^{d}$. There is an anti-isomorphism [^2] That is $\phi_{g_{1}g_{2}}=\phi_{g_{2}}\phi_{g_{1}}$ for all $g_{1},g_{2}\in P^{\prime}$. between $P^{\prime}$ and $\operatorname{Sim}(\mathbb{R}^{d})^{+}$ given by
$g\in P^{\prime}\mapsto\phi_{g}\in\operatorname{Sim}(\mathbb{R}^{d})^{+},$
where
$\phi_{g}(\boldsymbol{s})=\mathtt{r}_{g}O_{g}\boldsymbol{s}+\mathtt{b}_{g}.$
Throughout this paper, we fix a probability measure $\lambda$ on $\operatorname{Sim}(\mathbb{R}^{d})^{+}$ and denote by $\mu$ the corresponding probability measure on $P^{\prime}$ via the above anti-isomorphism. Note that $\lambda$ and $\mu$ determine each other.
We assume that $\lambda$, and equivalently $\mu$, has finite exponential moment, which means that there exists $\varepsilon>0$ such that
$\int_{P^{\prime}}|\mathtt{r}_{g}|^{\varepsilon}+|\mathtt{r}_{g}^{-1}|^{\varepsilon}+\lVert\mathtt{b}_{g}\rVert^{\varepsilon}d\mu(g)<\infty.$
We assume that $\lambda$ is strongly irreducible. This means that for every set $E\subseteq\mathbb{R}^{d}$ which is a finite union of affine subspaces and satisfies $\phi E=E$ for all $\phi\in\operatorname{supp}\lambda$, we have $E=\mathbb{R}^{d}$.
Self-similar measure $\sigma$. We fix a probability measure $\sigma$ on $\mathbb{R}^{d}$ which is $\lambda$-stationary, i.e.
$\sigma=\int_{\operatorname{Sim}(\mathbb{R}^{d})^{+}}\phi_{\star}\sigma\,\mathrm{d}\lambda(\phi).$
Lyapunov exponent. Let $\operatorname{Ad}:G\to\operatorname{Aut}(\mathfrak{g})$ be the adjoint representation. The quantity $\ell$ given by
(2.1) $\ell=-\int_{P^{\prime}}\log\mathtt{r}_{g}\,\mathrm{d}\mu(g)$
is the top Lyapunov exponent associated to $\operatorname{Ad}_{\star}\mu$.
By a theorem of Bougerol-Picard [12, Theorem 2.5], the existence of a $\lambda$-stationary probability measure is equivalent to the condition $\ell>0$, i.e. the random walk on $\mathbb{R}^{d}$ driven by $\lambda$ is contractive in average. Moreover, in this case, the $\lambda$-stationary probability measure is unique, see [12, Corollary 2.7].
Finite time approximation. For any $n\in\mathbb{N}$, let $\sigma^{(n)}=\lambda^{*n}*\delta_{\boldsymbol{0}}$ be a probability measure on $\mathbb{R}^{d}$, where $\delta_{\boldsymbol{0}}$ is the Dirac measure at $\boldsymbol{0}\in\mathbb{R}^{d}$. Note that $\sigma^{(n)}$ is the image measure of $\mu^{*n}$ under the map $g\in P^{\prime}\mapsto\mathtt{b}_{g}\in\mathbb{R}^{d}$. It is known that $\sigma^{(n)}$ converges to $\sigma$ exponentially fast. More precisely, denote by $\operatorname{Lip}(\mathbb{R}^{d})$ the space of bounded Lipschitz functions on $\mathbb{R}^{d}$ equipped with the Lipschitz norm:
$\lVert f\rVert_{\operatorname{Lip}}=\lVert f\rVert_{\infty}+\sup_{\boldsymbol{s_{1}}\neq\boldsymbol{s}_{2}}\frac{|f(\boldsymbol{s}_{1})-f(\boldsymbol{s}_{2})|}{\lVert\boldsymbol{s}_{1}-\boldsymbol{s}_{2}\rVert}.$
We then have by [7, Lemma 2.2] [^3] The proof of [7] is formulated for $d=1$, but it easily carries to arbitrary $d$ by using Lemma 3.1 below in the place of [7, Lemma 2.1 (i)].:
Lemma 2.1 ([7]).
There exist constants $C,\varepsilon>0$ such that for all $n\geq 0$, $f\in\operatorname{Lip}(\mathbb{R}^{d})$, we have
$|\sigma^{(n)}(f)-\sigma(f)|\leq Ce^{-\varepsilon n}\|f\|_{\operatorname{Lip}}.$
Intervals. For real numbers $a<b$, we write $\llbracket a,b\rrbracket$ to denote $\mathbb{Z}\cap[a,b]$. Similarly, we set $\rrbracket a,b\rrbracket:=\mathbb{Z}\cap(a,b]$. We also write $\mathbb{N}_{\geq a}:=\mathbb{N}\cap[a,+\infty)$.
[…]
3. Regularity of self-similar measures
We recall that $\sigma$ has a finite moment of positive order. We then establish that $\sigma$ cannot be concentrated near proper affine subspaces, and more generally near proper algebraic subvarieties of $\mathbb{R}^{d}$.
We start with the control of the tail probabilities. It can be seen as a non-concentration property near infinity.
Lemma 3.1 (Finite moment).
There exists $\gamma>0$ such that
$\int_{\mathbb{R}^{d}}\lVert s\rVert^{\gamma}\,\mathrm{d}\sigma(s)<\infty.$
Proof.
As $\lambda$ has finite exponential moment and is contractive in average, this follows from Kloeckner [39, Theorem 3.1 & Lemma 3.9]. ∎
We now provide a Hölder control on the mass granted by $\sigma$ to neighborhoods of affine subspaces.
Proposition 3.2 (Non-concentration near affine subspaces).
There exists $C,c>0$ such that for every $\varepsilon>0$, every proper affine subspace ${\mathcal{L}}\subsetneq\mathbb{R}^{d}$,
$\sigma({\mathcal{L}}^{(\varepsilon)})\leq C\varepsilon^{c},$
where ${\mathcal{L}}^{(\varepsilon)}$ is the $\varepsilon$-neighborhood of $\mathcal{L}$} in $\mathbb{R}^{d}$.
In the case where $\sigma$ arises from a self-similar IFS which is finite, contractive, and satisfies the open set condition, the result is a consequence of [33, Lemmas 8.2, 8.3]. In [33], the argument is roughly as follows. Assume for simplicity $\lambda$ has equal contraction ratios $\varepsilon\in(0,1)$. Consider $\underline{\phi}=(\phi_{i})_{i\geq 1}\sim\lambda^{\otimes\mathbb{N}}$ and observe $x_{{\underline{\phi}}}:=\lim_{k}\phi_{1}\dots\phi_{k}(0)$ has law $\sigma$. As $\|x_{{\underline{\phi}}}-\phi_{1}\dots\phi_{n}(0)\|\ll\varepsilon^{n}$, the inclusion $x_{{\underline{\phi}}}\in{\mathcal{L}}^{(\varepsilon^{n})}$ implies for every $i\leq n$ that $\phi_{1}\dots\phi_{i}(0)\in{\mathcal{L}}^{(O(\varepsilon^{i}))}$, i.e. $\phi_{i}(0)\in(\phi_{1}\dots\phi_{i-1})^{-1}{\mathcal{L}}^{(O(\varepsilon))}$. Up to taking $\varepsilon$ small enough (by considering $\mu^{*n}$ instead of $\mu$), the probability of the latter event conditionally to previous steps is smaller than $1/2$, whence the desired decay.
In our situation, the support of $\sigma$ may be unbounded, hence knowing that $x_{{\underline{\phi}}}$ is close to $\mathcal{L}$} does not mean that all steps leading to $x_{{\underline{\phi}}}$ will be as well. To deal with this difficulty, we propose an induction scheme that both takes into the count the position of $\phi_{1}\dots\phi_{n}(0)$ with respect to $\mathcal{L}$} but also how far $\mathcal{L}$} is from the origin.
Proof.
Given $s\in(0,1)$ and $T\in\mathbb{R}_{\geq 0}$, we set
$I_{s,T}:=\sup\{\sigma(L^{(s)})\,:\,{\mathcal{L}}\subsetneq\mathbb{R}^{d}\text{ affine with }\operatorname{dist}(0,L)\geq T\}.$
We will show that for some $C,c>0$ depending on $\sigma$ only, for all $s\in(0,1)$, $T\in\mathbb{R}_{\geq 0}$, we have
$I_{s,T}\leq Cs^{c}(1+T)^{-c}.$
Taking $T=0$, we obtain in particular Proposition˜3.2. (In fact there is an equivalence, using that $\sigma$ has finite positive moment - Lemma˜3.1 - and interpolation).
We exhibit some relations between the terms $I_{s,T}$ that will allow to perform an inductive argument. We consider parameters $\varepsilon,c\in(0,1)$ to be chosen later depending on $\sigma$. We aim to show a bound of the form:
(3.1) $\forall s\geq\varepsilon^{k},\,T\geq 0,\,\,\,\,\,\,\,I_{s,T}\leq C_{k}s^{c}(1+T)^{-c}.$
where $C_{k}$ can be formulated in terms of $\sigma,\varepsilon,c$. We argue by induction on $k$. The case $k=0$ is clear because $\sigma$ has a moment of positive order (Lemma˜3.1).
We assume now the result holds for $1,\dots,k-1$, and establish it for $k$. We write $\tau_{\varepsilon}:(\operatorname{Sim}(\mathbb{R}^{d})^{+})^{\mathbb{N}}\rightarrow\mathbb{N}\cup\{\infty\}$ the stopping time defined by
$\tau_{\varepsilon}({\underline{\phi}})=\inf\{n\geq 1\,:\,\mathtt{r}_{\phi_{1}\dots\phi_{n}}<\varepsilon\}$
As $\lambda$ is contracting in average, we have that $\tau_{\varepsilon}$ is $\lambda^{\mathbb{N}}$-almost surely finite. We set $\lambda^{*\tau_{\varepsilon}}$ the distribution of $\phi_{1}\dots\phi_{\tau_{\varepsilon}({\underline{\phi}})}$ as ${\underline{\phi}}\sim\lambda^{\mathbb{N}}$. It is known that $\sigma$ is $\lambda^{*\tau_{\varepsilon}}$-stationary, see e.g [3, Lemma A.2]. By the finite exponential moment of $\lambda$, it can be shown that the variable $\varepsilon/\mathtt{r}_{\phi}$ where $\phi\sim\lambda^{*\tau_{\varepsilon}}$ has a moment of positive order independently of $\varepsilon$: there exists $\gamma=\gamma(\sigma)>0$ such that $\sup_{\varepsilon>0}\int(\varepsilon/\mathtt{r}_{\phi})^{\gamma}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)<2$, see e.g. the proof of [4, Proposition A.18].
Recalling the notation $\mathtt{b}_{\phi}=\phi(0)$. Since $\lambda$ is strongly irreducible, $\sigma$ gives zero mass to proper affine subspaces. Combined with Lemmas 2.1, 3.1, this allows to introduce (small enough) $\delta,\gamma^{\prime}\in(0,1/2)$ depending on $\sigma$ only, and such that
$\sup_{\varepsilon\leq\delta}\sup_{{\mathcal{L}}}\lambda^{*\tau_{\varepsilon}}\{\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\leq\delta\}<\frac{1}{10^{4}},$
and for all $T\geq\delta^{-1}/2$,
$\sup_{\varepsilon\leq\delta}\lambda^{*\tau_{\varepsilon}}\{\phi\,:\,\|\mathtt{b}_{\phi}\|\geq T\}<\frac{1}{10^{4}}(1+T)^{-\gamma^{\prime}}.$
Let $s\in[\varepsilon^{k+1},\varepsilon^{k})$, $T\geq 0$, and consider $\mathcal{L}$} such that $\operatorname{dist}(0,{\mathcal{L}})\geq T$. Set $T^{\prime}=\delta^{2}\max(1,T)$. Using $\lambda^{*\tau_{\varepsilon}}*\sigma=\sigma$ and distinguishing according to the position of $\mathtt{b}_{\phi}$, we have
$\sigma({\mathcal{L}}^{(s)})=\underbrace{\int_{\{\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\geq T^{\prime}\}}\phi_{\star}\sigma({\mathcal{L}}^{(s)})\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)}_{A}\,\,\,+\,\,\,\underbrace{\int_{\{\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})<T^{\prime}\}}\phi_{\star}\sigma({\mathcal{L}}^{(s)})\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)}_{B}.$
Using the induction hypothesis, we establish an upper bound for $A$. More precisely, noting that $\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})\geq T^{\prime}$ implies $\operatorname{dist}(0,{\phi}^{-1}{\mathcal{L}})\geq\mathtt{r}_{\phi}^{-1}T^{\prime}$, then using that $\mathtt{r}_{\phi}^{-1}s\geq\varepsilon^{k}$ for all $\phi\in\operatorname{supp}\lambda^{*\tau_{\varepsilon}}$ to apply the induction hypothesis, we find
$A$ $\leq\int_{\operatorname{Sim}(\mathbb{R}^{d})^{+}}I_{\mathtt{r}_{\phi}^{-1}s,\mathtt{r}_{\phi}^{-1}T^{\prime}}\,\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)$
$\leq C_{k-1}\int_{\operatorname{Sim}(\mathbb{R}^{d})^{+}}(\mathtt{r}_{\phi}^{-1}s(1+\mathtt{r}_{\phi}^{-1}T^{\prime})^{-1})^{c}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)$
$\leq C_{k-1}2\delta^{-2c}s^{c}(1+T)^{-c}.$
We now bound $B$. We have
$B$ $\leq\int_{\{\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})<T^{\prime}\}}I_{\mathtt{r}_{\phi}^{-1}s,0}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)$
$\leq C_{k-1}s^{c}\varepsilon^{-c}\int_{\{\phi\,:\,\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})<T^{\prime}\}}(\varepsilon/\mathtt{r}_{\phi})^{c}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)$
$\leq C_{k-1}s^{c}\varepsilon^{-c}2\sqrt{\lambda^{*\tau_{\varepsilon}}\{\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})<T^{\prime}\}}$
where the last inequality relies on Cauchy-Schwarz, and assumes $c\leq\gamma/2$ to guarantee $\int_{\operatorname{Sim}(\mathbb{R}^{d})^{+}}(\varepsilon/\mathtt{r}_{\phi})^{2c}\,\mathrm{d}\lambda^{*\tau_{\varepsilon}}(\phi)\leq 2$. Note also, by distinguishing the cases $T^{\prime}\leq\delta$ (i.e. $T\leq\delta^{-1}$) and $T^{\prime}>\delta$ (i.e. $T>\delta^{-1}$), that the definition of $\delta$ yields
$\sqrt{\lambda^{*\tau_{\varepsilon}}\{\operatorname{dist}(\mathtt{b}_{\phi},{\mathcal{L}})<T^{\prime}\}}\leq\frac{1}{50}(1+\delta T)^{-\gamma^{\prime}/2}.$
Finally
$B\leq C_{k-1}s^{c}\varepsilon^{-c}\frac{1}{25}(1+\delta T)^{-\gamma^{\prime}/2}.$
From the above, we have obtained
$I_{s,T}\leq C^{\prime}s^{c}(1+T)^{-c}$
where
$C^{\prime}$ $=2(\delta^{-2c}+\varepsilon^{-c}\frac{1}{50}(1+T)^{c}(1+\delta T)^{-\gamma^{\prime}/2})C_{k-1}$
$\leq 2\delta^{-2c}(1+\varepsilon^{-c}\frac{1}{50})C_{k-1}$
where the second inequality assumed $c<\gamma^{\prime}/2$. We now check that $c,\delta$ and $\varepsilon$ can in fact be chosen so that
$2\delta^{-2c}(1+\varepsilon^{-c}\frac{1}{50})\leq\varepsilon^{-c/2}.$
Indeed, it suffices to select $c\lll_{\delta}1$ such that $2\delta^{-2c}<21/10$, then $\varepsilon$ such that $\varepsilon^{c/2}=2/5$. We have thus established
$C^{\prime}\leq\varepsilon^{-c/2}C_{k-1}.$
The above justifies (3.1), with constants $C_{k}$ given by $C_{k}=\varepsilon^{-kc/2}C_{0}$ where $C_{0}=C_{0}(\sigma)>1$. It follows that for every $s>0,T\geq 0$,
$I_{s,T}\leq C^{\prime}_{0}s^{c/3}(1+T)^{-c},$
where $C_{0}^{\prime}=C_{0}\varepsilon^{-c}$ is bounded depending on $\sigma$ only, and the proof of the proposition is complete. ∎
We upgrade Proposition˜3.2 into a Hölder control on the concentration of $\sigma$ near algebraic subvarieties of $\mathbb{R}^{d}$. Given $l\in\mathbb{N}$, we set $\mathcal{P}_{d,l}$ the vector space of real polynomial functions on $\mathbb{R}^{d}$ of degree at most $l$. We equip $\mathcal{P}_{d,l}$ with the supremum norm on the coefficients, which we write $\|\cdot\|$.
Theorem 3.3 (Non-concentration near subvarieties).
For every $l\in\mathbb{N}$, $P\in\mathcal{P}_{d,l}$ with $\|P\|\geq 1$, and $\varepsilon>0$, we have
$\sigma(\boldsymbol{s}\in\mathbb{R}^{d}\,:\,|P(\boldsymbol{s})|\leq\varepsilon)\leq C\varepsilon^{c},$
where $C,c>0$ depend only on $\sigma,l$.
The idea of the proof of ˜3.3 is to use the self-similarity of $\sigma$ to write $\sigma$ as a convex combination of measures $(\sigma_{j})_{j}$ obtained from $\sigma$ by pushing via affine maps, and with each $\sigma_{j}$ living at scale $\varepsilon^{3/4}$, say roughly supported in $B(x_{j},\varepsilon^{3/4})$. For each $\sigma_{j}$, we may approximate the polynomial map $P$ by its Taylor expansion up to order $1$ at $x_{j}$, and apply non-concentration near affine hyperplanes to deduce the estimate in ˜3.3. In fact, to make this argument work, we also need to make sure that most of the $x_{j}$’s are located where the gradient $\nabla P$ of $P$ is not too small. But $\nabla P$ is polynomial of smaller degree, and the distribution of the $x_{j}$’s resembles that of $\sigma$, whence we may guarantee this using an inductive approach. A related strategy is exploited in [33, Section 7] in the context of absolutely decaying measures.
Remark.
In the same manner, one can show that $\sigma$ is not concentrated near submanifolds $M$ of $\mathbb{R}^{d}$ such that $\dim M<d$, as long as $M$ is not too badly approximated by its tangent subspaces (e.g. if $\exp_{x}:B^{T_{x}M}_{1}\rightarrow M$ has uniformly bounded order $2$ derivatives for all $x$ in the manifold).
Proof of ˜3.3.
We argue by induction on the degree $l$. The case $l=0$ is clear. We assume the result known for degrees $0,\dots,l-1$ and prove it for $l\geq 1$. We fix $P\in{\mathcal{P}}_{d,l}$ with $\|P\|=1$, given $\varepsilon\in(0,1)$, we set
$E_{\varepsilon}:=\{\boldsymbol{s}\in\mathbb{R}^{d}:|P(\boldsymbol{s})|\leq\varepsilon\}.$
We use the shorthand $B_{r}=B^{\mathbb{R}^{d}}_{r}$. Given a function $f:\mathbb{R}^{d}\rightarrow V$ where $V$ is a normed vector space, we write $\|f\|_{B_{r}}:=\sup_{x\in B_{r}}\|f(x)\|$ the supremum norm of the restriction $f_{|B_{r}}$. Note that the family $\|\cdot\|_{B_{r}}$ induces a collection of norms on ${\mathcal{P}}_{d,l}$, and these norms are mutually equivalent by finite dimensionality of ${\mathcal{P}}_{d,l}$.
Let $\alpha,\beta\in(0,1)$ be parameters to be specified later, with $\alpha$ depending only on $\sigma,l$, and $\beta$ absolute. For convenience, we will write $R:=\varepsilon^{-\alpha}$ and $\delta=\varepsilon^{\beta}$. We also set
$F_{\delta}:=\{\boldsymbol{s}\in\mathbb{R}^{d}:\|\nabla P(\boldsymbol{s})\|\leq\delta\}$
where $\nabla P:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ refers to the gradient of $P$.
We decompose $\sigma$ as a combination of measures living at scale $\varepsilon^{3/4}$, and group them into 3 categories, distinguishing whether they are centered around a point outside of $B_{R}$, or within $B_{R}$, and in the second case whether the point is in $F_{\delta}$ or not. More formally, recalling the Lyapunov exponent $\ell$ from (2.1), we set $n=\lfloor\frac{3}{4\ell}\lvert\log\varepsilon\rvert\rfloor$, so that for $\phi\sim\lambda^{*n}$, we have $r_{\phi}$ close to $\varepsilon^{3/4}$. We then have by $\lambda$-stationarity of $\sigma$:
$\sigma(E_{\varepsilon})$ $=\int_{\operatorname{Sim}(\mathbb{R}^{d})}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi)=I_{1}+I_{2}+I_{3}$
where
$I_{1}=\int_{\phi(0)\notin B_{R}}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi),\,\,\,\,\,\,\,\,I_{2}=\int_{\phi(0)\in B_{R}\cap F_{\delta}}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi),$
$I_{3}=\int_{\phi(0)\in B_{R}\smallsetminus F_{\delta}}\phi_{\star}\sigma(E_{\varepsilon})\,\mathrm{d}\lambda^{*n}(\phi).$
Combining Lemma˜3.1 and Lemma˜2.1, we have for some $\gamma=\gamma(\lambda)>0$,
(3.2) $I_{1}\ll R^{-\gamma}.$
We now bound $I_{2}$. As a preliminary, note that the assumption $\|P\|=1$ implies $\|\nabla P\|_{B_{R+1}}\ll_{l}R^{l}$. On the other hand, we may assume
(3.3) $\|\nabla P\|\gg_{l}R^{-(1+l)}.$
To see why, note first that $\|P\|=1$ implies $\sup_{B_{1}}|P|\geq\eta$ where $\eta=\eta(d,l)>0$. Now if $\|\nabla P\|_{B_{R}}\leq R^{-1}\eta/4$, we get $\inf_{B_{R}}|P|\geq\eta/2$. In this scenario, provided $\varepsilon\leq\eta/4$, we have $E_{\varepsilon}\subseteq\mathbb{R}^{d}\smallsetminus B_{R}$, and the result follows from the finite moment of $\sigma$ (Lemma˜3.1). Hence we may suppose $\|\nabla P\|_{B_{R}}>R^{-1}\eta/4$, and (3.3) follows by passing to the norm $\|\cdot\|$.
Let $c_{1}>0$. Note that the upper bound $\|\nabla P\|_{B_{R+1}}\ll_{l}R^{l}$ implies that the $\varepsilon^{c_{1}}$-neighborhood of $B_{R}\cap F_{\delta}$ is included in $F_{\delta+O_{l}(\varepsilon^{c_{1}}R^{l})}$. Choosing $c_{1}=c_{1}(\sigma)$ small enough, we deduce from Lemma˜2.1 that
$I_{2}$ $\leq\sigma^{(n)}(B_{R}\cap F_{\delta})\leq\sigma(F_{\delta+O_{l}(\varepsilon^{c_{1}}R^{l})})+O(\varepsilon^{c_{1}}).$
Applying the induction hypothesis and the lower bound (3.3) on $\|\nabla P\|$, we deduce
(3.4) $I_{2}\ll_{l}R^{(1+l)c_{2}}(\delta+\varepsilon^{c_{1}}R^{l})^{c_{2}}+\varepsilon^{c_{1}}$
where $c_{2}=c_{2}(\sigma,l-1)>0$.
We now bound $I_{3}$. Using that $\sigma$ has a finite moment (Lemma˜3.1), and the large deviation principle for $(r_{\phi})_{\phi\sim\lambda^{*n}}$, we find for some $c_{3}=c_{3}(\sigma)>0$,
$\lambda^{*n}\{\phi\,:\,\phi_{\star}\sigma(\mathbb{R}^{d}\smallsetminus B_{\varepsilon^{1/2}}(\phi(0)))\geq\varepsilon^{c_{3}}\}\ll\varepsilon^{c_{3}},$
whence
$I_{3}\leq$ $\int_{\phi(0)\in B_{R}\smallsetminus F_{\delta}}\phi_{\star}\sigma[E_{\varepsilon}\cap B_{\varepsilon^{1/2}}(\phi(0))]\,\mathrm{d}\lambda^{*n}(\phi)+O(\varepsilon^{c_{3}}).$
Consider $\phi$ as in the above integral, and $\boldsymbol{s}\in E_{\varepsilon}\cap B_{\varepsilon^{1/2}}(\phi(0))$. By Taylor expansion, and the fact that $\|\cdot\|_{C^{2}(B_{R+1})}\ll_{l}R^{l}\|\cdot\|$ on ${\mathcal{P}}_{d,l}$, the conditions $\|P\|=1$ and $\|\boldsymbol{s}-\phi(0)\|<\varepsilon^{1/2}$ imply
$|P(\boldsymbol{s})-P(\phi(0))-\big{\langle}\nabla P(\phi(0)),\,\,\boldsymbol{s}-\phi(0)\big{\rangle}|\ll_{l}\varepsilon R^{l}.$
As $|P(\boldsymbol{s})|\leq\varepsilon$, setting $v_{P,\phi}:=P(\phi(0))-\big{\langle}\nabla P(\phi(0)),\,\,\phi(0)\big{\rangle}$, we deduce
$|\big{\langle}\nabla P(\phi(0)),\,\,\boldsymbol{s}\big{\rangle}-v_{P,\phi}|\ll_{l}\varepsilon R^{l}$
But $\|\nabla P(\phi(0))\|>\delta$ by assumption, whence $\boldsymbol{s}$ belongs to the $O_{l}(\varepsilon R^{l}\delta^{-1})$-neighborhood of an affine hyperplane. Applying $\phi^{-1}$, and the non-concentration near affine subspaces from Proposition˜3.2, we obtain
$\phi_{\star}\sigma[E_{\varepsilon}\cap B_{\varepsilon^{1/2}}(\phi(0))]\ll_{l}(\varepsilon R^{l}\delta^{-1}r_{\phi}^{-1})^{c_{4}}$
where $c_{4}=c_{4}(\sigma)>0$. Using the large deviation principle for $(r_{\phi})_{\phi\sim\lambda^{*n}}$, we may integrate to obtain
(3.5) $I_{3}\ll_{l}(\varepsilon^{1/8}R^{l}\delta^{-1})^{c_{4}}+\varepsilon^{c_{3}}.$
In the end, combining (3.2), (3.4), (3.5), and choosing $\delta=\varepsilon^{1/16}$, $R=\varepsilon^{-\alpha}$ with $\alpha\lll_{\sigma,l}1$, we have proven that for $c^{\prime}\lll_{\gamma,c_{1},\dots,c_{4}}1$,
$\sigma(E_{\varepsilon})\ll\varepsilon^{c^{\prime}}.$
This proves the induction step, whence the theorem. ∎
It is easy to deduce from the previous theorem that a product $\sigma^{\otimes k}$ is not concentrated near subvarieties of $\mathbb{R}^{dk}$. We record this observation for future use.
Corollary 3.4.
Let $k,l\geq 1$. Let $P:\mathbb{R}^{dk}\rightarrow\mathbb{R}$ be a polynomial map of degree at most $l$ and such that $\|P\|\geq 1$. Then for every $\varepsilon>0$,
$\sigma^{\otimes k}\{(\boldsymbol{s}_{i})_{i=1}^{k}\in(\mathbb{R}^{d})^{k}\,:\,|P(\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{k})|\leq\varepsilon\}\leq C\varepsilon^{c}$
where $C=C(\sigma,k,l)>1$ and $c=c(\sigma,l)>0$.
Proof.
We argue by induction on $k$. The case $k=1$ is ˜3.3. Now given $k\geq 2$, we assume the result holds up to k1k-1 and prove it for $k$. Consider $Q:\mathbb{R}^{d}\rightarrow{\mathcal{P}}_{d(k-1),l},\boldsymbol{s}\mapsto P(\boldsymbol{s},\,\cdot\,)$, which is a polynomial map whose coordinates have degree at most $l$. Moreover, endowing ${\mathcal{P}}_{d(k-1),l}$ with the standard basis, $Q$ and $P$ have same coefficients, in particular $\|Q\|=\|P\|\geq 1$. It follows from the $k=1$ case that for some $c_{1}=c_{1}(\sigma,l)>0$, for every $\delta>0$,
$\sigma\{\boldsymbol{s}\in\mathbb{R}^{d}\,:\,\|Q(\boldsymbol{s})\|\leq\delta\}\ll_{l}\delta^{c_{1}}.$
We deduce
$\sigma^{\otimes k}\{(\boldsymbol{s}_{i})_{i=1}^{k}\in(\mathbb{R}^{d})^{k}\,:\,|P(\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{k})|\leq\varepsilon\}$
$\leq$ $\sigma^{\otimes k}\{(\boldsymbol{s}_{i})_{i=1}^{k}\in(\mathbb{R}^{d})^{k}\,:\,|P(\boldsymbol{s}_{1},\dots,\boldsymbol{s}_{k})|\leq\varepsilon\text{ and }\|Q(\boldsymbol{s}_{1})\|>\varepsilon^{1/2}\}+O(\varepsilon^{c_{1}/2})$
$\ll$ $_{k,l}\varepsilon^{c_{2}/2}+\varepsilon^{c_{1}/2},$
where the last inequality follows from the induction hypothesis with parameter k1k-1, and exponent $c_{2}=c_{2}(\sigma,l)$>0. This concludes the proof.
Finite time consequences. Recalling that the finite time approximation $\sigma^{(n)}:=\lambda^{*n}*\delta_{\bf 0}$ converges to $\sigma$ exponentially fast, we may transfer the regularity properties of $\sigma$ to $\sigma^{(n)}$ provided we look at scales above an exponentially small threshold.
Lemma 3.5.
For $\gamma\lll 1$ and all $n\geq 1$, we have
  • (i)
    $\int_{\mathbb{R}^{d}}\lVert\boldsymbol{s}\rVert^{\gamma}\,\mathrm{d}\sigma^{(n)}(\boldsymbol{s})\ll 1,$
  • (ii)
    $\forall\varepsilon>e^{-n},\quad\sup_{\boldsymbol{s}\in\mathbb{R}^{d}}\sigma^{(n)}(\boldsymbol{s}+[-\varepsilon,\varepsilon]^{d})\ll\varepsilon^{\gamma}.$
Moreover, for $l\geq 1$, $c\lll_{l}1$, $P\in{\mathcal{P}}_{d,l}$ with $\|P\|=1$, $n\geq 1$, $\varepsilon>e^{-n}$, we have
  • (iii)
    $\sigma^{(n)}\{\boldsymbol{s}\in\mathbb{R}^{d}\,:\,|P(\boldsymbol{s})|\leq\varepsilon\}\ll_{l}\varepsilon^{c}.$
Proof.
In view of Lemma˜2.1, Items (i) and (ii) respectively follow from Lemma˜3.1 and Proposition˜3.2, see the proof of [7, Lemma 2.3] for more details. For item (iii), given $R>1$, $\varepsilon>0$, set $E_{R,\varepsilon}:=\{\boldsymbol{s}\in B_{R}\,:\,|P(\boldsymbol{s})|\leq\varepsilon\}$ where $B_{R}:=B^{\mathbb{R}^{d}}_{R}$. By Lemma˜2.1, we have for $1\ggg\varepsilon>e^{-n}$, for some $\gamma=\gamma(\lambda)\in(0,1)$,
$\sigma^{(n)}(E_{R,\varepsilon})\leq\sigma(E_{2R,\,\varepsilon+\varepsilon^{\gamma}\|\nabla P\|_{B_{2R}}})+e^{-\gamma n}.$
Observe $\|\nabla P\|_{B_{2R}}\ll R^{l}$. Therefore, taking $R=\varepsilon^{-\alpha}$ with $\alpha>0$, we have by ˜3.3
$\sigma(E_{2R,\,\varepsilon+\varepsilon^{\gamma}\|\nabla P\|_{B_{2R}}})\ll_{l}\varepsilon^{(\gamma-l\alpha)c}$
where $c=c(\sigma,l)>0$. The result follows by taking $\alpha\lll_{l}1$, and applying Lemma˜3.5 (i) to allow restriction to $B_{\varepsilon^{-\alpha}}$. ∎
4. Effective recurrence of the $\mu$-walk
In this section, we establish that the $n$-step distribution of the $\mu$-random walk on $X$ is not concentrated near infinity, provided $n$ is large enough in terms of the starting point. We recall notations have been set up in Section˜2, in particular $\operatorname{inj}(x)$ denotes the injectivity radius of $X$ at the point $x$.
Proposition 4.1.
There exist constants $C,c>0$ such that for every $x\in X$, $n\in\mathbb{N}$, $\rho>0$,
$\mu^{*n}*\delta_{x}\{\operatorname{inj}\leq\rho\}\ll\rho^{c}(e^{-cn}\operatorname{inj}(x)^{-C}+1).$
For $d=1$, a short self-contained proof is given in [7, Section 2.3]. For arbitrary $d$, we explain how to deduce Proposition˜4.1 from Prohaska-Sert-Shi [47], which is itself inspired by the works of Benoist-Quint [8], Eskin-Margulis [23], Eskin-Margulis-Mozes [24].
Lemma 4.2.
Denote by $H_{\mu}$ the Zariski closure of the semigroup generated by $\operatorname{supp}\mu$. Then $U\subseteq H_{\mu}$.
Proof.
Recall that every $g\in P^{\prime}$ can be written uniquely as $g=k_{g}^{-1}a(\mathtt{r}_{g}^{-1})u(\mathtt{b}_{g})$, where $k_{g}=\operatorname{diag}(O_{g},1)\in K^{\prime},\mathtt{r}_{g}>0,\mathtt{b}_{g}\in\mathbb{R}^{d}$. We first deal with a particular case of the lemma.
Case ()(*): there exists $g^{\prime}\in\operatorname{supp}\mu$ such that $\mathtt{r}_{g^{\prime}}\in(0,1)$ and $\mathtt{b}_{g^{\prime}}=0$. In this case, choose a sequence $n_{i}\to+\infty$ such that $k_{g^{\prime}}^{-n_{i}}\to\operatorname{Id}$ as $i\to+\infty$. Then for every $g\in\operatorname{supp}\mu$, we have $\lim_{i\to+\infty}g^{\prime-n_{i}}gg^{\prime n_{i}}=k_{g}^{-1}a(\mathtt{r}_{g}^{-1})$, from which it follows that
(4.1) $k_{g}^{-1}a(\mathtt{r}_{g}^{-1})\in H_{\mu},\quad\quad\quad u(\mathtt{b}_{g})\in H_{\mu}.$
Write $S$ the set of vectors $\boldsymbol{s}\in\mathbb{R}^{d}$ such that $u(\boldsymbol{s})\in H_{\mu}$. Note $S$ is a Zariski-closed subgroup of $\mathbb{R}^{d}$, i.e. $S$ is a subspace. Using that $U\cap H_{\mu}$ is normalized by $H_{\mu}$, Equation (4.1), and the relation $a(\mathtt{r}_{g})k_{g}u(\boldsymbol{s})k_{g}^{-1}a(\mathtt{r}_{g}^{-1})u(\mathtt{b}_{g})=u(\mathtt{r}_{g}O_{g}\boldsymbol{s}+\mathtt{b}_{g})$, we get that $S$ is invariant under $\operatorname{supp}\lambda$. By irreducibility of $\lambda$, we deduce $S=\mathbb{R}^{d}$. This finishes the proof of Case ()(*).
General case. We now reduce the general case to Case ()(*). As $\lambda$ is contractive in average (see discussion after (2.1)), there exists $g^{\prime}\in\operatorname{supp}\mu$ such that $\mathtt{r}_{g^{\prime}}\in(0,1)$. In particular, the vector
$\boldsymbol{s}_{0}=(\operatorname{Id}_{\mathbb{R}^{d}}-\mathtt{r}_{g^{\prime}}O_{g^{\prime}})^{-1}\mathtt{b}_{g^{\prime}}$
is well defined. By direct computation, we find that
$u(\boldsymbol{s}_{0})g^{\prime}u(-\boldsymbol{s}_{0})=k_{g^{\prime}}^{-1}a(\mathtt{r}_{g^{\prime}}^{-1}).$
The measure $\mu^{\prime}:=\delta_{u(\boldsymbol{s}_{0})}*\mu*\delta_{u(-\boldsymbol{s}_{0})}$ on $P^{\prime}$ corresponds to some (strongly) irreducible randomized self-similar IFS $\lambda^{\prime}$. By the analysis of Case ()(*), we know that $H_{\mu^{\prime}}\supseteq U$. On the other hand, $H_{\mu^{\prime}}=u(\boldsymbol{s}_{0})H_{\mu}u(-\boldsymbol{s}_{0})$, whence $H_{\mu}\supseteq U$ as well. ∎
The previous lemma allows us to apply [47] to obtain some Margulis function, i.e. a proper positive function on $X$ which is uniformly contracted by the random walk and satisfies some growth control under the action of $G$.
Lemma 4.3 (Height function [47]).
There exists a function $\beta:X\to[1,+\infty)$ which is proper and satisfies:
  • (1)
    Contraction property: There exist $m\in\mathbb{N}$ and $\theta,M>0$ such that for all $x\in X$,
    $\mu^{*m}*\delta_{x}(\beta)\leq e^{-\theta}\beta(x)+M.$
  • (2)
    Growth control: $\beta(gx)\leq\|\operatorname{Ad}(g)\|^{O(1)}\beta(x)$ for all $g\in G,x\in X$.
Proof.
The combination of Lemma˜4.2 and [47, Corollary 3.8] guarantees that $\mu$ is $G$-expanding in the sense of [47, Definition 2.7]. This allows to apply [47, Theorem 6.1] to obtain the desired function $\beta$.
We justify that the above Margulis function can be compared to the injectivity radius.
Lemma 4.4.
For any proper function $\Upsilon:X\to[1,+\infty)$ satisfying properties (1) (2) of Lemma˜4.3, there exist $C,c>0$ such that for all $x\in X$,
$\Upsilon(x)^{-C}\ll\operatorname{inj}(x)\ll\Upsilon(x)^{-c}.$
Proof.
Using the assumptions on $\Upsilon$, the comparison between $\operatorname{inj}(x)$ and $\Upsilon(x)$ follows from the same argument as for [6, Lemmas 3.13, 3.14]. ∎
We are now able to conclude the proof of the effective recurrence property.
Proof of Proposition˜4.1.
Let $\beta$ and $m,\theta,M$ as in Lemma˜4.3. We first replace $\beta$ by a suitable $\beta^{\prime}$ satisfying the contraction property with $m=1$. For that, given $f:X\rightarrow\mathbb{R}_{\geq 0}$, set $P_{\mu}f=\int_{G}f(g.)\,\mathrm{d}\mu(g)$. Let $\kappa>0$, consider $\beta^{\prime}=\beta+e^{\kappa}P_{\mu}\beta+\dots+e^{(m-1)\kappa}P_{\mu}^{m-1}\beta$. By Lemma˜4.3 (1), taking $\kappa:=\theta/m$, we have
(4.2) $P_{\mu}\beta^{\prime}\leq e^{-\kappa}\beta^{\prime}+e^{(m-1)\kappa}M.$
Now, iterating (4.2), we find for every $n\geq 0$,
$P_{\mu}^{n}\beta^{\prime}\leq e^{-\kappa n}\beta^{\prime}+M^{\prime}$
where $M^{\prime}=e^{(m-1)\kappa}M/(1-e^{-\kappa})$. Using the Markov inequality, we deduce for every $\rho>0$,
$\mu^{*n}*\delta_{x}\{y:\beta^{\prime}(y)>\rho^{-1}\}\leq(e^{-\kappa n}\beta^{\prime}(x)+M^{\prime})\rho.$
The proposition then follows from the comparison Lemma˜$4.$ ∎
As a direct corollary of Proposition˜4.1, we bound the Haar measure of cusp neighborhoods. This estimate will be useful in Section˜8.
Lemma 4.5.
There are constants $C>1$ and $c>0$ depending on $\Lambda$ such that
  1. (1)
    for all $\rho>0$,
    $m_{X}\{\operatorname{inj}\leq\rho\}\leq C\rho^{c};$
  2. (2)
    writing $x_{0}=\Lambda/\Lambda$ for the basepoint of $X$, we have for all $r>0$,
    $m_{X}\{\operatorname{dist}(\cdot,x_{0})\geq r\}\leq Ce^{-cr}.$
Proof.
The Haar measure $m_{X}$ is an ergodic $\mu$-stationary measure. Hence for $m_{X}$-almost every $x\in X$, the sequence $(\mu^{*n}*\delta_{x})_{n\geq 0}$ converges to $m_{X}$ in Cesáro average for the weak-* topology. Then the first estimate follows immediately from Proposition˜4.1. Note the constants $C$, $c$ only depend on $\Lambda$ (and not $\mu$) because $\mu$ does not play a role in the statement.
By [6, Lemma 3.14], we have for $x\in X$,
(4.3) $\lvert\log\operatorname{inj}(x)\rvert-1\ll\operatorname{dist}(x,x_{0})\ll\lvert\log\operatorname{inj}(x)\rvert+1.$
The second estimate then follows from the first. ∎
5. Positive dimension
In this section, we show that the $n$-step distribution of the $\mu$-random walk has positive dimension provided we look at scales above an exponentially small threshold and $n$ is large enough in terms of the starting point.
Proposition 5.1 (Positive dimension).
There exists $A,\kappa>0$ such that for every $x\in X$, $\rho>0$, $n\geq\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}(x)\rvert$, we have
$\forall y\in X,\quad\mu^{*n}*\delta_{x}(B_{\rho}y)\ll\rho^{\kappa}.$
The case $d=1$ corresponds to [7, Proposition 3.1]. For arbitrary $d$, the argument of [7] goes through with a few adaptations to deal with the rotation component $K^{\prime}$. We provide the proof for completeness.
Proof.
Let $\kappa>0$ be a parameter to be specified later. Let $x\in X$, $\rho\in(0,1)$, $n\geq\lvert\log\rho\rvert$. Assume by contradiction that there exists some $y\in X$ such that
(5.1) $\mu^{*n}*\delta_{x}(B_{\rho}y)>\rho^{\kappa}.$
Let $\alpha=\frac{1}{10(\ell+1)}$ and $m=\lfloor\alpha\lvert\log\rho\rvert\rfloor$. Write
$\mu^{*n}*\delta_{x}=\mu^{*m}*\mu^{*(n-m)}*\delta_{x}$
and
$Z:=\bigl{\{}\,z\in X\,:\,\mu^{*m}*\delta_{z}(B_{\rho}y)\geq\rho^{2\kappa}\,\bigr{\}}.$
Then (5.1) implies that $\mu^{*(n-m)}*\delta_{x}(Z)\geq\rho^{2\kappa}$, provided $\rho\lll_{\kappa}1$.
We are going to show that points in $Z$ have small injectivity radius. Fix $z\in Z$. By definition we have
(5.2) $\mu^{*m}\{g:gz\in B_{\rho}y\}\geq\rho^{2\kappa}.$
On the other hand, by the large deviation principle, there exists $\varepsilon=\varepsilon(\mu)>0$ such that for $\rho\lll 1$,
(5.3) $\mu^{*m}\{g\,:\,\log\mathtt{r}_{g}\in[-(\ell+1)m,-(\ell-1)m]\}\geq 1-\rho^{\alpha\varepsilon}.$
Furthermore, considering $\gamma=\gamma(\mu)>0$ as in Lemma˜3.5, we have for all $\rho\lll_{\kappa}1$,
(5.4) $\mu^{*m}\{g\,:\,\lVert\mathtt{b}_{g}\rVert\leq\rho^{-4\gamma^{-1}\kappa}\}\geq 1-\rho^{3\kappa}.$
Let $C>1$ be a large parameter to be specified below depending on $\mu$ only. Partition the product set
$K^{\prime}\times[-(\ell+1)m,-(\ell-1)m]\times[-\rho^{-4\gamma^{-1}\kappa},\rho^{-4\gamma^{-1}\kappa}]^{d}$
into subsets $(S_{i})_{i\in I}$ of the form $S_{i}=S_{i,1}\times S_{i,2}\times S_{i,3}$ where each $S_{i,j}$ has diameter less than $\rho^{C\kappa}$. Note we can arrange the number of elements in the partition to be controlled via
$|I|\ll\rho^{-d(d-1)C\kappa/2}\cdot m\rho^{-C\kappa}\cdot\rho^{-4d\gamma^{-1}\kappa}\rho^{-dC\kappa}.$
By the pigeonhole principle and (5.2), (5.3), (5.4), there exists $i_{0}\in I$ such that the set
$E:=\{g\,:\,gz\in B_{\rho}y\text{ and }(k_{g},\log\mathtt{r}_{g},\mathtt{b}_{g})\in S_{i_{0}}\}$
satisfies
(5.5) $\mu^{*m}(E)\geq\frac{\rho^{2\kappa}-\rho^{\alpha\varepsilon}-\rho^{3\kappa}}{|I|}\geq\rho^{10d^{2}C\kappa},$
provided that $C\geq\gamma^{-1}$, $3\kappa\leq\alpha\varepsilon$ and $\rho\lll_{\kappa}1$.
Let $g_{1},g_{2}\in E$. Note that $\operatorname{dist}(g_{1}z,g_{2}z)\ll\rho$. By the choice of $m$ and the bounds on $\mathtt{r}_{g_{1}},\mathtt{b}_{g_{1}}$, we have $\lVert\operatorname{Ad}(g_{1}^{-1})\rVert\leq\rho^{-1/2}$ provided $\kappa\lll 1$. It follows that
(5.6) $\operatorname{dist}(z,g_{1}^{-1}g_{2}z)\ll\lVert\operatorname{Ad}(g_{1}^{-1})\rVert\rho\ll\rho^{1/2}.$
Using that $K^{\prime}$ and $A^{\prime}$ commute, we can further write $g_{1}^{-1}g_{2}$ as
$g_{1}^{-1}g_{2}=u(-\mathtt{b}_{g_{2}})hu(\mathtt{b}_{g_{2}})\quad\text{where }h=u(\mathtt{b}_{g_{2}}-\mathtt{b}_{g_{1}})a(\mathtt{r}_{g_{1}}\mathtt{r}_{g_{2}}^{-1})k_{g_{1}}k_{g_{2}}^{-1}.$
To deduce an estimate on the injectivity radius of $X$ at $z$, we now show that $g_{1},g_{2}$ can be chosen so that $g_{1}^{-1}g_{2}$ is not too close to $\operatorname{Id}$, but still at a distance less than a small power of $\rho$ from one another. First, using the non-concentration estimate Lemma˜3.5 (ii) and (5.5), we can choose $g_{1},g_{2}\in E$ such that
(5.7) $\lVert\mathtt{b}_{g_{1}}-\mathtt{b}_{g_{2}}\rVert\geq\rho^{11\gamma^{-1}d^{2}C\kappa},$
provided $\kappa\ll_{C}1$ to ensure that $\rho^{11\gamma^{-1}d^{2}C\kappa}\geq e^{-m}\simeq\rho^{\alpha}$ as required in Lemma˜3.5, and $\rho\lll_{\kappa}1$. For $g_{1},g_{2}\in E$ satisfying (5.7), we have
$\operatorname{dist}(h,\operatorname{Id})\simeq\lVert k_{g_{1}}-k_{g_{2}}\rVert+\lVert\mathtt{b}_{g_{1}}-\mathtt{b}_{g_{2}}\rVert+|1-\mathtt{r}_{g_{1}}\mathtt{r}_{g_{2}}^{-1}|\in[\rho^{11\gamma^{-1}d^{2}C\kappa},10\rho^{C\kappa}].$
Recalling that $\lVert\mathtt{b}_{g_{2}}\rVert\leq\rho^{-4\gamma^{-1}\kappa}$, we get
(5.8) $\rho^{1/4}\ll\rho^{(11d^{2}C+4(d+1))\gamma^{-1}\kappa}\ll\operatorname{dist}(g_{1}^{-1}g_{2},\operatorname{Id})\ll\rho^{(C-4(d+1)\gamma^{-1})\kappa},$
where the lower bound assumes $\kappa\ll_{C}1$. Provided $C>8(d+1)\gamma^{-1}$, Equations (5.6), (5.8) yield $\operatorname{inj}(z)\ll\rho^{C\kappa/2}+\rho^{1/2}$. When $\kappa\lll_{C}1$ and $\rho\lll_{\kappa}1$, this gives
$\operatorname{inj}(z)\leq\rho^{C\kappa/4}.$
In conclusion, we have shown that for $C\ggg 1$, for $\kappa\lll_{C}1$, $\rho\lll_{\kappa}1$, and $n\geq m=\lfloor\alpha\lvert\log\rho\rvert\rfloor$, we have
$\mu^{*(n-m)}*\delta_{x}\{\operatorname{inj}\leq\rho^{C\kappa/4}\}\geq\rho^{2\kappa}.$
By the effective recurrence statement from Proposition˜4.1, this is absurd if $n-m\ggg{\lvert\log\operatorname{inj}(x)\rvert}$. The proof of the proposition is complete. ∎
6. Dimensional Bootstrap
In this section, we show that the $n$-step distribution $\mu^{*n}*\delta_{x}$ becomes high dimensional in $X$ exponentially fast as $n$ goes to infinity. The next definition will be useful.
Definition 6.1 (Robust measure).
Let $\alpha>0,\tau\geq 0$ and $I\subseteq(0,1]$. A Borel measure $\nu$ on $X$ is said to be $(\alpha,{\mathcal{B}}_{I},\tau)$-robust if $\nu$ can be written as the sum of two Borel measures $\nu=\nu^{\prime}+\nu^{\prime\prime}$ such that $\nu^{\prime\prime}(X)\leq\tau$, and $\nu^{\prime}$ satisfies
(6.1) $\nu^{\prime}\{\operatorname{inj}<\sup I\}=0,$
as well as for all $\rho\in I$, $y\in X$,
(6.2) $\nu^{\prime}(B_{\rho}y)\leq\rho^{\alpha\dim X}$
If $I$ is a singleton $I=\{\rho\}$, we simply say that $\nu$ is $(\alpha,{\mathcal{B}}_{\rho},\tau)$-robust.
We aim to show the following.
Proposition 6.2 (High dimension).
Let $\kappa\in(0,1/10)$. For $\eta,\rho\lll_{\kappa}1$ and for all $n\ggg_{\kappa}\lvert\log\rho\rvert+\lvert\log\operatorname{inj}(x)\rvert$, the measure $\mu^{*n}*\delta_{x}$ is $(1-\kappa,B_{\rho},\rho^{\eta})$-robust.
6.1. Non-concentration inequalities
The strategy to prove Proposition˜6.2 is to show that convolution by $\mu$ (or a suitable power $\mu^{*n}$) improves the dimensional properties of any given Frostman measure $\nu$ on $X$. Iterating this phenomenon allows to reach high dimension. The dimensional increment property for random walks is rooted in the following key observation:
$\mu^{*n}*\nu(B_{\rho}x)=\int_{G}\nu(g^{-1}B_{\rho}x)\,\,\mathrm{d}\mu^{*n}(g)$
and $g^{-1}B_{\rho}x$ can be seen, in some exponential chart, as a Euclidean box $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho}$, varying randomly with $g\sim\mu^{*n}$. Provided this random box satisfies suitable non-concentration properties, we can then derive a small dimensional increment via a multislicing theorem (which itself boils down to the sum product phenomenon).
The required non-concentration concerns the partial flag carrying the box. Let us see what it is in our setting. Consider the weight spaces decomposition $\mathfrak{g}=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}\oplus\mathfrak{g}_{+}$ for $A^{\prime}$. More precisely, $\mathfrak{g}_{+},\mathfrak{g}_{-}$ are respectively the Lie algebras of $U$ and $U^{-}$, where $U^{-}$ denotes the transpose of $U$, and $\mathfrak{g}_{0}$ is their orthogonal complement in $\mathfrak{g}$. Recalling $g=k_{g}^{-1}a(\mathtt{r}_{g}^{-1})u(\mathtt{b}_{g})$ and the norm on $\mathfrak{g}$ is $\operatorname{Ad}(K^{\prime})$-invariant, the box $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho}$ can be written
[…]
(6.3) $\overset{L_{g}}{\simeq}B^{\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-}}_{\mathtt{r}^{-1}_{g}\rho}+B^{\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0}}_{\rho}+B^{\mathfrak{g}}_{\mathtt{r}_{g}\rho}$
where $\mathfrak{g}_{\leq 0}:=\mathfrak{g}_{-}\oplus\mathfrak{g}_{0}$, $L_{g}=O((1+\|\mathtt{b}_{g}\|)^{d+1})$, and the notation $A\overset{L}{\simeq}B$ means that $A$ can be covered by less than $L$ additive translates of $B$, and conversely. Note the norm $\|\mathtt{b}_{g}\|$ is controlled via Lemma˜3.5 (i). We are left to examine the non-concentration properties of the partial flag given by $V_{1}(g):=\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-}$ and $V_{2}(g):=\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0}$ as $g$ varies with law $\mu^{*n}$.
In [7] about the case $d=1$ (as well as in [6]), a similar approach is exploited, but the non-concentration at disposal therein is very strong, namely: for $i=1,2$, any subspace $W\subseteq\mathfrak{g}$ with $\dim V_{i}+\dim W=\dim\mathfrak{g}$, for $\mu^{*n}$-many $g$, we have $V_{i}(g)\cap W=\{0\}$ with a large angle between $V_{i}(g)$ and $W$. This non-concentration requirement is natural, as it is the hypothesis of the projection theorems à la Bourgain which are at the heart of the multislicing estimates from [6]. Unfortunately, such property fails for $d\neq 1$, as we see in the next lemma.
Lemma 6.3 (Obstacle).
Assume $d\geq 2$. Let $W=\{\,M\in\mathfrak{g}\,:\,Me_{1}=0\,\}$ be the subspace of matrices in $\operatorname{M}_{d}(\mathbb{R})$ with zero trace and null first column. Then $\operatorname{codim}W=d+1>\dim\mathfrak{g}_{-}$ but for every $g\in G$, we have
$\operatorname{Ad}(g)\mathfrak{g}_{-}\cap W\neq\{0\}.$
Proof.
Observe that
$\mathfrak{g}_{-}=\oplus_{j=1}^{d}\mathbb{R}E_{d+1,j}=\begin{pmatrix}0&0&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&0\\ *&*&\cdots\,*&0\end{pmatrix}$.
Therefore, $\operatorname{Ad}(g)\mathfrak{g}_{-}$ corresponds to the collection of endomorphisms of $\mathbb{R}^{d+1}$ which are $2$-step nilpotent and with image in $g\mathbb{R}e_{d+1}$. Consider $m,m^{\prime}\in\operatorname{Ad}(g)\mathfrak{g}_{-}$ non colinear. As $me_{1},m^{\prime}e_{1}$ are colinear, there must exist $(t,t^{\prime})\in\mathbb{R}^{2}\smallsetminus\{0\}$ with $(tm+t^{\prime}m^{\prime})e_{1}=0$, whence $tm+t^{\prime}m^{\prime}\in W$. This justifies that any $2$-dimensional subspace of $\operatorname{Ad}(g)\mathfrak{g}_{-}$ intersects $W$, whence $\dim(\operatorname{Ad}(g)\mathfrak{g}_{-}\cap W)\geq d-1$. ∎
In this section, we consider an arbitrary $d\geq 1$ and show that the random subspaces $V_{1}(g),V_{2}(g)$ where $g\sim\mu^{*n}$ still satisfy a weak form of non-concentration. It is presented as Proposition˜6.4 below. We will explain afterward how this can be utilized to perform the dimensional bootstrap.
Given subspaces $F_{1},\dots,F_{k}\in\operatorname{Gr}(\mathfrak{g})$, we write
$\|F_{1}\wedge\dots\wedge F_{k}\|:=\|v_{1}\wedge\dots\wedge v_{k}\|$
where $v_{i}\in\bigwedge^{*}\mathfrak{g}$ is a unit vector spanning the line $\bigwedge^{\dim F_{i}}F_{i}.$
Proposition 6.4 (Mild non-concentration).
Let $W\in\operatorname{Gr}(\mathfrak{g},d)$. Then for $n\geq 1$, $r\geq e^{-n}$,
$(\mu^{*n})^{\otimes d+1}\left\{(g_{i})_{i=1}^{d+1}\,:\,\|V_{1}(g_{1})\wedge\dots\wedge V_{1}(g_{d+1})\wedge W\|\leq r\right\}\leq Cr^{c}$
where $C,c>0$ are constants depending on $\mu$ only.
Observing $\dim\mathfrak{g}=d(d+2)$, Proposition˜6.4 means that for most parameters $(g_{i})_{i=1}^{d+1}$ selected by $(\mu^{*n})^{\otimes d+1}$, we have $\mathfrak{g}=\oplus_{i}V_{1}(g_{i})\bigoplus W$, and each subspace makes a rather large angle with the complementary sum.
We may derive a similar non-concentration property for $V_{2}(g)^{\perp}$ as $g\sim\mu^{*n}$.
Corollary 6.5.
Let $W\in\operatorname{Gr}(\mathfrak{g},d)$. Then for $n\geq 1$, $r\geq e^{-n}$,
$(\mu^{*n})^{\otimes d+1}\left\{(g_{i})_{i=1}^{d+1}\,:\,\|V_{2}(g_{1})^{\perp}\wedge\dots\wedge V_{2}(g_{d+1})^{\perp}\wedge W\|\leq r\right\}\leq Cr^{c}$
where $C,c>0$ are constants depending on $\mu$ only.
Proof of Corollary˜6.5.
Recall that the Lie algebra $\mathfrak{g}=\mathfrak{sl}_{d+1}(\mathbb{R})$ is equipped with the scalar product given by $\langle A,B\rangle=\operatorname{tr}(A^{T}B)$ where $A^{T}$ denotes the transpose of the matrix $A$. Using $\operatorname{tr}(AB)=\operatorname{tr}(BA)$, it is direct to check for every $g\in G$, the adjoint of $\operatorname{Ad}(g)\in\operatorname{End}(\mathfrak{g})$ for this Euclidean structure is given by $\operatorname{Ad}(g)^{*}=\operatorname{Ad}(g^{T})$. Moreover, the eigenspaces $\mathfrak{g}_{-}$, $\mathfrak{g}_{0}$ and $\mathfrak{g}_{+}$ are mutually orthogonal. It follows that for $g\in P^{\prime}$, we have
$V_{2}(g)^{\perp}=(\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0})^{\perp}=\operatorname{Ad}(u(\mathtt{b}_{g})^{T})\mathfrak{g}_{+}=\left(\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-}\right)^{T}.$
As the map $\mathfrak{g}\mapsto\mathfrak{g},A\mapsto A^{T}$ is an isometry, the claim follows from Proposition˜6.4. ∎
We now focus on establishing Proposition˜6.4. We first reduce to a purely geometric version of that result.
Proposition 6.6 (Geometric reduction).
Let $W\in\operatorname{Gr}(\mathfrak{g},d)$. Then there exists $(u_{i})_{i=1}^{d+1}\in U^{d+1}$ such that
$\operatorname{Ad}(u_{1})\mathfrak{g}_{-}\oplus\operatorname{Ad}(u_{2})\mathfrak{g}_{-}\oplus\dots\oplus\operatorname{Ad}(u_{d+1})\mathfrak{g}_{-}\oplus W=\mathfrak{g}$
Proposition˜6.6 is geometric in the sense that no random variable is involved. It turns out to be equivalent to Proposition˜6.4.
Proof that Proposition˜6.4 $\iff$ Proposition˜$6.$
The direct implication is clear, therefore we assume Proposition˜6.6 and check Proposition˜6.4. We write $P(\boldsymbol{s}):=\operatorname{Ad}(u(-\boldsymbol{s}))$ for conciseness. Observe that the angle function $(\mathbb{R}^{d})^{d}\rightarrow\bigwedge^{\dim\mathfrak{g}}\mathfrak{g}\simeq\mathbb{R},(\boldsymbol{s}_{i})_{i}\mapsto\lVert P(\boldsymbol{s}_{1})\mathfrak{g}_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})\mathfrak{g}_{-}\wedge W\rVert$ is Lipschitz continuous. In view of Lemma˜2.1, it suffices to show the existence of constants $C,c>0$ depending only on $\sigma$ such that for every $r>0$,
$\sigma^{\otimes d+1}\left\{(\boldsymbol{s}_{i})_{i=1}^{d+1}\,:\,\|P(\boldsymbol{s}_{1})\mathfrak{g}_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})\mathfrak{g}_{-}\wedge W\|\leq r\right\}\leq Cr^{c}.$
Let $v_{-},w\in\bigwedge^{*}\mathfrak{g}$ be unit vectors spanning respectively the lines $\bigwedge^{\dim\mathfrak{g}_{-}}\mathfrak{g}_{-}$ and $\bigwedge^{\dim W}W$. Note that
(6.4) $\|P(\boldsymbol{s}_{1})\mathfrak{g}_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})\mathfrak{g}_{-}\wedge W\|=\frac{\|P(\boldsymbol{s}_{1})v_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})v_{-}\wedge w\|}{\|P(\boldsymbol{s}_{1})v_{-}\|\dotsm\|P(\boldsymbol{s}_{d+1})v_{-}\|}.$
As the map $\boldsymbol{s}\mapsto P(\boldsymbol{s})$ is polynomial, and $\sigma$ has finite moment of positive order (Lemma˜3.1), we have for some $\gamma=\gamma(\sigma)>0$,
(6.5) $\sigma^{\otimes d+1}\{\|P(\boldsymbol{s}_{1})v_{-}\|\dotsm\|P(\boldsymbol{s}_{d+1})v_{-}\|\geq r^{-1/2}\}\ll_{\sigma}r^{\gamma}.$
On the other hand, Proposition˜6.6 guarantees that the polynomial map $(\boldsymbol{s}_{i})_{i}\mapsto P(\boldsymbol{s}_{1})v_{-}\wedge\dots\wedge P(\boldsymbol{s}_{d+1})v_{-}\wedge w$ is non-zero. As it depends continuously on $w$ and $\operatorname{Gr}(\mathfrak{g},d)$ is compact, it must have the supremum norm on the coefficients bounded below by a constant $c_{d}>0$ depending only on $d$. Combined with Corollary˜3.4, this yields
(6.6) $\sigma^{\otimes d+1}\{\|P(\boldsymbol{s}_{1})v_{-}\wedge\dotsm\wedge P(\boldsymbol{s}_{d+1})v_{-}\wedge w\|<r^{1/2}\}\ll_{\sigma}r^{\gamma}$
up to taking $\gamma$ smaller. Proposition˜6.4 follows from the combination of (6.4), (6.5), (6.6). ∎
We further reduce to the case where the subspace $W$ is invariant under a Borel subgroup of $G$. We denote by $B$ the upper triangular subgroup of $G$.
Proposition 6.7 (Borel-invariant reduction).
Let $W\in\operatorname{Gr}(\mathfrak{g},d)$ be a subspace which is $\operatorname{Ad}(B)$-invariant. Then there exists $(g_{i})_{i=1}^{d+1}\in G^{d+1}$ such that
$\operatorname{Ad}(g_{1})\mathfrak{g}_{-}\oplus\operatorname{Ad}(g_{2})\mathfrak{g}_{-}\oplus\dots\oplus\operatorname{Ad}(g_{d+1})\mathfrak{g}_{-}\oplus W=\mathfrak{g}.$
Let us check that that Proposition˜6.6 and Proposition˜6.7 are equivalent.
Proof that Proposition˜6.6 $\iff$ Proposition˜6.7.
The direct implication is clear. We establish the converse. Assume by contradiction that Proposition˜6.6 fails for some $W\in\operatorname{Gr}(\mathfrak{g},d)$. Write $Z_{G}(A^{\prime})$ the centralizer of $A^{\prime}$ in $G$. Noting that $\mathfrak{g}_{-}$ is $Z_{G}(A^{\prime})U^{-}$-invariant, we obtain for every $(g_{i})_{i}\in(UZ_{G}(A^{\prime})U^{-})^{d+1}$ that
(6.7) $\operatorname{Ad}(g_{1})\mathfrak{g}_{-}+\operatorname{Ad}(g_{2})\mathfrak{g}_{-}+\dots+\operatorname{Ad}(g_{d+1})\mathfrak{g}_{-}+W\neq\mathfrak{g}.$
This is a Zariski-closed condition in the variable $(g_{i})_{i=1}^{d+1}$. By looking at Lie algebras, we see that $UZ_{G}(A^{\prime})\supseteq B$, so by Bruhat’s decomposition, $UZ_{G}(A^{\prime})U^{-}$ is Zariski-dense in $G$. It follows that (6.7) holds for all $(g_{i})_{i=1}^{d+1}\in G^{d+1}$. Applying another $\operatorname{Ad}(g)$ on both side we see the set
[…]
is preserved under the action of $\operatorname{Ad}(G)$, whence of $\operatorname{Ad}(B)$. It is moreover Zariski-closed. More precisely, it is the set of $\mathbb{R}$-points of a complete $\mathbb{R}$-variety. On the other hand, $B$ is the set of $\mathbb{R}$-points of a $\mathbb{R}$-split connected solvable linear algebraic group which acts $\mathbb{R}$-morphically on this variety. Thus, by a version of the Borel fixed point theorem ([11, Proposition 15.2]), the above set contains a fixed point for $\operatorname{Ad}(B)$. This contradicts Proposition˜6.7, thus finishing the proof of the converse implication. ∎
The advantage of reducing to Proposition˜6.7 is that it constrains $W$ to belong to a finite explicit family of subspaces.
Lemma 6.8.
Let $W\subseteq\mathfrak{g}$ be a subspace of dimension at most $d$. Then $W$ is $\operatorname{Ad}(B)$-invariant if and only if we can write
$W=\oplus_{(i,j)\in S}\mathbb{R}E_{i,j}$
where $S\subseteq\{(i,j)\,:\,1\leq i<j\leq d+1\}$ and $S$ is stable by the operations $(i,j)\mapsto(i-1,j)$ and $(i,j)\mapsto(i,j+1)$ (provided $i\geq 2$ and $j\leq d$ respectively).
In words, $W$ must be a sum of elementary subspaces that are strictly above the diagonal, and stable by moving upward or to the right in the matrix representation.
Proof.
Recall $\mathfrak{g}=\mathfrak{sl}_{d+1}$. Write $\mathfrak{b}$ the Lie algebra of $B$, i.e. the subspace of upper triangular matrices in $\mathfrak{g}$. Note that $W$ is $\operatorname{Ad}(B)$-invariant if and only if it is $\operatorname{ad}(\mathfrak{b})$-invariant. Observe the relation $E_{i,j},E_{k,l}]={\mathbbm{1}}_{j=k}E_{i,l}-{\mathbbm{1}}_{i=l}E_{k,j}$ for all $i,j,k,l\in\{1,\dots,d+1\}$. In particular, for $i<j$ and $k<l$, the matrix $E_{i,j},E_{k,l}]$ is either 0 or up to a sign an elementary matrix located either to the right or above $E_{i,j}$. This justifies the “if” direction in Lemma˜6.8.
We now assume $W$ to be $\operatorname{Ad}(B)$-invariant and establish the announced decomposition. By invariance under diagonal matrices, $W$ must be of the form $W=E\bigoplus\oplus_{(i,j)\in S}\mathbb{R}E_{i,j}$ where $E$ is a subspace of diagonal matrices, and $S$ does not intersect the diagonal. If $E\neq\{0\}$, then by $\operatorname{ad}(\mathfrak{b})$-invariance, $W$ must contain a line $\mathbb{R}E_{i,i+1}$ for some $i\in\{1,\dots,d\}$. But the $\operatorname{ad}(\mathfrak{b})$-invariant subspace spanned by such a line has dimension at least $d$, which is absurd because $\dim W\leq d$. Hence $E=\{0\}$. Noting that for every $i>j$, the $\operatorname{ad}(\mathfrak{b})$-invariant subspace spanned by $\mathbb{R}E_{i,j}$ intersects the diagonal subspace, we further deduce $S\subseteq\{(i,j)\,:\,1\leq i<j\leq d+1\}$. The final claim on $S$ follows from $\operatorname{ad}(\mathfrak{b})$-invariance and the bracket relation exhibited in the first paragraph. ∎
We are finally able to show Proposition˜6.7, thus completing the proofs of Propositions 6.4, 6.6.
Proof of Proposition˜6.7.
We shall prove this proposition by induction on $d$.
Base case $d=1$. Here $\mathfrak{g}=\mathfrak{sl}_{2}(\mathbb{R})$ and $\mathfrak{g}_{-}=\mathbb{R}E_{21}$. By assumption,
$W=\mathbb{R}E_{12}.$
Take $g_{1}=\operatorname{Id},g_{2}=\operatorname{Id}+E_{12}\in\operatorname{SL}_{2}(\mathbb{R})$. By direct computation, we can verify that
$\operatorname{Ad}(g_{1})\mathfrak{g}_{-}\oplus\operatorname{Ad}(g_{2})\mathfrak{g}_{-}\oplus W=\mathfrak{g}.$
Induction step. Let $d\geq 2$ be an integer. We assume the proposition has been proved for $\operatorname{SL}_{d}(\mathbb{R})$ and establish it for $\operatorname{SL}_{d+1}(\mathbb{R})$. Throughout the proof, we write $M_{d+1}$ the space of all $d+1$ by $d+1$ real matrices. We keep the notations $G,\mathfrak{g},\mathfrak{g}_{-}$ related to $\operatorname{SL}_{d+1}(\mathbb{R})$. We write $G^{\prime}=\operatorname{SL}_{d}(\mathbb{R})$, which we view as a subgroup of $G$ by embedding it in the lower-right corner (and imposing $1$ on the first diagonal entry). Accordingly, we define $\mathfrak{g}^{\prime}$ (resp. $\mathfrak{g}_{-}^{\prime}$) to be the intersection of $\mathfrak{g}$ (resp. $\mathfrak{g}_{-}$) with the lower-right $d$ by $d$ block of $M_{d+1}$. In particular,
$\mathfrak{g}_{-}^{\prime}=\mathbb{R}E_{d+1,2}\oplus\cdots\oplus\mathbb{R}E_{d+1,d}.$
We denote by $\operatorname{Proj}_{\mathfrak{g}^{\prime}}:M_{d+1}\to M_{d+1}$ the projection onto the lower right $d$ by $d$ block, and by $\operatorname{Proj}_{R_{1}}:M_{d+1}\to M_{d+1}$ the projection onto the the subspace of matrices with nonzero entries only on the first row.
Let $W\subseteq\mathfrak{g}$ be a $d$-dimensional linear subspace that is $\operatorname{Ad}(B)$-invariant. To make use of the induction hypothesis, our strategy is to choose a suitable element $g_{0}\in G$ such that $\operatorname{Ad}(g_{0})\mathfrak{g}_{-}\oplus W$ fills up the first row of $\mathfrak{g}$. This will enable us to work on $\mathfrak{g}^{\prime}$ and apply the induction hypothesis.
To begin with, we let $k=\dim\operatorname{Proj}_{R_{1}}(W)\geq 1$. We choose
$g_{0}=(\operatorname{Id}+E_{d+1,1})\cdot\omega_{1,d+1}.$
where $\omega_{1,d+1}$ is an element in the standard Weyl group of $G$ satisfying that left multiplication by $\omega_{1,d+1}$ exchanges the first and $(d+1)$-th row. Then by direct computation, we have
$\operatorname{Ad}(g_{0})\mathfrak{g}_{-}=\left\{\begin{pmatrix}a_{1}&a_{2}&\cdots&a_{d}&-a_{1}\\ 0&0&\cdots&0&0\\ \vdots&\vdots&\ddots&\vdots&\vdots\\ 0&0&\cdots&0&0\\ a_{1}&a_{2}&\cdots&a_{d}&-a_{1}\end{pmatrix}:a_{1},\cdots,a_{d}\in\mathbb{R}\right\}.$
Decompose $\mathfrak{g}_{-}=V_{0}\oplus V_{1}$, where $V_{0},V_{1}$ are linear subspaces defined by
$V_{0}=\mathbb{R}E_{d+1,1}\oplus\mathbb{R}E_{d+1,2}\oplus\cdots\oplus\mathbb{R}E_{d+1,d+1-k};$
$V_{1}=\mathbb{R}E_{d+1,d+2-k}\oplus\cdots\oplus\mathbb{R}E_{d+1,d}.$
Using Lemma˜6.8, we also decompose $W=\operatorname{Proj}_{R_{1}}(W)\oplus\mathrm{Proj_{\mathfrak{g}^{\prime}}}(W)$. Then we have
(6.8) $\operatorname{Ad}(g_{0})\mathfrak{g}_{-}\oplus W=\operatorname{Ad}(g_{0})V_{0}\oplus\mathrm{Proj_{R_{1}}(W)}\oplus V_{1}\oplus\mathrm{Proj_{\mathfrak{g}^{\prime}}(W)}.$
Here we used that $V_{1}$ and $\operatorname{Ad}(g_{0})V_{1}$ coincide modulo $W$. Let
$W^{\prime}=V_{1}\oplus\mathrm{Proj_{\mathfrak{g}^{\prime}}(W)}.$
Note that $W^{\prime}\subseteq\mathfrak{g}^{\prime}$ and
$\dim W^{\prime}=k-1+d-k=d-1.$
Hence, we can apply the induction hypothesis (more precisely its equivalent version from Proposition˜6.6) to the pair $(G^{\prime},W^{\prime})$ to obtain $g^{\prime}_{1},\cdots,g^{\prime}_{d}\in G^{\prime}$ such that
(6.9) $\operatorname{Ad}(g^{\prime}_{1})\mathfrak{g}^{\prime}_{-}\oplus\operatorname{Ad}(g^{\prime}_{2})\mathfrak{g}^{\prime}_{-}\oplus\cdots\oplus\operatorname{Ad}(g^{\prime}_{d})\mathfrak{g}^{\prime}_{-}\oplus W^{\prime}=\mathfrak{g}^{\prime}.$
Let $C\subseteq\mathfrak{g}$ be the linear subspace defined by
$C=\mathbb{R}E_{2,1}\oplus\mathbb{R}E_{3,1}\oplus\cdots\oplus\mathbb{R}E_{d+1,1}.$
Observe that $C$ is $\operatorname{Ad}(G^{\prime})$-invariant, and the adjoint representation of $G^{\prime}$ on $C$ is isomorphic to the standard representation $\mathbb{R}^{d}$ of $G^{\prime}=\operatorname{SL}_{d}(\mathbb{R})$. Therefore, we may find $g_{1}^{\prime\prime},g_{2}^{\prime\prime},\cdots,g_{d}^{\prime\prime}\in G^{\prime}$ such that
(6.10) $\operatorname{Ad}(g_{1}^{\prime\prime})\mathbb{R}E_{d+1,1}\oplus\operatorname{Ad}(g_{2}^{\prime\prime})\mathbb{R}E_{d+1,1}\oplus\cdots\oplus\operatorname{Ad}(g_{d}^{\prime\prime})\mathbb{R}E_{d+1,1}=C.$
Observe that the collection of elements $(g^{\prime}_{i})_{i=1}^{d}$ and $(g^{\prime\prime}_{i})_{i=1}^{d}$ satisfying respectively (6.9) and (6.10) are (non-empty) Zariski-open subsets of $G^{\prime d}$. By irreducibility of $G^{\prime d}$ for the Zariski topology, they are dense, whence must intersect. This allows to choose $(g^{\prime}_{1},\cdots,g^{\prime}_{d})=(g^{\prime\prime}_{1},\cdots,g^{\prime\prime}_{d})$ in (6.9) and (6.10). As
$\mathfrak{g}_{-}=\mathfrak{g}^{\prime}_{-}\oplus\mathbb{R}E_{d+1,1},$
we obtain
(6.11) $\operatorname{Ad}(g^{\prime}_{1})\mathfrak{g}_{-}\oplus\operatorname{Ad}(g^{\prime}_{2})\mathfrak{g}_{-}\oplus\cdots\oplus\operatorname{Ad}(g^{\prime}_{d})\mathfrak{g}_{-}\oplus W^{\prime}=\mathfrak{g}^{\prime}\oplus C.$
On the other hand, observing that the restriction of $\operatorname{Proj}_{R_{1}}$ to $\operatorname{Ad}(g_{0})V_{0}\oplus\mathrm{Proj_{R_{1}}}(W)$ is injective while its restriction to $\mathfrak{g}^{\prime}\oplus C$ vanishes, we have $(\operatorname{Ad}(g_{0})V_{0}\oplus\mathrm{Proj_{R_{1}}}(W))\cap(\mathfrak{g}^{\prime}\oplus C)=\{0\}$, or equivalently
(6.12) $\operatorname{Ad}(g_{0})V_{0}\oplus\mathrm{Proj_{R_{1}}}(W)\oplus\mathfrak{g}^{\prime}\oplus C=\mathfrak{g}$
because dimensions match. Combining (6.8), (6.11), (6.12), we obtain
$\operatorname{Ad}(g_{0})\mathfrak{g}_{-}\oplus\operatorname{Ad}(g^{\prime}_{1})\mathfrak{g}_{-}\oplus\operatorname{Ad}(g^{\prime}_{2})\mathfrak{g}_{-}\oplus\cdots\oplus\operatorname{Ad}(g^{\prime}_{d})\mathfrak{g}_{-}\oplus W=\mathfrak{g}.$
This validates the induction step and completes the proof. ∎
6.2. Linear multislicing
It remains to see how the non-concentration property established for the subspace $\operatorname{Ad}(u(-\mathtt{b}_{g})\mathfrak{g}^{-})$ in the previous subsection can be exploited to obtain a dimensional gain. In this subsection, we study this question in an abstract linear setting. We place ourselves in $\mathbb{R}^{D}$ where $D\geq 2$. We encapsulate the non-concentration property via the following definition.
Definition 6.9.
Let $k\in\llbracket 1,D-1\rrbracket$, let $C,c,\rho>0$. Let $\Xi$ be a probability measure on $\operatorname{Gr}(\mathbb{R}^{D},k)$. We say $\Xi$ satisfies the mild non-concentration property $\operatorname{(MNC)}$ with parameters $(\rho,C,c)$ if there exist integers $q,m\in\mathbb{N}$ such that $D=qk+m$ and for every $W\in\operatorname{Gr}(\mathbb{R}^{D},m)$, $r\geq\rho$,
$\Xi^{\otimes q}\left\{(F_{i})\,:\,\|F_{1}\wedge\dots\wedge F_{q}\wedge W\|\leq r\right\}\leq Cr^{c}.$
We also say $\Xi$ satisfies $\operatorname{(MNC)}^{\perp}$ with parameters $(\rho,C,c)$ if its image under $F\mapsto F^{\perp}$ satisfies $\operatorname{(MNC)}$ with parameters $(\rho,C,c)$.
Our aim is to show that $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ allow for a supercritical multislicing estimate, see Proposition˜6.14. For that, we first present a submodular inequality for covering numbers (Lemma˜6.10) and use it to connect $\operatorname{(MNC)}$ and $\operatorname{(MNC)}^{\perp}$ with the properties of the individual projectors $\pi_{F_{i}}$ (Lemma˜6.11). We then deduce that a random subspace $F$ whose law $\Xi$ has the property $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ must enjoy both supercritical and subcritical projection theorems (Lemmas 6.12, 6.13). Those estimates refine that of Bourgain [13] and He [26] which were established under the stronger non-concentration condition that $\Xi$ satisfies $\operatorname{(MNC)}$ with $q=1$. From there, we use [6] to combine our projection theorems into the multislicing estimate Proposition˜6.14
We now introduce the submodular inequality for covering numbers that we need. Let $\mathcal{P}$} and $\mathcal{Q}$} denote partitions of $\mathbb{R}^{D}$, let $A$ be a subset of $\mathbb{R}^{D}$. We write ${\mathcal{P}}(A)$ the set of cells of $\mathcal{P}$} that meet $A$, that is,
${\mathcal{P}}(A):=\{\,P\in{\mathcal{P}}:P\cap A\neq\varnothing\,\},$
and set ${\mathcal{N}}_{\mathcal{P}}(A)$ the cardinality of ${\mathcal{P}}(A)$. We say $\mathcal{Q}$} roughly refines $\mathcal{P}$} with parameter $L\geq 1$, and write ${\mathcal{P}}\overset{L}{\prec}{\mathcal{Q}}$, if
$\max_{Q\in{\mathcal{Q}}}\,{\mathcal{N}}_{\mathcal{P}}(Q)\leq L.$
We also use the notation ${\mathcal{P}}\overset{L}{\simeq}{\mathcal{Q}}$ to say that both ${\mathcal{P}}\overset{L}{\prec}{\mathcal{Q}}$ and ${\mathcal{Q}}\overset{L}{\prec}{\mathcal{P}}$ hold. Finally, we denote by ${\mathcal{P}}\vee{\mathcal{Q}}$ the partition obtained by taking the intersections of $\mathcal{P}$}-cells and $\mathcal{Q}$}-cells.
Lemma 6.10 (Submodular inequality).
Let ${\mathcal{P}},{\mathcal{Q}},{\mathcal{R}},{\mathcal{S}}$ be partitions of $\mathbb{R}^{D}$, and $A$ a subset of $\mathbb{R}^{D}$. Let $L\geq 1$. Assume that ${\mathcal{R}}\overset{L}{\simeq}{\mathcal{P}}\vee{\mathcal{Q}}$, and ${\mathcal{S}}\overset{L}{\prec}{\mathcal{P}}$, ${\mathcal{S}}\overset{L}{\prec}{\mathcal{Q}}$. Then for every $c>0$, there is a subset $A^{\prime}\subseteq A$ such that ${\mathcal{N}}_{{\mathcal{R}}}(A^{\prime})\geq\frac{1-c}{L^{3}}{\mathcal{N}}_{{\mathcal{R}}}(A)$ and
(6.13) ${\mathcal{N}}_{\mathcal{P}}(A){\mathcal{N}}_{\mathcal{Q}}(A)\geq\frac{c^{2}}{4L^{3}}{\mathcal{N}}_{\mathcal{R}}(A){\mathcal{N}}_{{\mathcal{S}}}(A^{\prime}).$
In the case where $L=1$, the result is due to [6, Lemma 2.6]. We deduce the refinement presented in Lemma˜6.10. Such an upgrade is convenient to deal with situations where partitions ${\mathcal{P}},{\mathcal{Q}},{\mathcal{R}},{\mathcal{S}}$ do not exactly fit together.
Proof.
We start with a few general observations on the relation $\overset{L}{\prec}$. Note it is transitive in the sense that ${\mathcal{P}}\overset{L}{\prec}{\mathcal{P}}^{\prime}$ and ${\mathcal{P}}^{\prime}\overset{L^{\prime}}{\prec}{\mathcal{P}}^{\prime\prime}$ implies ${\mathcal{P}}\overset{LL^{\prime}}{\prec}{\mathcal{P}}^{\prime\prime}$. It is also compatible with taking common refinements, that is, ${\mathcal{P}}\overset{L}{\prec}{\mathcal{Q}}$ and ${\mathcal{P}}^{\prime}\overset{L^{\prime}}{\prec}{\mathcal{Q}}^{\prime}$ implies ${\mathcal{P}}\vee{\mathcal{P}}^{\prime}\overset{LL^{\prime}}{\prec}{\mathcal{Q}}\vee{\mathcal{Q}}^{\prime}$. Finally, observe that ${\mathcal{P}}\overset{L}{\prec}{\mathcal{Q}}$ implies ${\mathcal{N}}_{\mathcal{P}}(A)\leq L{\mathcal{N}}_{\mathcal{Q}}(A)$ for any subset $A$.
Now consider ${\mathcal{P}}_{0}={\mathcal{P}}\vee{\mathcal{S}}$, ${\mathcal{Q}}_{0}={\mathcal{Q}}\vee{\mathcal{S}}$ and ${\mathcal{R}}_{0}={\mathcal{P}}\vee{\mathcal{Q}}\vee{\mathcal{S}}$. By [6, Lemma 2.6] applied to ${\mathcal{P}}_{0}$, ${\mathcal{Q}}_{0}$, ${\mathcal{R}}_{0}$ and $\mathcal{S}$}, there is a subset $A^{\prime}\subseteq A$ such that ${\mathcal{N}}_{{\mathcal{R}}_{0}}(A^{\prime})\geq(1-c){\mathcal{N}}_{{\mathcal{R}}_{0}}(A)$ and
(6.14) ${\mathcal{N}}_{{\mathcal{P}}_{0}}(A){\mathcal{N}}_{{\mathcal{Q}}_{0}}(A)\geq\frac{c^{2}}{4}{\mathcal{N}}_{{\mathcal{R}}_{0}}(A){\mathcal{N}}_{{\mathcal{S}}}(A^{\prime}).$
Using the properties of $\overset{L}{\prec}$ recalled above, we derive from the assumptions that ${\mathcal{P}}_{0}\overset{L}{\prec}{\mathcal{P}}$ and ${\mathcal{Q}}_{0}\overset{L}{\prec}{\mathcal{Q}}$, as well as ${\mathcal{R}}\overset{L}{\prec}{\mathcal{R}}_{0}$ and ${\mathcal{R}}_{0}\overset{L^{2}}{\prec}{\mathcal{R}}$. The inequality (6.13) then follows from (6.14). ∎
In the next lemma, we consider a family of projectors of $\mathbb{R}^{D}$ whose images (resp. kernels) are in direct sum with controlled angle. Given a set $A\subseteq\mathbb{R}^{D}$, we relate the product of covering numbers of the projections of $A$ with the covering number of the projection of $A$ onto (resp. parallel to) the sum of the images (resp. kernels). Given $F\in\operatorname{Gr}(\mathbb{R}^{d})$, we let $\pi_{F},\pi_{||F}:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ denote respectively the orthogonal projectors of image or kernel $F$. For $\rho>0$, we denote by ${\mathcal{N}}_{\rho}(A)$ the least number of open balls of radius $\rho$ needed to cover $A$.
Lemma 6.11.
Let $(F_{i})_{i=1,\dots,q}$ be a (non-necessarily generating) collection of subspaces in $\mathbb{R}^{D}$. Assume for some $r\in(0,1/2)$ that
(6.15) $\lVert F_{1}\wedge\dots\wedge F_{q}\rVert\geq r.$
Then for any set $A\subseteq\mathbb{R}^{D}$, one has
(6.16) $\prod_{i=1}^{q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}A)\geq r^{O_{D}(1)}{\mathcal{N}}_{\rho}(\pi_{\oplus_{i}F_{i}}A),$
and
(6.17) $\prod_{i=1}^{q}{\mathcal{N}}_{\rho}(\pi_{||F_{i}}A)\geq r^{O_{D}(1)}{\mathcal{N}}_{\rho}(A)^{q-1}{\mathcal{N}}_{\rho}(\pi_{||\oplus_{i}F_{i}}A^{\prime})$
for some subset $A^{\prime}\subseteq A$ satisfying ${\mathcal{N}}_{\rho}(A^{\prime})\geq r^{O_{D}(1)}{\mathcal{N}}_{\rho}(A)$.
Proof.
The first inequality (6.16) is a simple counting, see e.g. [26, Lemma 15]. We focus on proving (6.17).
By induction on $q$ together with the observation that
$\lVert(F_{1}\oplus\dots\oplus F_{i})\wedge F_{i+1}\rVert\geq\lVert F_{1}\wedge\dotsc\wedge F_{q}\rVert\geq r,$
the proof of (6.17) reduces to the case $q=2$.
Let ${\mathcal{D}}_{\rho}$ denote the partition corresponding to the tiling of $\mathbb{R}^{D}$ by the cube $0,\rho)^{D}$ and its $\rho\mathbb{Z}^{D}$-translates. Consider ${\mathcal{P}}=\pi_{||F_{1}}^{-1}({\mathcal{D}}_{\rho})$, ${\mathcal{Q}}=\pi_{||F_{2}}^{-1}({\mathcal{D}}_{\rho})$, ${\mathcal{R}}={\mathcal{D}}_{\rho}$ and ${\mathcal{S}}=\pi_{||F_{1}\oplus F_{2}}^{-1}({\mathcal{D}}_{\rho})$, so that ${\mathcal{N}}_{\mathcal{P}}(A)\simeq_{D}{\mathcal{N}}_{\rho}(\pi_{||F_{1}}A)$, ${\mathcal{N}}_{\mathcal{Q}}(A)\simeq_{D}{\mathcal{N}}_{\rho}(\pi_{||F_{2}}A)$, ${\mathcal{N}}_{\mathcal{R}}(A)\simeq_{D}{\mathcal{N}}_{\rho}(A)$ and ${\mathcal{N}}_{\mathcal{S}}(A)\simeq_{D}{\mathcal{N}}_{\rho}(\pi_{||F_{1}\oplus F_{2}}A)$.
From $\lVert F_{1}\wedge F_{2}\rVert\geq r$ we know that ${\mathcal{R}}\overset{r^{-O_{D}(1)}}{\prec}{\mathcal{P}}\vee{\mathcal{Q}}$ and it is always true that ${\mathcal{P}}\vee{\mathcal{Q}}\overset{O_{D}(1)}{\prec}{\mathcal{R}}$ and ${\mathcal{S}}\overset{O_{D}(1)}{\prec}{\mathcal{P}}$ and ${\mathcal{S}}\overset{O_{D}(1)}{\prec}{\mathcal{Q}}$. Thus, the inequality (6.17) follows from Lemma˜6.10. ∎
We show that $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ is a sufficient condition for the supercritical projection theorem.
Lemma 6.12 (Supercritical projection).
Let $k\in\llbracket 1,D-1\rrbracket$, let $c,\varepsilon,\rho>0$. Let $\Xi$ be a probability measure on $\operatorname{Gr}(\mathbb{R}^{D},k)$ satisfying either $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ with parameters $(\rho,\rho^{-\varepsilon},c)$.
Let $A\subseteq B^{\mathbb{R}^{D}}_{1}$ be any subset satisfying for some $\alpha\in[c,1-c]$,
(6.18) ${\mathcal{N}}_{\rho}(A)\geq\rho^{-D\alpha+\varepsilon},$
and for $r\geq\rho$,
(6.19) $\sup_{v\in\mathbb{R}^{D}}{\mathcal{N}}_{\rho}(A\cap B^{\mathbb{R}^{D}}_{r}(v))\leq\rho^{-\varepsilon}r^{c}{\mathcal{N}}_{\rho}(A).$
If $\varepsilon,\rho\lll_{D,c}1$, then the exceptional set
$\begin{split}{\mathcal{E}}:=\{\,F\in\operatorname{Gr}(\mathbb{R}^{D},k)\,:\,\exists A^{\prime}\subseteq A\,\,&\text{ with }\,\,{\mathcal{N}}_{\rho}(A^{\prime})\geq\rho^{\varepsilon}{\mathcal{N}}_{\rho}(A)\\ &\text{ and }\,\,{\mathcal{N}}_{\rho}(\pi_{F}A^{\prime})<\rho^{-\alpha k-\varepsilon}\}\end{split}$
satisfies $\Xi({\mathcal{E}})\leq\rho^{\varepsilon}$.
Proof.
We focus on the scenario where $\Xi$ satisfies $\operatorname{(MNC)}^{\perp}$. The case where $\Xi$ satisfies $\operatorname{(MNC)}$ can be handled similarly, and is only easier to justify as it involves (6.16) instead of (6.17).
Let $(q,m)$ be the couple of integers playing a role in the assumption $\operatorname{(MNC)}^{\perp}$ for $\Xi$. If $q=1$, the result is known. It is indeed the higher rank version of Bourgain’s projection theorem [13], due to the second-named author [26]. We deduce from there the general case $q\geq 1$. Note that throughout the proof, we may assume $A$ to be $2\rho$-separated. We may also allow the upper bound on $\rho$ to depend [^4] Indeed, if we establish the lemma for a pair $(\varepsilon,\rho)$ then it is automatically valid for $(\varepsilon^{\prime},\rho)$ with $\varepsilon^{\prime}\in(0,\varepsilon)$, because when passing from $\varepsilon$ to $\varepsilon^{\prime}$, assumptions get stronger and the conclusion gets weaker. on $\varepsilon$ (not only $D,c$). We let $\varepsilon_{1},\varepsilon_{2}>0$ be parameters to specify below in terms of $D$ and $c$. We use the shorthand $\mathcal{G}:=\operatorname{Gr}(\mathbb{R}^{D},k)$.
Provided $\varepsilon+\varepsilon_{1}\leq c$, the assumption that $\Xi$ enjoys $\operatorname{(MNC)}^{\perp}$ with parameters $(\rho,\rho^{-\varepsilon},c)$ implies
${\mathcal{E}}_{1}:=\{\,{\underline{F}}\in\mathcal{G}^{q}:\|F_{1}^{\perp}\wedge\dots\wedge F_{q}^{\perp}\|\leq\rho^{(\varepsilon+\varepsilon_{1})/c}\}\,\,\,\text{ satisfies}\,\,\,\Xi^{\otimes q}({\mathcal{E}}_{1})\leq\rho^{\varepsilon_{1}}.$
Let ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus{\mathcal{E}}_{1}$. Up to assuming $\varepsilon\leq\varepsilon_{1}$, Equation (6.17) implies that for every set $S\subseteq A$, there exists a subset $S^{\prime}\subseteq S$ such that $|S^{\prime}|\geq\rho^{O_{D,c}(\varepsilon_{1})}|S|$ and
(6.20) $\prod_{i=1}^{q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}S)\geq\rho^{O_{D,c}(\varepsilon_{1})}|S|^{q-1}{\mathcal{N}}_{\rho}(\pi_{\cap_{i}F_{i}}S^{\prime}).$
Taking $S$ to be a not too small subset of $A$, we use (6.20) to obtain an explicit lower bound on $\max_{i=1,\dots,q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}S)$, see (6.22). As a lower bound on $|S|^{q-1}$ comes directly from the assumption (6.18), we focus on ${\mathcal{N}}_{\rho}(\pi_{\cap_{i}F_{i}}S^{\prime})$.
Let $\Upsilon$ denote the restriction of $\Xi^{\otimes q}$ to $\mathcal{G}^{q}\smallsetminus{\mathcal{E}}_{1}$, renormalised into a probability measure. Note that the random Dq(Dk)D-q(D-k)-plane $(\cap_{i}F_{i})_{{\underline{F}}\sim\Upsilon}$ satisfies the non-concentration condition $\operatorname{(MNC)}^{\perp}$ with parameters $(\rho,\rho^{-\varepsilon-\varepsilon_{1}},c)$ and $q=1$. Therefore, provided $\varepsilon+\varepsilon_{1},\varepsilon_{2}\lll_{D,c}1$ and $\rho\lll_{D,c}1$, there is an event ${\mathcal{E}}_{2}\subseteq\mathcal{G}^{q}$ such that $\Upsilon({\mathcal{E}}_{2})\leq\rho^{\varepsilon_{2}}$, and for all ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus{\mathcal{E}}_{2}$, for all $A^{\prime}\subseteq A$ with $|A^{\prime}|\geq\rho^{\varepsilon_{2}}|A|$, we have
(6.21) ${\mathcal{N}}_{\rho}(\pi_{\cap_{i}F_{i}}A^{\prime})\geq\rho^{-\alpha(D-q(D-k))-\varepsilon_{2}}.$
Combining (6.20), (6.18) and (6.21), we obtain that for all ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus({\mathcal{E}}_{1}\cup{\mathcal{E}}_{2})$, and $A^{\prime\prime}\subseteq A$ with $|A^{\prime\prime}|\geq\rho^{\varepsilon_{1}}|A|$, we have
(6.22) $\max_{i=1,\dots,q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}A^{\prime\prime})\geq\rho^{-\alpha k-\frac{1}{2q}\varepsilon_{2}},$
provided $\varepsilon\leq\varepsilon_{1}\lll_{D,c}\varepsilon_{2}\lll_{D,c}1$.
To conclude the proof, we argue by contradiction, assuming $\Xi({\mathcal{E}})>\rho^{\varepsilon}$. For each $F\in{\mathcal{E}}$, let $A^{\prime}_{F}\subseteq A$ be such that $\lvert A^{\prime}_{F}\rvert\geq\rho^{\varepsilon}\lvert A\rvert$ and ${\mathcal{N}}_{\rho}(\pi_{F}A^{\prime}_{F})<\rho^{-\alpha k-\varepsilon}$. By a Fubini argument such as [26, Lemma 19], $q$ independent copies of $A^{\prime}_{F}$ are likely to intersect in a rather large subset:
${\mathcal{E}}_{3}:=\bigl{\{}\,{\underline{F}}\,:\,\lvert A^{\prime}_{F_{1}}\cap\dots\cap A^{\prime}_{F_{q}}\rvert\leq\rho^{2q\varepsilon}|A|\,\bigr{\}}\,\,\,\text{ satisfies}\,\,\,\Xi^{\otimes q}({\mathcal{E}}_{3})\leq 1-\rho^{4q\varepsilon}.$
Note that $\cup_{i=1}^{3}{\mathcal{E}}_{i}$ has $\Xi^{\otimes q}$-measure bounded above by $\rho^{\varepsilon_{1}}+\rho^{\varepsilon_{2}}+1-\rho^{4q\varepsilon}$ which is strictly less than $1$ provided $\varepsilon\lll_{D}\varepsilon_{1}\leq\varepsilon_{2}$ and $\rho\lll_{\varepsilon}1$. In particular, we may consider ${\underline{F}}\in\mathcal{G}^{q}\smallsetminus\cup_{i=1}^{3}{\mathcal{E}}_{i}$. Setting $A^{\prime\prime}=A^{\prime}_{F_{1}}\cap\dots\cap A^{\prime}_{F_{q}}$, we have $|A^{\prime\prime}|\geq\rho^{2q\varepsilon}|A|$ because ${\underline{F}}\notin{\mathcal{E}}_{3}$, while the inclusions $A^{\prime\prime}\subseteq A^{\prime}_{F_{i}}$ yield
$\max_{i=1,\dots q}{\mathcal{N}}_{\rho}(\pi_{F_{i}}A^{\prime\prime})<\rho^{-\alpha k-\varepsilon}.$
This is in contradiction with (6.22) for $\varepsilon\lll_{D}\varepsilon_{1}\leq\varepsilon_{2}$. ∎
Without non-concentration assumption on $A$, we still derive from $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ a subcritical projection theorem.
Lemma 6.13 (Subcritical projection).
Let $k\in\llbracket 1,D-1\rrbracket$, let $C>1$ and $c,\varepsilon,\rho\in(0,1/2]$. Let $\Xi$ be a probability measure on $\operatorname{Gr}(\mathbb{R}^{D},k)$ satisfying either $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ with parameters $(\rho,\rho^{-\varepsilon},c)$. Let $A\subseteq B^{\mathbb{R}^{D}}_{1}$ be any subset.
If $C\ggg_{D,c}1$ and $\rho\lll_{D,c,\varepsilon}1$, then
$\begin{split}{\mathcal{E}}:=\{\,F\in\operatorname{Gr}(\mathbb{R}^{D},k)\,:\,\exists A^{\prime}\subseteq A\,\,&\text{ with }\,\,{\mathcal{N}}_{\rho}(A^{\prime})\geq\rho^{\varepsilon}{\mathcal{N}}_{\rho}(A)\\ &\text{ and }\,\,{\mathcal{N}}_{\rho}(\pi_{F}A^{\prime})<\rho^{C\varepsilon}{\mathcal{N}}_{\rho}(A)^{\frac{k}{D}}\}\end{split}$
satisfies $\Xi({\mathcal{E}})\leq\rho^{\varepsilon}$.
Proof.
The proof is similar to that of Lemma˜6.12, using the subcritical projection theorem [6, Proposition A.2] instead of the supercritical projection theorem. ∎
We now combine Lemmas 6.12, 6.13 into a multislicing estimate. We place ourselves in $\mathbb{R}^{D}$ where $D\geq 3$. We consider $d_{1},d_{2}\in\mathbb{N}$ such that $1\leq d_{1}<d_{2}<D$ and ${\bf t}=(t_{1},t_{2},t_{3})\in\mathbb{R}^{3}$ such that $0\leq t_{1}<t_{2}<t_{3}\leq 1$. We set $\mathcal{F}$} the collection of couples $\mathscr{V}=(V_{1},V_{2})$ where $V_{i}\in\operatorname{Gr}(\mathbb{R}^{d},d_{i})$ for $i=1,2$ and $V_{1}\subseteq V_{2}$. Given $\mathscr{V}\in{\mathcal{F}}$ and $\rho\in(0,1)$, we set
$B_{\rho^{\bf t}}^{\mathscr{V}}=B^{V_{1}}_{\rho^{t_{1}}}+B^{V_{2}}_{\rho^{t_{2}}}+B^{\mathbb{R}^{d}}_{\rho^{t_{3}}}.$
Therefore $B_{\rho^{\bf t}}^{\mathscr{V}}$ represents a Euclidean box carried by the partial flag $\mathscr{V}$ and of side length parameters $\rho^{t_{1}}>\rho^{t_{2}}>\rho^{t_{3}}$. The multislicing theorem below considers a random partial flag $\mathscr{V}$, and a measure $\nu$ which is Frostman above scale $\rho$. For most realizations of $\mathscr{V}$, it gives an upper bound on the mass granted by $\nu$ to all translates of $B_{\rho^{\bf t}}^{\mathscr{V}}$. It requires a certain assumption on $\mathscr{V}$, namely that each component of $\mathscr{V}$ satisfies $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ above scale $\rho$.
Proposition 6.14 (Supercritical multislicing).
Let $D\geq 3$ and $d_{1},d_{2},{\bf t}$ be as above. Let $c,\varepsilon,\rho>0$.
Let $\Xi$ be a probability measure on $\mathcal{F}$}. Assume that for each $i=1,2$, the distribution of the component $V_{i}$ as $\mathscr{V}\sim\Xi$ satisfies either $\operatorname{(MNC)}$ or $\operatorname{(MNC)}^{\perp}$ with parameters $(\rho,\rho^{-\varepsilon},c)$.
Let $\nu$ be a Borel measure on $B^{\mathbb{R}^{D}}_{1}$ such that for some $\alpha\in[c,1-c]$, for all $v\in\mathbb{R}^{D}$, and $r\geq\rho$, we have
$\nu(B^{\mathbb{R}^{D}}_{r}+v)\leq\rho^{-\varepsilon}r^{D\alpha}.$
If $\varepsilon,\rho\lll_{D,{\bf t},c}1$, then there exists an event $\mathcal{E}$$\subseteq{\mathcal{F}}$ such that $\Xi({\mathcal{E}})\leq\rho^{\varepsilon}$ and for $\mathscr{V}\in{\mathcal{F}}\smallsetminus{\mathcal{E}}$, there is a set $A_{\mathscr{V}}\subseteq\mathbb{R}^{D}$ with $\nu(\mathbb{R}^{D}\smallsetminus A_{\mathscr{V}})\leq\rho^{\varepsilon}$ and such that for every $v\in\mathbb{R}^{D}$,
$\nu_{|A_{\mathscr{V}}}\left(B^{\mathscr{V}}_{\rho^{\bf t}}+v\right)\leq\operatorname{Leb}\left(B^{\mathscr{V}}_{\rho^{\bf t}}\right)^{\alpha+\varepsilon}.$
Proof.
Lemmas 6.12, 6.13 guarantee that the random projectors $\pi_{||V_{1}}$ and $\pi_{||V_{2}}$ where $\mathscr{V}\sim\Xi$ satisfy respectively subcritical and supercritical estimates. Those can be combined as in the original paper [6, Section 2] into the above result. More formally, the deduction is a direct consequence of [5, Theorem 3.4].∎
6.3. Linearizing charts
Recall $X=G/\Lambda$ where $G=\operatorname{SL}_{d+1}(\mathbb{R})$ and $\Lambda$ is a fixed arbitrary lattice. We define on $X$ a covering of linearizing charts which do not deform balls much, and most importantly send any $g$-translate $gB_{r}x$ ($g\in G,r>0,x\in X$) to an additive translate of the box $\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}$ provided $\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}$ is not too distorted and lives at a suitable scale. We point out that contrary to our previous work [7] where $d=1$, those charts live at a microscopic scale. This linearizing scheme is extracted from [5, Lemma 6.3], which is itself inspired by Shmerkin [51].
Lemma 6.15 ([5]).
Let $0<\delta\leq\eta\lll 1$. There exists a measurable map $\varphi:\{\operatorname{inj}\geq\eta\}\rightarrow B^{\mathfrak{g}}_{1}$ satisfying the following.
  • 1)
    For every $r\in(0,\eta)$, $v\in\mathfrak{g}$, the preimage $\varphi^{-1}(B^{\mathfrak{g}}_{r}+v)$ is covered by $O(1)$ many balls $(B_{r}x)_{x\in X}$
  • 2)
    For every $r\in(0,\eta)$, $g\in G$ such that $B^{\mathfrak{g}}_{\delta^{2}}\subseteq\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}\subseteq B^{\mathfrak{g}}_{\delta}$, and $x\in X$, the translate $gB_{r}x\cap\{\operatorname{inj}\geq\eta\}$ is covered by $O(1)$ many preimages of boxes $(\varphi^{-1}(\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}+v))_{v\in\mathfrak{g}}$.
In this lemma, $\eta$ controls the region of $X$ which is linearized and the maximum scale at which linearization occurs. On the other hand, $\delta$ controls the distortion allowed on $G$-translates of balls $gB_{r}x$ to be well represented by additive translates of boxes $\operatorname{Ad}(g)B^{\mathfrak{g}}_{r}+v$ via the linearization.
6.4. Dimension increment and bootstrap
We combine the results of the three previous subsections to show that the dimensional properties of a prescribed measure $\nu$ on $X$ are improved under the action of the $\mu$-random walk on $X$. This is Proposition˜6.16. By iteration, we deduce the desired bootstrap to high dimension, Proposition˜6.2.
Recall that $\ell>0$ denotes the top Lyapunov exponent of the $\operatorname{Ad}_{\star}\mu$-random walk on $\mathfrak{g}$, see (2.1).
Proposition 6.16 (Dimension increment).
Let $\kappa,\varepsilon,\rho\in(0,1/10)$, $\alpha\in{[\kappa,1-\kappa]}$, $\tau\geq 0$ be some parameters. Consider on $X$ a Borel measure $\nu$ of mass at most $1$ and which is $(\alpha,{\mathcal{B}}_{[\rho,\rho^{\varepsilon}]},\tau)$-robust. Denote by $n_{\rho}\geq 0$ the integer part of $\frac{1}{10\ell}\lvert\log\rho\rvert$.
Assume $\varepsilon,\rho\lll_{\kappa}1$, then
$\text{$\mu^{*n_{\rho}}*\nu$ is $(\alpha+\varepsilon,{\mathcal{B}}_{\rho^{1/2}},\tau+\rho^{\varepsilon})$-robust}.$
Proof.
We may assume $\tau=0$. We write $n=n_{\rho}$. By Proposition˜4.1 integrated over $\nu$, we have
$\mu^{*n}*\nu\{\operatorname{inj}\leq\rho^{1/2}\}\ll\rho^{c/2}(e^{-cn}\rho^{-C\varepsilon}+1)$
for some constants $c>0,C>1$ depending on $\Lambda$ and $\mu$. We can require $\varepsilon\lll\frac{c}{C(\ell+1)}$ and $\rho\lll_{c}1$ so that this leads to $\mu^{*n}*\nu\{\operatorname{inj}\leq\rho^{1/2}\}\leq\frac{\rho^{\varepsilon}}{2}$. Thus, it remains to show that $\mu^{*n}*\nu$ can be written as a sum $\mu^{*n}*\nu=\nu^{\prime}+\nu^{\prime\prime}$ of Borel measures satisfying $\nu^{\prime\prime}(X)\leq\frac{\rho^{\varepsilon}}{2}$ and
(6.23) $\sup_{y\in X}\nu^{\prime}(B_{\rho^{1/2}}y)\leq\rho^{\frac{1}{2}(\alpha+\varepsilon)\dim X}.$
To this end, we first linearize the situation by looking through the covering of charts from §6.3. More precisely, we apply Lemma˜6.15 with parameters $\eta=\rho^{\varepsilon}$ and $\delta=\rho^{1/3}$. This yields a map $\varphi:\{\operatorname{inj}\geq\rho^{\varepsilon}\}\rightarrow B^{\mathfrak{g}}_{1}$, we set ${\tilde{\nu}}=\varphi_{\star}\nu$. The assumption that $\nu$ is $(\alpha,{\mathcal{B}}_{[\rho,\rho^{\varepsilon}]},0)$-robust and has mass at most $1$ implies, via Lemma˜6.15 item 1) and provided $\rho\lll_{\varepsilon}1$, that for every $r\geq\rho$,
$\sup_{v\in\mathfrak{g}}{\tilde{\nu}}(B^{\mathfrak{g}}_{r}+v)\leq\rho^{-\varepsilon\dim X}r^{\alpha\dim X}.$
We now aim to apply Proposition˜6.14 to the measure $\tilde{\nu}$}, and for the random box $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}$ where $g\sim\mu^{*n}$, or rather its close companion
$B^{\mathscr{V}_{g}}_{\rho^{\bf t}}:=B^{\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-}}_{\rho^{2/5}}+B^{\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0}}_{\rho^{1/2}}+B^{\mathfrak{g}}_{\rho^{3/5}}$
which is a good approximation of $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}$ (by (6.25) below), and whose partial flag we know how to control thanks to §6.1. Indeed, by Proposition˜6.4 and Corollary˜6.5, the distributions of $(\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{-})_{g\sim\mu^{*n}}$ and $(\operatorname{Ad}(u(-\mathtt{b}_{g}))\mathfrak{g}_{\leq 0})_{g\sim\mu^{*n}}$ satisfy respectively $\operatorname{(MNC)}$ and $\operatorname{(MNC)}^{\perp}$ with parameters $(e^{-n},C,c)$, or equivalently $(\rho,C,c)$ up to dividing $c$ by $11\ell$. Here $C,c>0$ are constants that only depend on $\mu$.
Provided $\rho,\varepsilon\lll 1$, the multislicing Proposition˜6.14 yields a subset $E_{1}\subseteq G$ and some constant $\varepsilon_{0}=\varepsilon_{0}(\mu)>0$ such that $\mu^{*n}(E_{1})\leq\rho^{\varepsilon_{0}}$ and for $g\in G\smallsetminus E_{1}$, there exists a set $\tilde{A}_{g}\subseteq\mathfrak{g}$ with ${\tilde{\nu}}(\mathfrak{g}\smallsetminus\tilde{A}_{g})\leq\rho^{\varepsilon_{0}}$ and such that for every $v\in\mathfrak{g}$,
(6.24) ${\tilde{\nu}}_{|\tilde{A}_{g}}\left(B^{\mathscr{V}_{g}}_{\rho^{\bf t}}+v\right)\leq\rho^{\varepsilon_{0}}\operatorname{Leb}\left(B^{\mathscr{V}}_{\rho^{\bf t}}\right)^{\alpha}.$
On the other hand, by the large deviation principle for $\log\mathtt{r}_{g}$ and Lemma˜3.5 item (i), there exists a subset $E_{2}\subseteq G$ and a constant $\gamma=\gamma(\mu,\varepsilon)>0$ such that $\mu^{*n}(E_{2})\ll\rho^{\gamma}$, and for every $g\in G\smallsetminus E_{2}$, we have $\mathtt{r}_{g}\in[\rho^{\frac{1}{10}+\varepsilon},\rho^{\frac{1}{10}-\varepsilon}]$ and $\|\mathtt{b}_{g}\|\leq\rho^{-\varepsilon}$. Combined with (6.3), we obtain
(6.25) $B^{\mathscr{V}_{g}}_{\rho^{\bf t}}\,\overset{\rho^{-O(\varepsilon)}}{\simeq}\,\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}.$
Equations (6.24) and (6.25) together imply that for every $g\in G\smallsetminus(E_{1}\cup E_{2})$ and $v\in\mathfrak{g}$,
(6.26) ${\tilde{\nu}}_{|\tilde{A}_{g}}\left(\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}+v\right)\leq\rho^{\varepsilon_{0}-O(\varepsilon)}\operatorname{Leb}\left(B^{\mathscr{V}}_{\rho^{\bf t}}\right)^{\alpha}.$
We now get back to $X$. To control the distortion of $\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}$, we observe for $g\in G\smallsetminus E_{2}$, we have $B^{\mathfrak{g}}_{\rho^{2/3}}\subseteq\operatorname{Ad}(g^{-1})B^{\mathfrak{g}}_{\rho^{1/2}}\subseteq B^{\mathfrak{g}}_{\rho^{1/3}}$ provided $\varepsilon\lll 1$. Applying Lemma˜6.15 item $2)$ and (6.26), we deduce that for all $g\in G\smallsetminus\cup_{i=1}^{2}E_{i}$, setting $A_{g}=\varphi^{-1}(\tilde{A}_{g})$, we have $\nu(G\smallsetminus A_{g})\leq\rho^{\varepsilon_{0}}$ and for every $y\in G$,
$g_{\star}\nu_{|A_{g}}(B_{\rho^{1/2}}y)=\nu_{|A_{g}}\left(g^{-1}B_{\rho^{1/2}}y\right)\ll\rho^{\varepsilon_{0}-O(\varepsilon)}\operatorname{Leb}\left(B^{\mathscr{V}}_{\rho^{\bf t}}\right)^{\alpha}\ll\rho^{\frac{1}{2}\alpha\dim X+\varepsilon_{0}-O(\varepsilon)}.$
Taking $\nu^{\prime}=\int_{G\smallsetminus\cup_{i=1}^{2}E_{i}}g_{\star}\nu_{|A_{g}}\,\mathrm{d}\mu^{*n}(g)$, and $\varepsilon\lll\varepsilon_{0}$, $\rho\lll_{\varepsilon}1$, this concludes the proof of (6.23), whence that of the proposition. ∎
We now deduce high dimension (Proposition˜6.2) from the combination of effective recurrence (Proposition˜4.1), initial positive dimension (Proposition˜5.1), and dimension increment (Proposition˜6.16).
Proof of Proposition˜6.2.
Let $A>0$ be a large enough constant depending on the initial data $\mu$. Combining Proposition˜5.1 and Proposition˜4.1, we may assume $\kappa>0$ small enough from the start, so that for any $M>0$, for every $\rho\lll_{M}1$ and $n\geq M\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}(x)\rvert$, the measure
$\text{{$\mu^{*n}*\delta_{x}$} is $(\kappa,{\mathcal{B}}_{[\rho^{M},\,\rho^{1/M}]},\rho^{\kappa/M})$-robust}.$
By Proposition˜6.16, there is some small constant $\varepsilon=\varepsilon(\mu,\kappa)>0$, such that up to imposing from the start $M\ggg_{\kappa}1$, we have for all $\rho\lll_{\kappa,M}1$, all $n\geq(\frac{1}{10\ell}+1)M\lvert\log\rho\rvert+A\lvert\log\operatorname{inj}(x)\rvert$ and $r\in[\rho^{2M\varepsilon},\rho^{1/(2M\varepsilon)}]$,
$\text{{$\mu^{*n}*\delta_{x}$} is $(\kappa+2\varepsilon,{\mathcal{B}}_{r^{1/2}},2\rho^{\kappa/M})$-robust}.$
These estimates for single scales can be combined using [6, Lemma 4.5] to get under the same conditions:
$\text{{$\mu^{*n}*\delta_{x}$} is $(\kappa+\varepsilon,{\mathcal{B}}_{[\rho^{M\varepsilon},\rho^{1/(M\varepsilon)}]},O_{\kappa,M}(\rho^{\kappa/M}))$-robust}.$
The argument in the last paragraph can be applied iteratively, adding at each step $k$ the value $\varepsilon$ to the dimension provided the latter is not yet above $1-\kappa$ and provided $M\ggg_{\kappa,k}1$. As the value of $\varepsilon$ only depends on $\mu,\kappa$, we reach dimension $1-\kappa$ in a finite number of steps. This concludes the proof. ∎
7. From high dimension to equidistribution
In this section, we establish ˜1.2 and ˜1.3. We further establish a double equidistribution estimate (Proposition˜7.3) which will be useful to prove the divergent case of ˜1.1.
We let $(\eta_{t})_{t>0}$ denote the one-parameter family of probability measures on $G$ defined by
$\mathrm{d}\eta_{t}:=a(t)u(\boldsymbol{s})\,\mathrm{d}\sigma(\boldsymbol{s}).$
The next proposition states that a probability measure $\nu$ on $X$ with dimension close to $\dim X$ equidistributes with exponential rate under convolution with $\eta_{t}$. It is in fact slightly more precise as the dimension assumption on $\nu$ concerns only a single scale $\rho$, and equidistribution is guaranteed for a corresponding interval of times $t\in[\rho^{-1/2},\rho^{-1/4}]$.
Proposition 7.1.
There exist $\kappa,\rho_{0}>0$ such that the following holds for all $\rho\in(0,\rho_{0}]$ and $\tau\in\mathbb{R}_{\geq 0}$.
Let $\nu$ be a Borel measure on $X$ which is $(1-\kappa,{\mathcal{B}}_{\rho},\tau)$-robust with $\nu(X)\leq 1$. Set $l=\lceil\frac{1}{2}\dim\operatorname{SO}(d+1)\rceil$. Then for all $t\in[\rho^{-1/2},\rho^{-1/4}]$, for all $f\in B_{\infty,l}^{\infty}(X)$ with $m_{X}(f)=0$, we have
(7.1) $|\eta_{t}*\nu(f)|\leq(\rho^{\kappa}+\tau){\mathcal{S}}_{\infty,l}(f).$
Proof.
The proof is similar to that of [7, Proposition 5.1]. We provide a sketch for completeness and refer the reader to [7] for details.
Denote by $(P_{\eta_{t}})_{t>0}$ the family of Markov operators on $L^{2}(X)$ associated to $(\eta_{t})_{t>0}$. It is defined by: $\forall f\in L^{2}(X)$,
$P_{\eta_{t}}f=\int_{G}f(g\,\cdot)\,\mathrm{d}\eta_{t}(g).$
The first step of the proof is to show a spectral gap property for $(P_{\eta_{t}})_{t>0}$ as $t\to+\infty$, namely: there exists $c=c(G,\Lambda,\sigma)>0$ such that for all $f\in B^{\infty}_{2,l}(X)$ with $m_{X}(f)=0$, all $t>1$, one has
(7.2) $\|P_{\eta_{t}}f\|_{L^{2}}\ll t^{-c}{\mathcal{S}}_{2,l}(f).$
The proof of (7.2) exploits the quantitative decay of matrix coefficients (see [2, Lemma 3] and [22, Equations (6.1), (6.9)]):
$\exists\delta_{0}=\delta_{0}(\Lambda)>0,\,\forall g\in G,\quad\lvert\langle f(g\,\cdot),f\rangle_{L^{2}}\rvert\ll\lVert g\rVert^{-\delta_{0}}{\mathcal{S}}_{2,l}(f)^{2},$
and the non-concentration property of $\sigma$ from Proposition˜3.2, see [7, Proposition 5.2] for details.
Once (7.2) is established, we obtain (7.1) as follows. We introduce $\nu_{\rho}$ the mollification of $\nu$ at scale $\rho$, namely
$\nu_{\rho}:=\frac{1}{m_{G}(B_{\rho})}\int_{B_{\rho}}g_{\star}\nu\,\mathrm{d}m_{G}(g).$
Given $f\in B_{\infty,l}^{\infty}(X)$ with $m_{X}(f)=0$, we then have for every $t>1$,
$|\eta_{t}*\nu(f)|=\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu\right\rvert\leq\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu-\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu_{\rho}\right\rvert\,+\,\left\lvert\int_{X}P_{\eta_{t}}f\,\mathrm{d}\nu_{\rho}\right\rvert$.
The first integral in the right hand side is bounded by $\rho\operatorname{Lip}(P_{\eta_{t}}f)\ll\rho t{\mathcal{S}}_{\infty,l}(f).$ To bound the second integral, note we may assume from the start $\tau=0$. Then $\nu_{\rho}$ satisfies $\mathrm{d}\nu_{\rho}(x)=\frac{\nu(B_{\rho}x)}{m_{G}(B_{\rho})}\,\mathrm{d}m_{X}(x)\ll\rho^{-\kappa\dim X}\,\mathrm{d}m_{X}(x)$. Applying (7.2), we find the second term in the right hand side is bounded by $O(t^{-c}\rho^{-\kappa\dim X}{\mathcal{S}}_{2,l}(f)).$ The proof is concluded by taking $t\in[\rho^{-1/2},\rho^{-1/4}]$, $\kappa$ small enough in terms of $c,\dim X$, and $\rho_{0}$ small enough in terms of $G$, $\kappa$. ∎
In the next lemma, we invoke the self-similarity of $\sigma$ to relate $\eta_{t}$ and convolution powers of $\mu$.
Lemma 7.2 ($\eta_{t}$-process vs $\mu$-walk).
Given $t>0$, $n\geq 0$, we have
$\eta_{t}=\int_{P^{\prime}}\delta_{k_{g}}*\eta_{t\mathtt{r}_{g}}*\delta_{g}\,\mathrm{d}\mu^{*n}(g).$
Proof.
We observe that for any $\boldsymbol{s}\in\mathbb{R}^{d}$ and $g\in P^{\prime}$,
$k_{g}a(t\mathtt{r}_{g})u(\boldsymbol{s})g$ $=a(t)a(\mathtt{r}_{g})k_{g}u(\boldsymbol{s})k_{g}^{-1}a(\mathtt{r}_{g}^{-1})u(\mathtt{b}_{g})$
$=a(t)u(\mathtt{r}_{g}O_{g}\boldsymbol{s}+\mathtt{b}_{g})$
$=a(t)u(\phi_{g}(\boldsymbol{s})).$
The lemma follows by the equality $\lambda^{*n}*\sigma=\sigma$. ∎
We are now able to conclude the proof of ˜1.2. The strategy is to use Lemma˜7.2 to decompose $\eta_{t}$ as a random walk part $\mu^{*n}$ (where $n=n(\mu,t)$) which generates high dimension thanks to Proposition˜6.2, followed by some $\eta_{t^{\prime}}$-part (with $t^{\prime}=t^{\prime}(\mu,t)$) which will convert this high dimension into equidistribution via Proposition˜7.1. The apparent obstruction is that the decomposition appearing Lemma˜7.2 does not separate the $\mu$ part and the $\eta$ part, because the term $\delta_{k_{g}}*\eta_{t\mathtt{r}_{g}}$ involves $g$. To deal with this obstacle, we partition the space of parameters $g$ into $O(\rho^{-\alpha})$ subsets ($\alpha\lll_{\kappa}1$) in which $\delta_{k_{g}}*\eta_{t\mathtt{r}_{g}}$ hardly depends on $g$.
Proof of ˜1.2.
Up to replacing $\lambda$ by a suitable convolution power until a stopping time (as in [7, Lemma 5.3]), we may assume that $\lambda$-almost every $\phi$ is orientation preserving. This way we place ourselves in the context of Section˜2, and all results obtained until now are applicable.
Observe that given $t,r_{0},r_{1}>0$ and $k_{0},k_{1}\in K^{\prime}$, we have
$\delta_{k_{0}}*\eta_{tr_{0}}=\delta_{k_{0}k_{1}^{-1}a(r_{0}r_{1}^{-1})}*\delta_{k_{1}}*\eta_{tr_{1}}.$
Hence, given any Borel measure $\nu$ on $X$ and $f\in B_{\infty,l}^{\infty}(X)$, we have
(7.3) $\lvert\delta_{k_{0}}*\eta_{tr_{0}}*\nu(f)-\delta_{k_{1}}*\eta_{tr_{1}}*\nu(f)\rvert\ll(\lVert\operatorname{Id}-k_{0}k_{1}^{-1}\rVert+\lvert\log r_{0}r_{1}^{-1}\rvert){\mathcal{S}}_{\infty,l}(f)\nu(X).$
Let $\alpha,\rho>0$ be parameters to be specified later, with $\alpha$ depending only on $\Lambda,\mu$, and $\rho$ on $\Lambda,\mu,t$. We discretize the set of $k_{g}$ and $\mathtt{r}_{g}$ for $g\in P^{\prime}$ as follows. We partition the compact group $K^{\prime}$ into $\rho^{-O(\alpha)}$ disjoint measurable sets $\{K^{\prime}_{i}:i\in I\}$ such that each $K^{\prime}_{i}$ is contained in a ball of radius $\rho^{\alpha}$ centered at some $k_{i}\in K^{\prime}$. We set
${\mathscr{R}}:=\{(1+\rho^{\alpha})^{k}\,:\,k\in\mathbb{Z}\}.$
For $i\in I,r\in{\mathscr{R}}$, we define
$P^{\prime}(i,r):=\{g\in P^{\prime}\,:\,k_{g}\in K^{\prime}_{i}\ \text{ and }\mathtt{r}_{g}\in[r,r(1+\rho^{\alpha})[\}.$
Then
$P^{\prime}=\bigsqcup_{\mathtt{r}\in{\mathscr{R}},i\in I}P^{\prime}(\mathtt{r},i).$
Hence, by Lemma˜7.2 and (7.3), we obtain for every $n\geq 1$,
$\lvert\eta_{t}*\delta_{x}(f)\rvert$ $=\lvert\int_{P^{\prime}}\delta_{k_{g}}*\eta_{t\mathtt{r}_{g}}*\delta_{g}*\delta_{x}(f)d\mu^{*n}(g)\rvert$
(7.4) $\leq\sum_{i\in I,r\in{\mathscr{R}}}\lvert\delta_{k_{i}}*\eta_{tr}*\mu^{*n}_{|P^{\prime}(i,r)}*\delta_{x}(f)\rvert+O(\rho^{\alpha}{\mathcal{S}}_{\infty,l}(f)).$
We now bound each term in the sum from (7). Let $\kappa=\kappa(\Lambda,\mu)>0$ as in Proposition˜7.1. Assume $\operatorname{inj}(x)\geq\rho$. By Proposition˜6.2, there are constants $C=C(\Lambda,\mu)>1$ and $\varepsilon_{1}=\varepsilon_{1}(\Lambda,\mu)>0$ such that, provided $\rho\lll 1$, the measure $\mu^{*n}*\delta_{x}$ on $X$ is $(1-\kappa,{\mathcal{B}}_{\rho},\rho^{\varepsilon_{1}})$-robust for any $n\geq C\lvert\log\rho\rvert$. For the rest of this proof, we choose $\rho$ and $n$ depending on $\Lambda,\mu,t$ so that
(7.5) $t=\rho^{-C\ell-3/8}\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,\,n=\lceil C\lvert\log\rho\rvert\rceil$,
where $\ell$ has been defined in (2.1). Consider
${\mathscr{R}}^{\prime}=\{\,r\in{\mathscr{R}}:\rho^{-1/4}\leq tr\leq\rho^{-1/2}\,\}={\mathscr{R}}\cap[\rho^{C\ell+1/8},\rho^{C\ell-1/8}].$
On the one hand, by the principle of large deviations and our choice for $n$, we have for some $\varepsilon_{2}=\varepsilon_{2}(\mu,C)>0$,
(7.6) $\mu^{*n}\{g\,:\,\mathtt{r}_{g}\notin{\mathscr{R}}^{\prime}\}\leq\rho^{\varepsilon_{2}}$
On the other hand, for $i\in I$, $\mathtt{r}\in{\mathscr{R}}^{\prime}$, observing that $\mu^{*n}_{|P^{\prime}(i,r)}*\delta_{x}\leq\mu^{*n}*\delta_{x}$ is $(1-\kappa,{\mathcal{B}}_{\rho},\rho^{\varepsilon_{1}})$-robust, and ${\mathcal{S}}_{\infty,l}(f\circ k_{i})\ll{\mathcal{S}}_{\infty,l}(f)$ , $m_{X}(f\circ k_{i})=m_{X}(f)$, we get via Proposition˜7.1, for $t\ggg 1$,
(7.7) $\lvert\eta_{tr}*\mu^{*n}_{|P^{\prime}(i,r)}*\delta_{x}(f)\rvert\leq(\rho^{\kappa}+\rho^{\varepsilon_{1}}){\mathcal{S}}_{\infty,l}(f).$
Combining (7.5), (7.6), (7.7) and choosing $\alpha$ small enough in terms of $\varepsilon_{1},\varepsilon_{2},\kappa$, we obtain the bound announced by ˜1.2. So far, we have worked under the condition $\operatorname{inj}(x)\geq t^{-(C\ell+3/8)^{-1}}$. Noting the claim is trivial otherwise, the proof of ˜1.2 is complete.
Proof of ˜1.3.
By a similar argument, we see Proposition˜7.1 is still valid with $(t,\eta_{t})$ replaced by $(e^{k},\mu^{k})$ Replacing (7.2) by the equality $\mu^{*(k+n)}=\mu^{*k}*\mu^{*n}$, we can then argue as in the proof of ˜1.2. Details are left to the reader. ∎
Double equidistribution. We conclude this section by upgrading ˜1.2 into a double equidistribution property. This upgrade will play a role to prove the divergent case of the Khintchine dichotomy.
Given bounded measurable functions $f_{1},f_{2}:X\to\mathbb{R}$ and $t_{2}\geq t_{1}\geq 0$, we introduce the double equidistribution coefficient
(7.8) $\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2}):=\left\lvert\int_{\mathbb{R}^{d}}f_{1}\bigl{(}a(t_{1})u(\boldsymbol{s})x_{0}\bigr{)}f_{2}\bigl{(}a(t_{2})u(\boldsymbol{s})x_{0}\bigr{)}\,\mathrm{d}\sigma(\boldsymbol{s})-m_{X}(f_{1})m_{X}(f_{2})\right\rvert$.
We recall that in the above, $x_{0}=\Lambda/\Lambda$ denotes the identity coset of $X$.
The following proposition gives a quantitative upper bound on $\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})$ provided the times $t_{1},t_{2}$ are sufficiently separated.
Proposition 7.3 (Effective double equidistribution of expanded fractals).
For every $\eta>0$, there exist $C,c>0$ such that for all $t_{1},t_{2}>1$ with $t_{2}\geq t^{1+\eta}_{1}$ and $f_{1},f_{2}\in B^{\infty}_{\infty,l}(X)$, we have
(7.9) $\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{\infty,l}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,l}(f_{1}){\mathcal{S}}_{\infty,l}(f_{2})t_{2}^{-c}.$
Remark.
A quantitative bound on $\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})$ without separation condition on $t_{1},t_{2}$ will be extrapolated below, see Equation (8.3).
Proof.
The proof is the same as that of [7, Proposition 6.1], using ˜1.2 in the place of [7, Theorem B’]. ∎
8. Quantitative Khintchine dichotomy in $\mathbb{R}^{d}$ from equidistribution
We show that an arbitrary probability measure $\xi$ on $\mathbb{R}^{d}$ obeys the Khintchine dichotomy provided that the pushfoward $a(t)u(\boldsymbol{s})\operatorname{SL}_{d+1}(\mathbb{Z})\,\mathrm{d}\xi(\boldsymbol{s})$ satisfies certain effective equidistribution properties in $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$ for large $t$. Combined with Proposition˜7.3, this yields ˜1.1.
Throughout the section, notations refer to Section˜2, and we further specify $\Lambda=\operatorname{SL}_{d+1}(\mathbb{Z})$, in particular $X=\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$. We also set $x_{0}=\Lambda/\Lambda\in X$.
Definition 8.1.
Let $\xi$ be a probability measure on $\mathbb{R}^{d}$. We say that $\xi$ satisfies the effective single equidistribution property on $X$ if there are constants $C,c>0$ and $l\in\mathbb{N}$ such that
(8.1) $\forall f\in B^{\infty}_{\infty,l}(X),\,\forall t>1,\\ \left\lvert\int_{\mathbb{R}^{d}}f\bigl{(}a(t)u(\boldsymbol{s})x_{0}\bigr{)}\,\mathrm{d}\xi(\boldsymbol{s})-m_{X}(f)\right\rvert\leq C{\mathcal{S}}_{\infty,l}(f)t^{-c}.$
We say that $\xi$ satisfies the effective double equidistribution property on $X$ if for every $\eta>0$, there are constants $C,c>0$ and $l\in\mathbb{N}$ such that
(8.2) $\forall f_{1},f_{2}\in B^{\infty}_{\infty,l}(X),\,\forall t_{1}>1,\,\forall t_{2}>t_{1}^{1+\eta},\\ \Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{\infty,l}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,l}(f_{1}){\mathcal{S}}_{\infty,l}(f_{2})t_{2}^{-c}.$
where the notation $\Delta^{\xi}_{f_{1},f_{2}}(t_{1},t_{2})$ is defined in (7.8).
Taking $f_{2}=1$ and $t_{2}\to+\infty$, we see that effective double equidistribution (8.2) implies effective single equidistribution (8.1). Note that (8.2) assumes some separation $t_{2}>t_{1}^{1+\eta}$ between $t_{1}$ and $t_{2}$. As it turns out, (8.2) (with small enough $\eta$) can in fact be automatically upgraded to the following full range estimate: there exist (potentially different) constants $C,c>0$, $l\in\mathbb{N}$, such that for every $f_{1},f_{2}\in B^{\infty}_{\infty,l}(X)$ and all $t_{2}\geq t_{1}>1$.
(8.3) $\Delta^{\sigma}_{f_{1},f_{2}}(t_{1},t_{2})\leq C{\mathcal{S}}_{2,l}(f_{1}){\mathcal{S}}_{2,l}(f_{2})t_{1}^{c}t_{2}^{-c}\\ +C{\mathcal{S}}_{\infty,l}(f_{1})\lvert m_{X}(f_{2})\rvert t_{1}^{-c}+C{\mathcal{S}}_{\infty,l}(f_{1}){\mathcal{S}}_{\infty,l}(f_{2})t_{2}^{-c}.$
The implication from (8.2) to (8.3) (again, parameters $C,c,l$ may differ) is explained in [7, Section 7.1]. The idea is that (8.2) implies single equidistribution, which in turn, by decay of matrix coefficients, yields effective double equidistribution in the short range regime $t_{1}\leq t_{2}\leq t_{2}^{1+\eta}$ for sufficiently small $\eta$.
The following result of Khalil-Luethi [28] guarantees that effective single equidistribution implies the convergent case of the Khintchine dichotomy.
Theorem 8.2 (Convergent case [28, Theorem 9.1]).
Let $\xi$ be a probability measure on $\mathbb{R}^{d}$ satisfying the effective single equidistribution property (8.1) on $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$. Then for every non-increasing function $\psi:\mathbb{N}\to\mathbb{R}_{\geq 0}$ such that $\sum_{q\in\mathbb{N}}\psi(q)^{d}<\infty$, we have
$\xi(W(\psi))=0.$
We show that effective double equidistribution implies the divergent case of the Khintchine dichotomy. Our result is in fact quantitative: we provide the asymptotic of the number of solutions of the Khintchine inequality when bounding the denominator.
Theorem 8.3 (Divergent case).
Let $\xi$ be a probability measure on $\mathbb{R}^{d}$ satisfying the effective double equidistribution property (8.2) on $\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$. Let $\psi:\mathbb{N}\to\mathbb{R}_{\geq 0}$ be a non-increasing function such that $\sum_{q\in\mathbb{N}}\psi(q)^{d}=\infty$.
Then $\xi(W(\psi))=1$ and for $\xi$-a.e. $\boldsymbol{s}\in\mathbb{R}^{d}$, we have as $N\to+\infty$:
(8.4) $\big{|}\{(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\llbracket 1,N\rrbracket\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(q)\}\big{|}\sim_{\boldsymbol{s},\psi}\sum_{q=1}^{N}\psi(q)^{d}.$
Remark.
A light variation of the proof allows to estimate the number of primitive solutions of the Khintchine inequality. More precisely, consider ${\mathcal{P}}(\mathbb{Z}^{d+1}):=\mathbb{Z}^{d+1}\smallsetminus\cup_{k\geq 2}k\mathbb{Z}^{d+1}$ the set of primitive vectors in $\mathbb{Z}^{d+1}$. Set ${\mathcal{P}}(\mathbb{Z}^{d+1})_{N}={\mathcal{P}}(\mathbb{Z}^{d+1})\cap(\mathbb{Z}^{d}\times\llbracket 1,N\rrbracket)$. Then for $N\to+\infty$, we have
(8.5) $\big{|}\{(\boldsymbol{p},q)\in{\mathcal{P}}(\mathbb{Z}^{d+1})_{N}\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(q)\}\big{|}\sim_{\boldsymbol{s},\psi}\zeta(d+1)^{-1}\sum_{q=1}^{N}\psi(q)^{d}.$
where $\zeta(t)=\sum_{n\geq 1}n^{-t}$ denotes the Riemann zeta function.
Note also that in both cases (non-primitive and primitive), given an arbitrary subset of subscripts $I\subseteq\llbracket 1,d\rrbracket$, we may replace the condition $0\leq qs_{i}-p_{i}<\psi(q)$ ($i\in I$) in the above sets by $0\leq p_{i}-qs_{i}<\psi(q)$ ($i\in I$) without affecting the asymptotic.
The work conducted in previous sections guarantees that self-similar measures satisfy the conditions required in Theorems 8.2, 8.3. Provided ˜8.3 holds, we directly deduce ˜1.1.
Proof of ˜1.1.
The convergent case follows from combining Theorems 1.2, 8.2. The divergent case is a consequence of Proposition˜7.3 and ˜8.3 (along with its subsequent remark). ∎
It remains to show ˜8.3. We will distinguish the cases where $d\geq 2$ and $d=1$. The reason for that distinction is that the Siegel transform of the characteristic function of a ball in $\mathbb{R}^{d}$ has finite second moment when $d\geq 2$ and infinite second moment when $d=1$. We present the case $d\geq 2$ and explain afterward how the proof can be adapted, by the mean of a suitable truncation, to obtain the case $d=1$.
8.1. Case $d\geq 2$, the lower bound
We show the lower asymptotic in ˜8.3.
Given $N\geq 1$ and $\boldsymbol{s}\in\mathbb{R}^{d}$, we denote the left-hand side of (8.4) by
$\mathscr{T}_{N}(\boldsymbol{s}):=\big{|}\{(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\llbracket 1,N\rrbracket\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(q)\}\big{|}.$
We extend $\psi$ to a non-increasing function on $\mathbb{R}_{\geq 0}$ by setting $\psi(q)=\psi(\lceil q\rceil)$. From now on we fix a parameter $\tau\in(1,2]$. For any $k\in\mathbb{N},\boldsymbol{s}\in\mathbb{R}^{d}$, let
$\mathscr{S}_{k}(\boldsymbol{s}):=\big{|}\{(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\rrbracket\tau^{k-1},\tau^{k}\rrbracket\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(\tau^{k})\}\big{|}.$
For $N\geq 1$, letting $n\in\mathbb{N}$ such that $\tau^{n}\leq N<\tau^{n+1}$, and using that $\psi$ is non-increasing, we have
(8.6) $\mathscr{T}_{N}(\boldsymbol{s})\geq\mathscr{T}_{\tau^{n}}(\boldsymbol{s})\geq\sum_{k=1}^{n}\mathscr{S}_{k}(\boldsymbol{s}).$
We will obtain the lower bound for $\mathscr{T}_{N}(\boldsymbol{s})$ via an asymptotic lower bound for $\sum_{k=1}^{n}\mathscr{S}_{k}(\boldsymbol{s})$. More precisely, we will show
Proposition 8.4.
Under the assumptions of ˜8.3 and with $d\geq 2$, we have for every $\tau\in(1,2]$, for $\xi$-almost all $\boldsymbol{s}\in\mathbb{R}^{d}$, for every $\eta>0$, for all large enough $n$,
$\sum_{k=1}^{n}\mathscr{S}_{k}(\boldsymbol{s})\geq(1-\tau^{-1}-\eta)\sum_{k=1}^{n}\psi(\tau^{k})^{d}\tau^{k}.$
The lower bound in ˜8.3 follows directly:
Proof of lower bound in (8.4) using Proposition˜8.4.
Let $\varepsilon\in(0,1/2)$. As $\psi$ is non-increasing, we have for large $k$,
$(1-\tau^{-1})\psi(\tau^{k})^{d}\tau^{k+1}\geq(1-\varepsilon)(\lceil\tau^{k+1}\rceil-\lceil\tau^{k}\rceil)\psi(\tau^{k})^{d}\geq(1-\varepsilon)\sum_{q=\lceil\tau^{k}\rceil}^{\lceil\tau^{k+1}\rceil-1}\psi(q)^{d}$
Summing over $k$ and using the divergence $\sum_{q\in\mathbb{N}}\psi(q)^{d}=\infty$, we obtain that for every large enough $N$, and $n\geq 1$ such that $\tau^{n}\leq N<\tau^{n+1}$,
$(1-\tau^{-1})\tau\sum_{k=1}^{n}\psi(\tau^{k})^{d}\tau^{k}\geq(1-2\varepsilon)\sum_{q=1}^{N}\psi(q)^{d}$
Choose $\tau$ close enough to $1$ so that $\tau^{-1}\geq(1-\varepsilon)$. Choose $\eta>0$ with $\eta\lll_{\tau,\varepsilon}1$ so that $(1-\tau^{-1}-\eta)\geq(1-\varepsilon)(1-\tau^{-1})$. Using Proposition˜8.4, then (8.6), we obtain that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for large enough $N$,
$\mathscr{T}_{N}(\boldsymbol{s})\geq(1-2\varepsilon)^{3}\sum_{q=1}^{N}\psi(q)^{d}.$
This justifies the lower bound asymptotic $\liminf_{N\to+\infty}\frac{\mathscr{T}_{N}(\boldsymbol{s})}{\sum_{q=1}^{N}\psi(q)^{d}}\geq 1$. ∎
We now focus on the proof of Proposition˜8.4. For that, we need a dynamical interpretation of $\mathscr{S}_{k}(\boldsymbol{s})$. Recall $X=\operatorname{SL}_{d+1}(\mathbb{R})/\operatorname{SL}_{d+1}(\mathbb{Z})$ throughout the section. Given a measurable non-negative function $f:\mathbb{R}^{d+1}\rightarrow\mathbb{R}_{\geq 0}$, we denote by $\widetilde{f}:X\rightarrow[0,+\infty]$ its Siegel transform. It is given by: for $g\in G$,
$\widetilde{f}(gx_{0})=\sum_{v\in\mathbb{Z}^{d+1}\smallsetminus\{0\}}f(gv).$
We interpret $\mathscr{S}_{k}(\boldsymbol{s})$ dynamically by the mean of a Siegel transform. We fix some $\tau\in(1,2)$. For each $k\in\mathbb{N}$, define $r_{k},t_{k}\in\mathbb{R}_{>0}$ by the relations
$\psi(\tau^{k})=r_{k}t_{k}^{-\frac{1}{d+1}},\quad\tau^{k}=t_{k}^{\frac{d}{d+1}}r_{k},$
or equivalently,
(8.7) $\psi(\tau^{k})^{d}\tau^{k}=r_{k}^{d+1},\quad\tau^{k}\psi(\tau^{k})^{-1}=t_{k}.$
Consider the box
$R_{k}=[0,r_{k})^{d}\times(\tau^{-1}r_{k},r_{k}]\subseteq\mathbb{R}^{d+1}.$
By direct computation, we have
(8.8) $\mathscr{S}_{k}(\boldsymbol{s})=\widetilde{\mathbbm{1}}_{R_{k}}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}.$
Let $\gamma_{1}\in(0,1/2)$ be a small parameter to be specified later. We partition the subscripts $k$’s into two families : $\operatorname{K_{big}}:=\{k\geq 3\,:\,r_{k}>t_{k}^{\gamma_{1}}\}$, and $\operatorname{K_{small}}:=\mathbb{N}_{\geq 3}\smallsetminus\operatorname{K_{big}}$. Given $n\geq 1$, we set $\operatorname{K_{big}}(n)=\operatorname{K_{big}}\cap\llbracket 1,n\rrbracket$, and $\operatorname{K_{small}}(n)=\operatorname{K_{small}}\cap\llbracket 1,n\rrbracket$. We will establish the lower asymptotics required by Proposition˜8.4 for the sums $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$ and $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$ separately.
[…]
We first show that a typical geodesic trajectory sampled by $\xi$ has at most a very slow escape to infinity along the sequence of times $(t_{k})_{k\geq 1}$.
Lemma 8.5.
For $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for all sufficiently large $k\geq 1$ (depending on $\boldsymbol{s}$), we have
$\operatorname{dist}(a(t_{k})u(\boldsymbol{s})x_{0},x_{0})\ll\log\log t_{k}.$
Proof.
For $r>1$, let $f_{r}:X\rightarrow[0,1]$ be a smooth function such that ${\mathcal{S}}_{\infty,l}(f_{r})\ll 1$ and ${\mathbbm{1}}_{\operatorname{dist}(\cdot,x_{0})\geq r}\leq f_{r}\leq{\mathbbm{1}}_{\operatorname{dist}(\cdot,x_{0})\geq r/2}$. Then by Lemma˜4.5,
$m_{X}(f_{r})\leq m_{X}\{\operatorname{dist}(\cdot,x_{0})\geq r/2\}\leq Me^{-r/M}$
for some $M=M(d)>0$. Applying effective single equidistribution (8.1) with test function $f_{r}$ at time $t>1$, we get
$\xi\{\boldsymbol{s}\,:\,\operatorname{dist}(a(t)u(\boldsymbol{s})x_{0},x_{0})\geq r\}\leq\int_{\mathbb{R}^{d}}f_{r}(a(r)u(\boldsymbol{s})x_{0})\,\mathrm{d}\xi(\boldsymbol{s})\ll Me^{-r/M}+t^{-c}.$
Recalling $t_{k}\gg\tau^{k}$ where $\tau>1$, the right hand side has converging series over $(t,r)\in\{(t_{k},(M+1)\log\log t_{k})\,:\,k\geq 1\}$, and the claim follows by the Borel-Cantelli Lemma. ∎
The next lemma expresses that the counting measure on a covolume $1$ lattice of $\mathbb{R}^{d+1}$ is a good volume estimate for a box in $\mathbb{R}^{d+1}$, provided the box has large enough sidelength depending on the distorsion of the lattice.
Lemma 8.6.
Let $R\subseteq\mathbb{R}^{d+1}$ be a subset of the form
$R=v+\prod_{i=1}^{d+1}[0,T_{i}]$
where $v\in\mathbb{R}^{d+1}$ and $(T_{i})_{i=1}^{d+1}\in\mathbb{R}_{>0}^{d+1}$. Let $g\in G$ with $\lVert g\rVert\leq\min_{1\leq i\leq d+1}\frac{T_{i}}{\sqrt{d+1}}$. Then
$\left\lvert\lvert g\mathbb{Z}^{d+1}\cap R\rvert-\operatorname{Leb}(R)\right\rvert\leq 2^{d+1}\sqrt{d+1}\max_{1\leq i\leq d+1}\frac{\lVert g\rVert}{T_{i}}\operatorname{Leb}(R).$
Proof.
Set $Q:=g(-\frac{1}{2},\frac{1}{2})^{d+1}$. The symmetric difference of $(g\mathbb{Z}^{d+1}\cap R)+Q$ and $R$ is contained in $\partial R+Q$. Taking the volume, we obtain
$\left\lvert\lvert g\mathbb{Z}^{d+1}\cap R\rvert-\operatorname{Leb}(R)\right\rvert\leq\operatorname{Leb}(\partial R+Q).$
Note that $Q\subseteq B^{\mathbb{R}^{d+1}}_{\rho}$ where $\rho:=\frac{\sqrt{d+1}}{2}\lVert g\rVert\leq\frac{1}{2}\min_{i}T_{i}$, in particular
$\partial R+Q\subseteq v\,+\,\prod_{i=1}^{d+1}[-\rho,T_{i}+\rho]\smallsetminus\prod_{i=1}^{d+1}(\rho,T_{i}-\rho).$
It follows that
$\frac{\operatorname{Leb}(\partial R+Q)}{\operatorname{Leb}(R)}$ $\leq\prod_{i=1}^{d+1}\bigl{(}1+\frac{2\rho}{T_{i}}\bigr{)}-\prod_{i=1}^{d+1}\bigl{(}1-\frac{2\rho}{T_{i}}\bigr{)}$
$\leq\bigl{(}1+\max_{i}\frac{2\rho}{T_{i}}\bigr{)}^{d+1}-\bigl{(}1-\max_{i}\frac{2\rho}{T_{i}}\bigr{)}^{d+1}$
$\leq 2^{d+2}\max_{i}\frac{\rho}{T_{i}}$
where the last bound is obtained by expanding the power $d+1$ and using $\frac{2\rho}{T_{i}}\leq 1$. This yields the desired estimate. ∎
We infer from Lemmas 8.5, 8.6 the asymptotic lower bound for $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$.
Lemma 8.7.
Assume $\sum_{q\in\operatorname{K_{big}}}\psi(\tau^{k})^{d}\tau^{k}=+\infty$. Then for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for every $\eta>0$, for large enough $n$,
$\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}_{k}(\boldsymbol{s})\geq(1-\tau^{-1}-\eta)\sum_{k\in\operatorname{K_{big}}(n)}\psi(\tau^{k})^{d}\tau^{k}$
Proof.
Recall that for every $\boldsymbol{s}\in\mathbb{R}^{d}$, $k\geq 3$, we have $\mathscr{S}_{k}(\boldsymbol{s})=\widetilde{{\mathbbm{1}}}_{R_{k}}(a(t_{k})u(\boldsymbol{s})x_{0})$. Assuming $k\in\operatorname{K_{big}}$, we have that $R_{k}$ is a box of minimal sidelength $(1-\tau^{-1})r_{k}\geq(1-\tau^{-1})t^{\gamma_{1}}_{k}$. Moreover Lemma˜8.5 guarantees that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for large enough $k$, say $k\geq k_{\boldsymbol{s}}$, we have $a(t_{k})u(\boldsymbol{s})\in g\operatorname{SL}_{d+1}(\mathbb{Z})$ where $\|g\|\leq(\log t_{k})^{O(1)}\leq t_{k}^{\gamma_{1}/100}$. Invoking Lemma˜8.6, we obtain in this context
$\widetilde{{\mathbbm{1}}}_{R_{k}}(a(t_{k})u(\boldsymbol{s})x_{0})\geq(1-t^{-\gamma_{1}/4}_{k})\operatorname{Leb}(R_{k}).$
Recalling from (8.7) that $\operatorname{Leb}(R_{k})=(1-\tau^{-1})\psi(\tau^{k})^{d}\tau^{k}$, we deduce
$\sum_{\begin{subarray}{c}k\in\operatorname{K_{big}}(n)\\ k\geq k_{\boldsymbol{s}}\end{subarray}}\mathscr{S}_{k}(\boldsymbol{s})\,\geq\,(1-\tau^{-1})\sum_{\begin{subarray}{c}k\in\operatorname{K_{big}}(n)\\ k\geq k_{\boldsymbol{s}}\end{subarray}}(1-t^{-\gamma_{1}/4}_{k})\psi(\tau^{k})^{d}\tau^{k}$
and the lemma follows using the right hand side is divergent by hypothesis. ∎
Lower asymptotic over $\operatorname{K_{small}}$. We now establish an asymptotic lower bound for the partial sums $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$. Let $\varepsilon,\gamma_{2}\in(0,1/2)$ be small parameters to be specified below. For $k\in\operatorname{K_{small}}$, set
$R^{-}_{k}=[\varepsilon r_{k},(1-\varepsilon)r_{k})^{d}\times((\tau^{-1}+\varepsilon)r_{k},(1-\varepsilon)r_{k}]$
the rectangle obtained from $R_{k}$ by shrinking sides via $\varepsilon r_{k}$. Let $\chi_{k}:X\rightarrow\{0,1\}$ be the truncation function given by
(8.9) $\chi_{k}(x)=\begin{cases}1,\quad\text{if }\operatorname{inj}(x)\geq t_{k}^{-\gamma_{2}},\\ 0,\quad\text{ otherwise.}\end{cases}$
Let $\theta:B_{\varepsilon/10}\rightarrow\mathbb{R}_{\geq 0}$ be a smooth bump function such that $m_{G}(\theta)=1$ and ${\mathcal{S}}_{\infty,l}(\theta)\leq\varepsilon^{-D}$ where $D=D(G,l)>0$. Set
$\varphi_{k}=\theta*(\chi_{k}\widetilde{{\mathbbm{1}}}_{R_{k}^{-}}).$
We view $\varphi_{k}$ as a bounded and smooth approximation of $\widetilde{{\mathbbm{1}}}_{R_{k}^{-}}$. Note that every $g\in B_{\varepsilon/10}$ satisfies $gR_{k}^{-}\subseteq R_{k}$, whence $\varphi_{k}\leq\widetilde{{\mathbbm{1}}}_{R_{k}}$, and in particular
$\mathscr{S}_{k}(\boldsymbol{s})\geq\varphi_{k}(a(t_{k})u(\boldsymbol{s})x_{0}).$
Therefore, we will focus on establishing a lower bound for the partial sums of terms $\varphi_{k}(a(t_{k})u(\boldsymbol{s})x_{0})$ as $k$ runs along $\operatorname{K_{small}}$.
[…]
We first recall well-known moment estimates for the Siegel transform of characteristic functions of bounded sets, see [50, pages 2-3]. We emphasize here that we work under the assumption $d\geq 2$ (otherwise (8.11) does not hold, see §8.3).
Fact 8.8 (Moments of Siegel transforms [50]).
Let $R\subseteq\mathbb{R}^{d+1}$ be a bounded measurable subset. Then
(8.10) $\int_{X}\widetilde{{\mathbbm{1}}}_{R}\,\mathrm{d}m_{X}=\operatorname{Leb}(R)$
(8.11) $\int_{X}(\widetilde{{\mathbbm{1}}}_{R})^{2}\,\mathrm{d}m_{X}=\operatorname{Leb}(R)^{2}+O(\operatorname{Leb}(R)).$
We also record that convolution with a (signed) bump function does not increase the $L^{2}$-norm.
Lemma 8.9.
For every measurable functions $\iota\in L^{1}(G)$, $F\in L^{2}(X)$, we have $\|\iota*F\|_{L^{2}}\leq\|\iota\|_{L^{1}}\|F\|_{L^{2}}.$
Proof.
Using the triangle inequality for the $L^{2}$-norm, and the fact that $\|g_{\star}F\|_{L^{2}}=\|F\|_{L^{2}}$, we have $\|\iota*F\|_{L^{2}}=\|\int_{G}\iota(g)g_{\star}F\,\mathrm{d}m_{G}(g)\|_{L^{2}}\leq\int_{G}\lvert\iota(g)\rvert\,\|g_{\star}F\|_{L^{2}}\,\mathrm{d}m_{G}(g)=\|\iota\|_{L^{1}}\|F\|_{L^{2}}$. ∎
We deduce from ˜8.8 and Lemma˜8.9 several moment estimates for the functions $\varphi_{k}$.
Lemma 8.10.
If $\gamma_{1}\lll\gamma_{2}$, then for some $M=M(d)>1$, every $k\in\operatorname{K_{small}}$, we have
(8.12) $m_{X}(\varphi_{k})=\operatorname{Leb}(R_{k}^{-})-O(t_{k}^{-\gamma_{2}/M}),$
(8.13) ${\mathcal{S}}_{\infty,l}(\varphi_{k})\ll t_{k}^{M\gamma_{2}},$
(8.14) ${\mathcal{S}}_{2,l}(\varphi_{k})\ll m_{X}(\varphi_{k})+\sqrt{m_{X}(\varphi_{k})}+t_{k}^{-\gamma_{2}/M}.\$
Proof.
Let us prove (8.12). Note first $m_{X}(\varphi_{k})=m_{X}(\chi_{k}\widetilde{{\mathbbm{1}}}_{R_{k}^{-}})$. Applying (8.10), followed by the Cauchy-Schwarz inequality, Lemma˜4.5 and (8.11), we find
$0\leq\operatorname{Leb}(R_{k}^{-})-m_{X}(\varphi_{k})$ $=\int_{X}{\mathbbm{1}}_{\operatorname{inj}(x)<t^{-\gamma_{2}}_{k}}\widetilde{{\mathbbm{1}}}_{R_{k}^{-}}(x)\,\,\mathrm{d}m_{X}(x)$
$=\sqrt{m_{X}\{\operatorname{inj}<t^{-\gamma_{2}}_{k}\}}\,\|\widetilde{{\mathbbm{1}}}_{R_{k}^{-}}\|_{L^{2}}$
$\ll t_{k}^{-\gamma_{2}/M}\max(\operatorname{Leb}(R_{k}^{-}),\sqrt{\operatorname{Leb}(R_{k}^{-})})$
for some $M=M(d)>1$. But $\operatorname{Leb}(R_{k}^{-})\leq(1-\tau^{-1})r_{k}^{d+1}\ll t_{k}^{O(\gamma_{1})}$ since $k\in\operatorname{K_{small}}$. Thus upon letting $\gamma_{1}\lll\gamma_{2}$, the right-hand side can be bounded by $t_{k}^{-\gamma_{2}/(2M)}$, validating (8.12).
We now deal with (8.13). Note that ${\mathcal{S}}_{\infty,l}(\varphi_{k})\ll{\mathcal{S}}_{\infty,l}(\theta)\|\chi_{k}\widetilde{1}_{R_{k}^{-}}\|_{L^{\infty}}$. By construction, we have ${\mathcal{S}}_{\infty,l}(\theta)\ll 1$. On the other hand, $\|\chi_{k}\widetilde{{\mathbbm{1}}}_{R_{k}^{-}}\|_{L^{\infty}}=\sup_{x\,:\,\operatorname{inj}(x)\geq t_{k}^{-\gamma_{2}}}\widetilde{{\mathbbm{1}}}_{R^{-}_{k}}(x)$. For such $x$, Equation (4.3) allows to write $x=gx_{0}$ where $g\in G$ satisfies $\lVert g^{-1}\rVert\ll t_{k}^{M\gamma_{2}}$ for some $M=M(d)>1$. Then $\widetilde{{\mathbbm{1}}}_{R^{-}_{k}}(x)=\lvert\mathbb{Z}^{d+1}\cap g^{-1}R^{-}_{k}\rvert$ with $g^{-1}R^{-}_{k}$ contained in a ball of radius $O(\lVert g^{-1}\rVert r_{k})=O(t_{k}^{M\gamma_{2}}r_{k})$. Recalling the assumption $k\in\operatorname{K_{small}}$, this implies $\widetilde{{\mathbbm{1}}}_{R^{-}_{k}}(x)\ll t_{k}^{(d+1)M(\gamma_{2}+\gamma_{1})}$. This shows (8.13).
Finally, we check (8.14). By Lemma˜8.9 followed by (8.11), we have
${\mathcal{S}}_{2,l}(\varphi_{k})\leq{\mathcal{S}}_{1,l}(\theta)\|\chi_{k}\widetilde{{\mathbbm{1}}}_{R_{k}^{-}}\|_{L^{2}}\ll\|\widetilde{{\mathbbm{1}}}_{R_{k}^{-}}\|_{L^{2}}\ll\operatorname{Leb}(R_{k}^{-})+\sqrt{\operatorname{Leb}(R_{k}^{-})}.$
Now (8.14) follows from (8.12). ∎
We consider $(\mathbb{R}^{d},\xi)$ as a probability space. Expectation $\mathbb{E}[\,\cdot\,]$ refers implicitely to this probability space. For every $k\in\operatorname{K_{small}}$, we introduce the random variable
$Y_{k}:\mathbb{R}^{d}\to\mathbb{R},\quad\boldsymbol{s}\mapsto\varphi_{k}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}.$
We write
$y_{k}=m_{X}(\varphi_{k})\in\mathbb{R}_{\geq 0}$
and set $Z_{k}=Y_{k}-y_{k}$ the (quasi-recentered) companion of $Y_{k}$. The next lemma bounds the second moment of $Z_{k}$ by $y_{k}$, provided $y_{k}$ is not too small. It relies on single effective equidistribution of expanding translates of $u(\boldsymbol{s})x_{0}\,\mathrm{d}\xi(\boldsymbol{s})$.
Lemma 8.11.
Assume $\gamma_{1}\lll\gamma_{2}\lll 1$. Then for every $k\in\operatorname{K_{small}}$ such that $y_{k}\geq 1$, we have
$\mathbb{E}[Z_{k}^{2}]\ll y_{k}.$
Proof.
By effective single equidistribution of expanding translates (8.1), we have
$\mathbb{E}[Z_{k}^{2}]$ $=\int_{\mathbb{R}^{d}}(\varphi_{k}(a(t_{k})u(\boldsymbol{s})x_{0})-y_{k})^{2}\,\mathrm{d}\xi(\boldsymbol{s})$
(8.15) $=\int_{X}(\varphi_{k}(x)-y_{k})^{2}\,\mathrm{d}m_{X}(x)+O({\mathcal{S}}_{\infty,l}([\varphi_{k}-y_{k}]^{2})t_{k}^{-c}).$
Let us bound the error term in (8.15). By (8.13), we have
${\mathcal{S}}_{\infty,l}([\varphi_{k}-y_{k}]^{2})\ll{\mathcal{S}}_{\infty,l}(\varphi_{k})^{2}+y^{2}_{k}\ll t_{k}^{2M\gamma_{2}}+y^{2}_{k}.$
Taking $\gamma_{2}\lll 1$, we have $t_{k}^{2M\gamma_{2}-c}\ll 1$. Taking $\gamma_{1}\lll 1$, and observing $y_{k}^{2}\leq t_{k}^{2(d+1)\gamma_{1}}$ by (8.12) and definition of $\operatorname{K_{small}}$, we also find $y^{2}_{k}t_{k}^{-c}\ll 1$. Therefore, the error term in (8.15) is bounded by $O(1)$.
We now estimate the main term of (8.15). By expanding the square, then using Lemma˜8.9 and $m_{G}(\theta)=1$, we see that
$\int_{X}(\varphi_{k}(x)-y_{k})^{2}\,\mathrm{d}m_{X}(x)=\lVert\varphi_{k}\rVert_{L^{2}}^{2}-y_{k}^{2}\leq\lVert\chi_{k}\tilde{{\mathbbm{1}}}_{R^{-}_{k}}\rVert_{L^{2}}^{2}-y_{k}^{2}\leq\lVert\tilde{{\mathbbm{1}}}_{R^{-}_{k}}\rVert_{L^{2}}^{2}-y_{k}^{2}.$
Using (8.11), (8.12) and $y_{k}\geq 1$, the main term is bounded by $O(y_{k})$. The result follows. ∎
From the effective double equidistribution hypothesis on $\xi$, we deduce an upper bound on the second moment of a sum of $Z_{k}$’s where $k\in\operatorname{K_{small}}$.
Proposition 8.12.
Assume $\gamma_{1}\lll\gamma_{2}\lll 1$ and
(8.16) $\psi(q)\geq q^{-1/d}\log^{-2/d}(q),\quad\forall q\in\operatorname{K_{small}}.$
Then for every finite subset $J\subseteq\operatorname{K_{small}}$ with $\inf J\ggg_{\gamma_{2}}1$, we have
$\mathbb{E}[(\sum_{j\in J}Z_{j})^{2}]\ll\left(1+\sum_{j\in J}y_{j}\right)^{3/2}.$
Proof.
We use the shorthand
$Y_{J}:=\sum_{k\in J}Y_{k},\quad\quad\quad\quad y_{J}:=\sum_{k\in J}y_{k},\quad\quad\quad\quad Z_{J}=Y_{J}-y_{J}.$
Set $J_{1}:=\{\,j\in J\,:\,y_{j}<1\,\}$, write $n:=\lvert J\smallsetminus J_{1}\rvert$. Further partition $J$ into $J=J_{1}\sqcup J_{2}\sqcup J_{3}$ where $J_{2},J_{3}$ are respectively determined by the condition $y_{j}\in[1,n^{2})$, and $y_{j}\in[n^{2},+\infty)$. Using the inequalities $(a+b)^{2}\leq 2(a^{2}+b^{2})$ and $a^{3/2}+b^{3/2}\leq(a+b)^{3/2}$ valid for all $a,b\in\mathbb{R}_{\geq 0}$, we just need to check the upper bound for the sum over each $J_{i}$ independently.
Case of $J_{1}$.
By definition, for all $j\leq k\in\operatorname{K_{small}}$,
(8.17) $\mathbb{E}[Z_{j}Z_{k}]=\mathbb{E}[Y_{j}Y_{k}]-y_{j}y_{k}-\mathbb{E}[Z_{j}]y_{k}-y_{j}\mathbb{E}[Z_{k}].$
By double equidistribution (8.3), we have
$\lvert\mathbb{E}[Y_{j}Y_{k}]-y_{j}y_{k}\rvert\ll{\mathcal{S}}_{2,l}(\varphi_{j}){\mathcal{S}}_{2,l}(\varphi_{k})t_{j}^{c}t_{k}^{-c}+{\mathcal{S}}_{\infty,l}(\varphi_{j})y_{k}t_{j}^{-c}+{\mathcal{S}}_{\infty,l}(\varphi_{j}){\mathcal{S}}_{\infty,l}(\varphi_{k})t_{k}^{-c}.$
Assume $j,k\in J_{1}$ so that $y_{j},y_{k}<1$. Using (8.13), (8.14), the above becomes
$\lvert\mathbb{E}[Y_{j}Y_{k}]-y_{j}y_{k}\rvert$ $\ll(\sqrt{y_{j}}+t_{j}^{-\gamma_{2}/M})(\sqrt{y_{k}}+t_{k}^{-\gamma_{2}/M})t_{j}^{c}t_{k}^{-c}+y_{k}t_{j}^{-c+M\gamma_{2}}+t_{j}^{M\gamma_{2}}t_{k}^{-c+M\gamma_{2}}$
(8.18) $\ll\sqrt{y_{j}y_{k}}t_{j}^{c}t_{k}^{-c}+y_{k}t_{j}^{-c/2}+t_{k}^{-c/2}+t_{j}^{c-\gamma_{2}/M}t_{k}^{-c}$
where the second inequality assumes $\gamma_{2}\leq c/(4M)$.
On the other hand, by single equidistribution (8.1) and the norm control (8.13), we have for $j\in J_{1}$,
(8.19) $\lvert\mathbb{E}[Z_{j}]\rvert\ll t_{j}^{-c+M\gamma_{2}}\ll t_{j}^{-c/2}.$
By expanding the square, using the above bounds (8.17), (8.18), (8.19), and recalling from (8.7) that $t_{j}\gg\tau^{j}$ and $t_{k}/t_{j}\geq\tau^{k-j}$ for $j\leq k$, we deduce
$\mathbb{E}[Z_{J_{1}}^{2}]\ll\sum\nolimits_{j,k\in J_{1},j\leq k}\left(\sqrt{y_{j}y_{k}}\tau^{-c(k-j)}+y_{k}\tau^{-cj/2}+\tau^{-ck/2}+\tau^{-c(k-j)-j\gamma_{2}/M}\right).$
Using $\sqrt{y_{j}y_{k}}\leq y_{j}+y_{k}$ and the convergence $\sum_{n=0}^{\infty}\tau^{-cn}<+\infty$, the first sum satisfies $\sum\nolimits_{j,k\in J_{1},j\leq k}\sqrt{y_{j}y_{k}}\tau^{-c(k-j)}\ll y_{J_{1}}$. The convergence $\sum_{n=0}^{\infty}\tau^{-cn}<+\infty$ bounds similarly the second sum. To bound the third sum, note that combining (8.7) with our assumption (8.16), we have
$\tau^{-cj}\ll(j\log\tau)^{-2/d}\leq r_{j}^{d+1}\ll y_{j},$
where the last inequality relies on the assumption $\inf J\ggg_{\gamma_{2}}1$ and (8.12). Hence $\tau^{-ck}\ll y_{j}\tau^{-c(k-j)}$, so $\sum\nolimits_{j,k\in J_{1},j\leq k}\tau^{-ck}\ll y_{J_{1}}$ as for the first sum. The fourth sum can be handled similarly to the third one. As $\sum_{J_{1}}y_{j}\leq(1+y_{J_{1}})^{3/2}$, we have justified the upper bound for $J_{1}$.
Case of $J_{2}$.
Set $m:=|J_{2}|$. We start with the case where $m$ is very small compared to $n=|J_{2}\sqcup J_{3}|$, more precisely we assume $m^{2}\leq n$. In this scenario, we have by the Cauchy-Schwarz inequality and Lemma˜8.11,
$\mathbb{E}\bigl{[}Z_{J_{2}}^{2}\bigr{]}\leq m\sum_{j\in J_{2}}\mathbb{E}[Z_{j}^{2}]\ll m\sum_{j\in J_{2}}y_{j}\leq m^{2}n^{2}\leq n^{3}\ll y_{J_{2}\sqcup J_{3}}^{3/2},$
whence the desired bound. Assume now $m^{2}>n$. Decompose $J_{2}$ into $J_{2}=J^{\prime}_{2}\sqcup J_{2}^{\prime\prime}$ according to whether $j\geq\sqrt{n}$ or not. The preceding argument gives
$\mathbb{E}\bigl{[}Z_{J^{\prime\prime}_{2}}^{2}\bigr{]}\ll y_{J_{2}\sqcup J_{3}}^{3/2}.$
We now focus on $J^{\prime}_{2}$. Note that
(8.20) $\mathbb{E}\bigl{[}Z_{J^{\prime}_{2}}^{2}\bigr{]}$ $=\sum_{|j-k|<\sqrt{m}}\mathbb{E}[Z_{j}Z_{k}]+\sum_{|j-k|\geq\sqrt{m}}\mathbb{E}[Z_{j}Z_{k}]$
(8.21) $\ll\sqrt{m}\sum_{j\in J^{\prime}_{2}}\mathbb{E}[Z_{j}^{2}]+\Bigl{\lvert}\sum_{|j-k|\geq\sqrt{m}}\mathbb{E}[Z_{j}Z_{k}]\Bigr{\rvert}$
where the second inequality uses the trivial bound $\mathbb{E}[Z_{j}Z_{k}]\leq\mathbb{E}[Z_{j}^{2}+Z_{k}^{2}]$ and the observation that each subscript $j$ in the first sum of (8.20) appears at most $O(\sqrt{m})$ many times.
For subscripts $j\leq k\in J_{2}^{\prime}$ such that $|j-k|\geq\sqrt{m}$, we have by (8.17), double equidistribution (8.3), and Lemma˜8.10, that
$\mathbb{E}[Z_{j}Z_{k}]\ll y_{j}y_{k}(t_{j}^{c}t_{k}^{-c}+t_{j}^{-c/2}+t_{k}^{-c/2})\leq n^{4}\tau^{-c\sqrt{m}/2}$
where the second inequality relies on the definition of $J^{\prime}_{2}$. Plugging this bound and Lemma˜8.11 into (8.21), we obtain
$\mathbb{E}\bigl{[}Z_{J^{\prime}_{2}}^{2}\bigr{]}\ll\sqrt{m}y_{J^{\prime}_{2}}+\underbrace{m^{2}n^{4}\tau^{-c\sqrt{m}/2}}_{O(1)}\,\ll\,y_{J_{2}}^{3/2}.$
Case of $J_{3}$.
We finally deal with $J_{3}$. Applying the Cauchy-Schwarz inequality then Lemma˜8.11, we obtain
$\mathbb{E}\bigl{[}Z_{J_{3}}^{2}\bigr{]}\leq|J_{3}|\sum_{j\in J_{3}}\mathbb{E}[Z_{j}^{2}]\ll|J_{3}|\sum_{j\in J_{3}}y_{j}\leq y_{J_{3}}^{3/2}.$
This concludes the proof. ∎
We also need the next lemma, which is a variant of [25, Lemma 1.5]. It converts a variance control as in Proposition˜8.12 into an asymptotic estimate.
Lemma 8.13.
Let $(Y_{j})_{j\geq 1}$ be a sequence of non-negative real random variables. Let $(y_{j})_{j\geq 1},(y^{\prime}_{j})_{j\geq 1}\in\mathbb{R}_{\geq 0}^{\mathbb{N}}$ be sequences of non-negative real numbers. Set $Z_{j}=Y_{j}-y_{j}$. Assume $y_{j}\leq y_{j}^{\prime}$ for all $j$, as well as $\sum_{j=1}^{\infty}y_{j}=+\infty$, and that for some $C_{1}\geq 1$, for all $n\geq m\geq C_{1}$,
(8.22) $\mathbb{E}\Bigl{[}\bigl{(}\sum_{j=m}^{n}Z_{j}\bigr{)}^{2}\Bigr{]}\leq C_{1}\bigl{(}1+\sum_{j=m}^{n}y^{\prime}_{k}\bigr{)}^{3/2}.$
Then almost surely, for large enough $n$, we have
$\Bigl{\lvert}\sum_{k=1}^{n}Z_{k}\Bigr{\rvert}\leq\Bigl{(}\sum_{k=1}^{n}y^{\prime}_{k}\Bigr{)}^{4/5}$
Proof.
For an interval $J\subseteq\mathbb{R}_{>0}$, we use the notation $Z_{J}=\sum_{j\in\mathbb{N}_{\geq 1}\cap J}Z_{j}$ and define similarly $y_{J}$, $y^{\prime}_{J}$. We prove the following slightly stronger statement : there is an almost-surely finite random variable $C_{2}$ such that for all $N\geq C_{2}$, we have
(8.23) $\lvert Z_{(0,N]}\rvert\leq(\log(y^{\prime}_{(0,N]}+2))^{2}(y^{\prime}_{(0,N]}+2)^{3/4}+C_{2}.$
For this, up to throwing away a finite number of terms, we may assume (8.22) holds for all $n\geq m\geq 1$.
Under this assumption, we have the following.
Lemma 8.14.
Let $0=N_{0}<N_{1}<N_{2}<\dotsb$ be an increasing sequence of integers such that
(8.24) $\forall i\geq 0,\quad y^{\prime}_{(N_{i},N_{i+1}]}\geq 1.$
Then almost-surely, for sufficiently large $i$,
(8.25) $Z_{(0,N_{i}]}^{2}\leq(\log y^{\prime}_{(0,N_{i}]})^{4}(y^{\prime}_{(0,N_{i}]})^{3/2}.$
Proof of Lemma˜8.14.
Denote by $\mathscr{D}$} the set of integers $T\geq 2$ such that the associated dyadic interval $(2^{T-1},2^{T}]$ meets the collection $(y^{\prime}_{(0,N_{i}]})_{i\geq 1}$. Define a sequence of integers $(M_{T})_{T\in{\mathscr{D}}}$ by
$M_{T}=\max\{N_{i}\,:\,y^{\prime}_{(0,N_{i}]}\in(2^{T-1},2^{T}]\}.$
For each $T\in{\mathscr{D}}$, consider the following collection of intervals
${\mathcal{K}}_{T}=\Bigl{\{}\,(N_{r2^{t}},N_{(r+1)2^{t}}]\,:\,r\geq 1,\,t\geq 0,\text{ and }N_{(r+1)2^{t}}\leq M_{T}\,\Bigr{\}}.$
By assumption, $y^{\prime}_{(0,N_{i}]}\geq i$ for every $i\geq 0$. It follows that $M_{T}\leq N_{2^{T}}$. Therefore, every integer in $\llbracket 1,M_{T}\rrbracket$ is contained in at most $T+1$ intervals of ${\mathcal{K}}_{T}$. Applying the assumption (8.22) and the inequality $y^{\prime}_{(N_{i},N_{i+1}]}\geq 1$, followed by the relation $a^{3/2}+b^{3/2}\leq(a+b)^{3/2}$ for all $a,b\in\mathbb{R}_{\geq 0}$, we deduce that
$\sum_{I\in{\mathcal{K}}_{T}}\mathbb{E}[Z_{I}^{2}]\leq{C_{1}}\left(2\sum_{I\in{\mathcal{K}}_{T}}y^{\prime}_{I}\right)^{3/2}\leq{C_{1}}\left(2(T+1)y^{\prime}_{(0,M_{T}]}\right)^{3/2}\leq 2^{3}{C_{1}}\left(T2^{T}\right)^{3/2}.$
By Markov’s inequality,
$\mathbb{P}\Bigl{[}\sum\nolimits_{I\in{\mathcal{K}}_{T}}Z_{I}^{2}>2^{-10}T^{3}2^{3T/2}\Bigr{]}\leq 2^{13}{C_{1}}T^{-3/2}.$
The latter being summable over $T\in{\mathscr{D}}$, we can use the Borel-Cantelli lemma to deduce that almost surely, for large enough $T$, we have
(8.26) $\sum_{I\in{\mathcal{K}}_{T}}Z_{I}^{2}\leq 2^{-10}T^{3}2^{3T/2}.$
Let $i\geq 1$. Assume $i$ large enough so that the unique element $T\in{\mathscr{D}}$ such that $2^{T-1}<y^{\prime}_{(0,N_{i}]}\leq 2^{T}$ satisfies (8.26) as well. By considering $i$ in base $2$ and using $N_{i}\leq M_{T}\leq N_{2^{T}}$, we may cover $(0,N_{i}]$ with at most $T$ non-overlapping intervals from ${\mathcal{K}}_{T}$. Let ${\mathcal{K}}_{T,i}$ be such a collection of intervals. Then by the Cauchy-Schwarz inequality and (8.26),
$Z_{(0,N_{i}]}^{2}=\Bigl{(}\sum_{I\in{\mathcal{K}}_{T,i}}Z_{I}\Bigr{)}^{2}\leq\lvert{\mathcal{K}}_{T,i}\rvert\sum_{I\in{\mathcal{K}}_{T,i}}Z_{I}^{2}\leq 2^{-10}T^{4}2^{3T/2}.$
We obtain the desired bound using $2^{T-1}<y^{\prime}_{(0,N_{i}]}$. ∎
To show (8.23), we provide lower and upper bounds for $Z_{(0,N]}$. Let $\mathscr{M}$} be the set of integers $m\geq 2$ such that the interval (m1,m](m-1,m] meets the collection $(y^{\prime}_{(0,j]})_{j\geq 1}$. Consider ${\mathscr{N}}=\{n_{1}<n_{2}<\dotsb\}\subseteq{\mathscr{M}}$ a subset satisfying $\inf_{i\neq j}|n_{i}-n_{j}|\geq 2$ and maximal conditionally to this property.
To obtain the lower bound, we set for $i\geq 1$,
$N_{i}:=\min\{j\geq 1\,:\,y^{\prime}_{(0,j]}\in(n_{i}-1,n_{i}]\}.$
The advantage of using $\mathscr{N}$} and not $\mathscr{M}$} to define $N_{i}$ is that we can guarantee (8.24). Thus we can apply Lemma˜8.14 to the sequence $(N_{i})$. We obtain that almost surely, if $N\geq 1$ is sufficiently large, then the unique $i\geq 1$ such that $N\in[N_{i},N_{i+1})$ satisfies (8.25). Recalling that the $Y_{j}$’s are almost-surely non-negative, we obtain that
$Z_{(0,N]}\geq Z_{(0,N_{i}]}-y_{(N_{i},N]}\geq-(\log y^{\prime}_{(0,N_{i}]})^{2}(y^{\prime}_{(0,N_{i}]})^{3/4}-y^{\prime}_{(N_{i},N]}.$
and the desired lower bound follows, noting that by construction, we have $0\leq y^{\prime}_{(N_{i},N]}\leq 2$.
The upper bound can be handled similarly, but for that we need to modify the sequence $(N_{i})_{i}$ into a certain $(N^{\prime}_{i})_{i}$, guaranteeing that when $N$ ranges within $(N^{\prime}_{i-1},N^{\prime}_{i}]$, the value of $y^{\prime}_{(0,N]}$ does not vary much. More precisely, we replace $N_{i}$ with
$N^{\prime}_{i}=\max\{j\geq 1\,:\,y^{\prime}_{(0,j]}\in(n_{i}-1,n_{i}]\}.$
Applying Lemma˜8.14 with $(N^{\prime}_{i})_{i}$, we have that almost-surely, for large enough $N$, for $i$ such that $(N^{\prime}_{i-1},N^{\prime}_{i}]$,
$Z_{(0,N]}\leq Z_{(0,N^{\prime}_{i}]}+y_{(N,N^{\prime}_{i}]}\leq(\log y^{\prime}_{(0,N^{\prime}_{i}]})^{2}(y^{\prime}_{(0,N^{\prime}_{i}]})^{3/4}+y_{(N,N^{\prime}_{i}]}.$
The desired lower bound follows, noting that by construction, we have $0\leq y^{\prime}_{(N,N_{i}^{\prime}]}\leq 2$. This finishes the proof of (8.23). ∎
Combining Proposition˜8.12 and Lemma˜8.13, we obtain the following counting estimate for parameters in $\operatorname{K_{small}}$. Given $n\geq 1$, we recall $\operatorname{K_{small}}(n)=\operatorname{K_{small}}\cap\llbracket 1,n\rrbracket$.
Corollary 8.15.
Assume $\sum_{k\in\operatorname{K_{small}}}\psi(\tau^{k})^{d}\tau^{k}=+\infty$ and (8.16), as well as $\gamma_{1}\lll 1$. Then for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}^{d}$, for large enough $n$,
(8.27) $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})\geq(1-\tau^{-1}-C(d)\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})^{d}\tau^{k}$
where $C(d)>0$ is a constant depending on $d$ only.
Proof.
By the convergent case of the Khintchine dichotomy, replacing $\psi(q)$ by $\max(\psi(q),q^{-1/d}\log^{-2/d}(q))$ may only perturb both sums in (8.27) by a bounded additive constant (which may depend on $\boldsymbol{s}$). As the right hand side of (8.27) is divergent by hypothesis, this perturbation does not affect asymptotics, so we may assume $\psi(q)\geq q^{-1/d}\log^{-2/d}(q)$ for all $q\in\operatorname{K_{small}}$. Recalling (8.8) and $\widetilde{{\mathbbm{1}}}_{R_{k}}\geq\varphi_{k}$, then combining Proposition˜8.12 and Lemma˜8.13, we obtain that $\xi$-almost surely, for large enough $n$,
$\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})\geq\sum_{k\in\operatorname{K_{small}}(n)}Y_{k}(\boldsymbol{s})\geq(1-\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}y_{k}.$
By (8.12), we have $\sum_{k\in\operatorname{K_{small}}(n)}y_{k}=O_{\gamma_{2}}(1)+\sum_{k\in\operatorname{K_{small}}(n)}\operatorname{Leb}(R_{k}^{-})$. Extracting from (8.7) that $\operatorname{Leb}(R_{k}^{-})=(1-\tau^{-1}-2\varepsilon)(1-2\varepsilon)^{d}\psi(\tau^{k})^{d}\tau^{k}$, and using that the associated series diverges by hypothesis, the result follows. ∎
Conclusion for the lower bound (case $d\geq 2$)
Proof of Proposition˜8.4.
It follows by combining Lemma˜8.7 and Corollary˜8.15. ∎
8.2. Case $d\geq 2$, the upper bound
The proof of the asymptotic upper bound in ˜8.3 (case $d\geq 2$) is similar to that of the lower bound. We briefly sketch the proof to highlight the relevant changes. We extend $\psi$ to $\mathbb{R}_{\geq 0}$ by setting $\psi(q)=\psi(\lfloor q\rfloor)$ for non-integer values of $q$. We let $\tau\in(1,2]$ and note that for all integers $N,n\geq 1$ such that $\tau^{n}\leq N<\tau^{n+1}$, for every $\boldsymbol{s}\in\mathbb{R}^{d}$, we have
$\mathscr{T}_{N}(\boldsymbol{s})\leq\sum_{k=0}^{n}\mathscr{S}_{k}^{+}(\boldsymbol{s})$
where for $k\geq 0$, we set
$\mathscr{S}^{+}_{k}(\boldsymbol{s}):=\big{|}\{(\boldsymbol{p},q)\in\mathbb{Z}^{d}\times\llbracket\tau^{k},\tau^{k+1}\llbracket\,:\,\forall i\in\llbracket 1,d\rrbracket,\,0\leq qs_{i}-p_{i}<\psi(\tau^{k})\}\big{|}.$
Then it suffices to show the upper bound analogue of Proposition˜8.4.
Lemma 8.16.
Under the assumptions of ˜8.3 and with $d\geq 2$, we have for every $\tau\in(1,2]$, for $\xi$-almost all $\boldsymbol{s}\in\mathbb{R}^{d}$, for every $\eta>0$, for all large enough $n$,
$\sum_{k=1}^{n}\mathscr{S}^{+}_{k}(\boldsymbol{s})\leq(\tau-1+\eta)\sum_{k=1}^{n}\psi(\tau^{k})^{d}\tau^{k}.$
To show Lemma˜8.16, we note that
$\mathscr{S}_{k}^{+}(\boldsymbol{s})=\widetilde{{\mathbbm{1}}}_{P_{k}}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}$
where $P_{k}:=[0,r_{k})^{d}\times[r_{k},\tau r_{k})$ and $r_{k},t_{k}>0$ have been defined in (8.7). Keep the notations $\gamma_{1}$, $\operatorname{K_{big}}$, $\operatorname{K_{small}}$, from the proof of the lower bound. We establish the upper bound announced in Lemma˜8.16 along the subsums $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})$ and $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})$ separately, and under the assumptions that the corresponding $\sum_{k\in\operatorname{K_{big}}}\psi(\tau^{k})^{d}\tau^{k}$, $\sum_{k\in\operatorname{K_{small}}}\psi(\tau^{k})^{d}\tau^{k}$ diverge. This is enough in view of the convergent case of the Khintchine dichotomy, ˜8.2.
The proof of the asymptotic upper bound for $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})$ is the same as that of Lemma˜8.7, but using this time the upper bound from Lemma˜8.6 instead of the lower bound.
To deal with $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})$, we recall the parameters $\gamma_{2},\varepsilon$, $\chi_{k}$, $\theta$ from the proof of the lower bound. We introduce the thickened box
$P_{k}^{+}:=[-\varepsilon r_{k},(1+\varepsilon)r_{k})^{d}\times[(1-\varepsilon)r_{k},(\tau+\varepsilon)r_{k})$
and note that $\widetilde{{\mathbbm{1}}}_{P_{k}}\leq\theta*\widetilde{{\mathbbm{1}}}_{P_{k}^{+}}$ (because $\theta$ is supported on $B_{\varepsilon/10}$). We consider the smooth truncated companion
$\varphi_{k}^{+}=\theta*(\chi_{k}\widetilde{{\mathbbm{1}}}_{P_{k}^{+}}),$
where $\chi_{k}$ is defined as in (8.9). In view of (4.3), $\varphi_{k}^{+}$ coincides with $\theta*\widetilde{{\mathbbm{1}}}_{P_{k}^{+}}$ on the subset $\{\operatorname{dist}(\cdot,x_{0})\leq\frac{\gamma_{2}}{M}\log t_{k}-M\}$ for some $M=M(d)>1$.
By Lemma˜8.5, for $\xi$-almost every $\boldsymbol{s}$, for large enough $k$, we have
$\operatorname{dist}(a(t_{k})u(\boldsymbol{s})x_{0},x_{0})\leq\log\log t_{k},$
and in particular,
$\mathscr{S}^{+}_{k}(\boldsymbol{s})\leq(\theta*\widetilde{{\mathbbm{1}}}_{P_{k}^{+}})(a(t_{k})u(\boldsymbol{s})x_{0})=\varphi^{+}_{k}(a(t_{k})u(\boldsymbol{s})x_{0}).$
Therefore, we only need to show the upper bound analogue of Corollary˜8.15:
(8.28) $\sum_{k\in\operatorname{K_{small}}(n)}\varphi^{+}_{k}(a(t_{k})u(\boldsymbol{s})x_{0})\leq(\tau-1+C(d)\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})^{d}\tau^{k}.$
where $C(d)>0$ is a constant depending on $d$ only. The estimate (8.28) follows mutatis mutandis from the argument establishing the lower bound for partial sums over $\operatorname{K_{small}}$, in which we replace $R_{k}^{-}$ by $P_{k}^{+}$.
We have thus established Lemma˜8.16, whence the asymptotic upper bound
$\limsup_{N\to+\infty}\frac{\mathscr{T}_{N}(\boldsymbol{s})}{\sum_{q=1}^{N}\psi(q)^{d}}\leq 1$
of ˜8.3 (case $d\geq 2$).
8.3. The case $d=1$.
It remains to establish ˜8.3 in the case where $d=1$. The proof is similar to the higher dimensional case but a certain number of refinements are required due to poorer moment estimates for Siegel transforms.
Let us start with the lower bound. Keep the notations $\tau$, $R_{k}$, $\gamma_{1}$, $\operatorname{K_{big}}$, $\operatorname{K_{small}}$, $\gamma_{2}$, $\chi_{k}$, $\varepsilon$, $R_{k}^{-}$, $\theta$, from §8.1. The asymptotic lower bound for $\sum_{k\in\operatorname{K_{big}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$ given in Lemma˜8.7 is still valid because the argument works without restriction on $d$.
We thus focus on the lower asymptotic for $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$, more precisely on extending Corollary˜8.15 to the case $d=1$. The difference with the higher dimensional case is that the Siegel transform of a ball in $\mathbb{R}^{2}$ does not have finite second moment, in particular (8.11) is not valid anymore. To deal with this obstacle, we restrict the Siegel transform by counting only lattice points $(p,q)$ which are bounded multiple of a primitive point. Namely, given $m>0$, we set
${\mathcal{P}}^{(m)}(\mathbb{Z}^{2}):=\{(p,q)\in\mathbb{Z}^{2}\smallsetminus\{0\}\,:\,\operatorname{GCD}(p,q)\leq m\}$
where $\operatorname{GCD}(p,q)\in\mathbb{N}_{\geq 1}$ denotes the greatest common divisor of $p$ and $q$. Given a measurable function $f:\mathbb{R}^{2}\rightarrow\mathbb{R}_{\geq 0}$, we define its restricted Siegel transform $\widetilde{f}^{(m)}:X\rightarrow[0,+\infty]$ by: $\forall g\in G$,
$\widetilde{f}^{(m)}(gx_{0}):=\sum_{v\in{\mathcal{P}}^{(m)}(\mathbb{Z}^{2})}f(gv).$
In this context, we have the following moment estimates. Their vocation is to replace ˜8.8 from the higher dimensional case.
Proposition 8.17 (Moments of restricted Siegel transforms).
Let $m\in\mathbb{N}_{\geq 1}$, let $k\in\operatorname{K_{small}}$. Let $c_{m}:=\zeta(2)^{-1}\sum_{t=1}^{m}t^{-2}$ and let $R\subseteq\mathbb{R}^{2}$ be a rectangle such as $R_{k}^{-}$ here and $P_{k}^{+}$ further below. We have
(8.29) $\int_{X}\widetilde{{\mathbbm{1}}}^{(m)}_{R}\,\mathrm{d}m_{X}=c_{m}\operatorname{Leb}(R)$
(8.30) $\int_{X}(\widetilde{{\mathbbm{1}}}^{(m)}_{R})^{2}\,\mathrm{d}m_{X}=c_{m}^{2}\operatorname{Leb}(R)^{2}+O\left(\operatorname{Leb}(R)\log(1+m)\right).$
Proof.
Those estimates appear in the litterature for primitive Siegel transforms, i.e. for $m=1$. We explain how to deduce the case $m\geq 1$. Note that
$\widetilde{{\mathbbm{1}}}^{(m)}_{R}=\sum_{q\in\llbracket 1,m\rrbracket}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q}$.
The Siegel summation formula for primitive Siegel transform [50, Equation ($8$)] guarantees $\int_{X}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q}\,\mathrm{d}m_{X}=\zeta(2)^{-1}\operatorname{Leb}(R/q)$, whence (8.29).
To justify (8.30), we note that
$\int_{X}(\widetilde{{\mathbbm{1}}}^{(m)}_{R})^{2}\,\mathrm{d}m_{X}=\sum_{q,q^{\prime}\in\llbracket 1,m\rrbracket}\int_{X}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q^{\prime}}\,\mathrm{d}m_{X}.$
By Rogers [48, Theorem 5], we have
$\int_{X}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q}\widetilde{{\mathbbm{1}}}^{(1)}_{R/q^{\prime}}\,\mathrm{d}m_{X}$
$=\zeta(2)^{-2}q^{-2}q^{\prime-2}\operatorname{Leb}(R)^{2}+\zeta(2)^{-1}\operatorname{Leb}(R/q\cap R/q^{\prime})+\zeta(2)^{-1}\operatorname{Leb}(R/q\cap(-R/q^{\prime})).$
Summing over $q,q^{\prime}\in\llbracket 1,m\rrbracket$ the first term of the right hand side gives
$\sum_{q,q^{\prime}\in\llbracket 1,m\rrbracket}\zeta(2)^{-2}q^{-2}q^{\prime-2}\operatorname{Leb}(R)^{2}=c_{m}^{2}\operatorname{Leb}(R)^{2}.$
On the other hand, looking at the second coordinate of points in $R$, we find for any $q\geq 1$,
$|\{q^{\prime}\geq 1\,:\,R/q\cap R/q^{\prime}\neq\varnothing\}|\ll q$
Using the trivial bound $\operatorname{Leb}(R/q\cap R/q^{\prime})\leq\operatorname{Leb}(R)/q^{2}$, we deduce
$\sum_{q,q^{\prime}\in\llbracket 1,m\rrbracket}\operatorname{Leb}(R/q\cap R/q^{\prime})\ll\sum_{q\in\llbracket 1,m\rrbracket}\operatorname{Leb}(R)/q\simeq\operatorname{Leb}(R)\log(1+m).$
As $R/q\cap(-R/q^{\prime})=\varnothing$, the last term does not contribute, and this concludes the proof of (8.30). ∎
From there, the proof of the higher dimensional case goes through with a few adaptations. Let $(m_{k})_{k\in\operatorname{K_{small}}}\in\mathbb{R}_{>0}^{\operatorname{K_{small}}}$ be the sequence satisfying
(8.31) $\forall k\in\operatorname{K_{small}},\quad\log(1+m_{k})=\Bigl{(}\sum_{j\in\operatorname{K_{small}}(k)}\psi(\tau^{j})\tau^{j}\Bigr{)}^{1/8}.$
Set $\varphi_{k}:=\theta*(\chi_{k}\widetilde{{\mathbbm{1}}}^{(m_{k})}_{R_{k}^{-}})$ where $\chi_{k}$ is as in (8.9), and define positive constants
$y_{k}:=m_{X}(\varphi_{k}),\qquad y^{\prime}_{k}:=m_{X}(\varphi_{k})\log(1+m_{k}).$
Noting that by hypothesis, we have $\lim_{k}m_{k}=+\infty$ as $k$ tends to infinity along $\operatorname{K_{small}}$, we find that $y_{k}\leq y^{\prime}_{k}$ for all large $k\in\operatorname{K_{small}}$.
Replacing ˜8.8 by Proposition˜8.17 in the proof of Lemma˜8.10, we obtain the following.
Lemma 8.18.
If $\gamma_{1}\lll\gamma_{2}$, then for some $M=M(d)>1$, every $k\in\operatorname{K_{small}}$, we have
(8.32) $y_{k}=c_{m_{k}}\operatorname{Leb}(R_{k}^{-})-O(t_{k}^{-\gamma_{2}/M}),$
${\mathcal{S}}_{\infty,l}(\varphi_{k})\ll t_{k}^{M\gamma_{2}},$
${\mathcal{S}}_{2,l}(\varphi_{k})\ll y_{k}+\sqrt{y^{\prime}_{k}}+O(t_{k}^{-\gamma_{2}/M}).$
Next, writing for $\boldsymbol{s}\in\mathbb{R}$,
$Y_{k}(\boldsymbol{s})=\varphi_{k}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)},\qquad Z_{k}(\boldsymbol{s})=Y_{k}(\boldsymbol{s})-y_{k},$
Lemma˜8.11 becomes (with a similar proof) the following.
Lemma 8.19.
Assume $\gamma_{1}\lll\gamma_{2}\lll 1$. Then for every $k\in\operatorname{K_{small}}$ such that $y^{\prime}_{k}\geq 1$, we have
$\mathbb{E}[Z_{k}^{2}]\ll y^{\prime}_{k}.$
We deduce the following replacement for Proposition˜8.12.
Proposition 8.20.
Assume $\gamma_{1}\lll\gamma_{2}\lll 1$ and
$\psi(q)\geq q^{-1}\log^{-2}(q),\quad\forall q\in\operatorname{K_{small}}.$
Then for every finite subset $J\subseteq\operatorname{K_{small}}$ with $\inf J\ggg_{\gamma_{1},\gamma_{2}}1$, we have
$\mathbb{E}\bigl{[}Z_{J}^{2}\bigr{]}\ll\left(1+y^{\prime}_{J}\right)^{3/2}.$
We recall that $Z_{J}=\sum_{j\in J}Z_{j}$ and $y^{\prime}_{J}=\sum_{j\in J}y^{\prime}_{j}$. Similar notations $Y_{J},y_{J}$ will be used below.
Proof.
Same as for Proposition˜8.12 but using $y_{j}^{\prime}$ to define the partition $J=J_{1}\sqcup J_{2}\sqcup J_{3}$, and noting that all the upper bounds in the proof of Proposition˜8.12 are valid with $y_{j}^{\prime}$ at the place of $y_{j}$ thanks to Lemmas 8.18, 8.19, and the inequality $y_{j}^{\prime}\geq y_{j}$ (which is valid thanks to the assumption $\inf J\ggg_{\gamma_{1}}1$). ∎
We can now combine Proposition˜8.20 and Lemma˜8.13 (note here that we allow $(y^{\prime}_{j})$ to be different from $(y_{j})$ in the latter) to obtain the following.
Lemma 8.21.
For $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $n$, we have
(8.33) $\lvert Z_{\operatorname{K_{small}}(n)}(\boldsymbol{s})\rvert\leq\left(y^{\prime}_{\operatorname{K_{small}}(n)}\right)^{4/5}.$
We now claim that the right-hand side in (8.33) is negligible compared to $y_{\operatorname{K_{small}}(n)}$. To see why, note first that by definition, the sequence $(m_{k})_{k\in\operatorname{K_{small}}}$ is non-decreasing, therefore
(8.34) $y^{\prime}_{\operatorname{K_{small}}(n)}\leq y_{\operatorname{K_{small}}(n)}\log(1+m_{\max\operatorname{K_{small}}(n)}).$
Moreover, by (8.7), (8.32), we have $\psi(\tau^{k})\tau^{k}\ll y_{k}+t^{-\gamma_{2}/M}_{k}$, so using the definition of $m_{k}$, we get
(8.35) $\forall k\in\operatorname{K_{small}},\quad\log(1+m_{k})\ll\left(y_{\operatorname{K_{small}}(k)}+O_{\gamma_{2}}(1)\right)^{1/8}.$
Equations (8.34) and (8.35) together justify the claim.
Lemma˜8.21 and the above claim yield in particular that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for sufficiently large $n$,
$Y_{\operatorname{K_{small}}(n)}(\boldsymbol{s})\geq(1-\varepsilon)y_{\operatorname{K_{small}}(n)}.$
By construction, we know that $\mathscr{S}_{k}(\boldsymbol{s})\geq Y_{k}(\boldsymbol{s})$. On the other hand, it follows from (8.32) that $y_{\operatorname{K_{small}}(n)}\geq\sum_{k\in\operatorname{K_{small}}(n)}c_{m_{k}}\operatorname{Leb}(R_{k}^{-})-O_{\gamma_{2}}(1)$, where $c_{m_{k}}\to_{k}1$ as $k$ goes to infinity along $\operatorname{K_{small}}$. Using (8.7) to see that $\operatorname{Leb}(R_{k}^{-})=(1-\tau^{-1}-2\varepsilon)(1-2\varepsilon)\psi(\tau^{k})\tau^{k}$, we infer that for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for all large enough $n$,
$\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})\geq(1-\tau^{-1}-6\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})\tau^{k}.$
This concludes the proof of the lower bound.
Let us now justify the upper bound in the case $d=1$. Similarly to the higher dimensional case, we can mimic the proof of the lower bound estimate in the case $d=1$ to establish an upper bound estimate. However, this upper bound concerns the restricted Siegel transforms which are used in the proof, while we aim for an upper bound without restriction on $\operatorname{GCD}(p,q)$ when counting solutions $(p,q)$ of the Khintchine inequality. We explain below how to deal with this obstacle.
We keep the notations of §8.2, in particular we fix $\tau\in(1,2]$, and consider for $k\geq 0$, $\boldsymbol{s}\in\mathbb{R}$,
$\mathscr{S}_{k}^{+}(\boldsymbol{s})$ $=\big{|}\{(p,q)\in\mathbb{Z}\times\llbracket\tau^{k},\tau^{k+1}\llbracket\,:\,0\leq q\boldsymbol{s}-p<\psi(\tau^{k})\}\big{|}$
$=\widetilde{{\mathbbm{1}}}_{P_{k}}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}.$
where $P_{k}=[0,r_{k})\times[r_{k},\tau r_{k})$. The goal is to show Lemma˜8.16 when $d=1$. Provided $\sum_{k\in\operatorname{K_{big}}}\psi(\tau^{k})\tau^{k}=+\infty$, the argument for Lemma˜8.7 yields the desired upper bound for the partial sums $\sum_{\operatorname{K_{big}}(n)}\mathscr{S}_{k}^{+}(\boldsymbol{s})$. Therefore we only need to deal with $\sum_{\operatorname{K_{small}}(n)}\mathscr{S}_{k}^{+}(\boldsymbol{s})$, and under the assumption $\sum_{k\in\operatorname{K_{small}}}\psi(\tau^{k})\tau^{k}=+\infty$ (as noted for Lemma˜8.16).
Recall from §8.2 that $\varepsilon>0$ is an arbitrarily small number and $P_{k}^{+}$ denotes the $\varepsilon r_{k}$-thickening of $P_{k}$. Define for $k\in\operatorname{K_{small}}$ and $m>0$,
$\varphi^{+(m)}_{k}=\theta*(\chi_{k}\widetilde{{\mathbbm{1}}}^{(m)}_{P_{k}^{+}})\quad\text{and}\quad Y^{+(m)}_{k}(\boldsymbol{s})=\varphi^{+(m)}_{k}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}.$
Let $(m_{k})_{k\in\operatorname{K_{small}}}$ be as in (8.31). We use the shorthand $\varphi^{+}_{k}:=\varphi^{+(m_{k})}_{k}$, $Y^{+}_{k}:=Y^{+(m_{k})}_{k}$. The argument used for the lower bound (case $d=1$) then shows that, provided $\gamma_{1}\lll 1$, we have for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $n$,
(8.36) $\sum_{k\in\operatorname{K_{small}}(n)}Y^{+}_{k}(\boldsymbol{s})\leq(\tau-1+10\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})\tau^{k}.$
We now compare the left hand side of (8.36) with $\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}_{k}(\boldsymbol{s})$. In other terms, we need to show that the truncation of the cusp induced by $\chi_{k}$, and most importantly the reduction of the Siegel transform to counting $m_{k}$-primitive lattice points does not affect too much the asymptotic of the partial sums. The next lemma is a first step to replace $m_{k}$ by a term $m^{\prime}_{k}$ which grows exponentially in $k$.
Lemma 8.22.
Let $m^{\prime}_{k}:=\max\{m_{k},t_{k}^{\gamma_{2}}\}$. Provided that $\gamma_{1},\gamma_{2}\lll 1$, we have for every $k\in\operatorname{K_{small}}$,
$\mathbb{E}\bigl{[}Y^{+(m^{\prime}_{k})}_{k}-Y^{+}_{k}\bigr{]}\ll r_{k}^{2}m^{-1}_{k}+t_{k}^{-\gamma_{2}}.$
Proof.
We may assume $m_{k}\leq t_{k}^{\gamma_{2}}$. Unfolding definitions, then using effective single equidistribution (8.1) while noting that ${\mathcal{S}}_{\infty,l}(\varphi_{k}^{+(m^{\prime}_{k})}-\varphi_{k}^{+})\ll t_{k}^{M\gamma_{2}}$ for some $M$ as in Lemma˜8.18, we have
$\mathbb{E}\bigl{[}Y^{+(m^{\prime}_{k})}_{k}-Y^{+}_{k}\bigr{]}$ $=\int_{\mathbb{R}^{d}}\theta*[\chi_{k}(\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}^{+}}-\widetilde{{\mathbbm{1}}}^{(m_{k})}_{P_{k}^{+}})]\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}\,\mathrm{d}\xi(\boldsymbol{s})$
$\ll m_{X}(\widetilde{{\mathbbm{1}}}_{P_{k}^{+}}-\widetilde{{\mathbbm{1}}}^{(m_{k})}_{P_{k}^{+}})\,+\,t_{k}^{-c+M\gamma_{2}}$
$\ll r_{k}^{2}m^{-1}_{k}+t_{k}^{-\gamma_{2}},$
where the last inequality uses (8.29), the definition of $c_{m_{k}}$ and assumes $\gamma_{2},\gamma_{1}$ small enough depending on $c$. ∎
We deduce that (8.36) is still valid for $Y^{+(m^{\prime}_{k})}_{k}$ in the place of $Y^{+}_{k}$.
Lemma 8.23.
For $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $n$, we have
$\sum_{k\in\operatorname{K_{small}}(n)}Y^{+(m^{\prime}_{k})}_{k}(\boldsymbol{s})\leq(\tau-1+11\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})\tau^{k}.$
Proof.
Let $j\geq 1$, set $I_{j}:=\{n\in\operatorname{K_{small}}\,:\,\sum_{\operatorname{K_{small}}(n)}r_{k}^{2}\in(j,j+1]\}$. We have by Lemma˜8.22
$\mathbb{E}[\sum_{k\in I_{j}}(Y^{+(m^{\prime}_{k})}_{k}-Y^{+}_{k})]\ll\sum_{k\in I_{j}}(r_{k}^{2}m^{-1}_{k}+t_{k}^{-\gamma_{2}})\leq\frac{j+1}{e^{j^{1/8}}-1}+\sum_{k\in I_{j}}t_{k}^{-\gamma_{2}},$
[…]
To conclude, we show the reduction prescribed by $m^{\prime}_{k}$ is loose enough not to affect the counting.
Lemma 8.24.
Assume $\gamma_{1}\lll\gamma_{2}$. For $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $k$, we have
$\mathscr{S}^{+}_{k}(\boldsymbol{s})\leq Y^{+(m^{\prime}_{k})}_{k}(\boldsymbol{s}).$
Proof.
Note that we have $\theta*(\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}^{+}})\geq\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}}$, because $gP^{+}_{k}\supseteq P_{k}$ for all $g\in\operatorname{supp}\theta$. Therefore, recalling the truncation function $\chi_{k}={\mathbbm{1}}_{\{\operatorname{inj}\geq t_{k}^{-\gamma_{2}}\}}$, and writing $\chi^{\prime}_{k}$ the characteristic function of the set of points $x$ such that $B_{1}x\subseteq\operatorname{supp}\chi_{k}$, we have $\theta*(\chi_{k}\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}^{+}})\geq\chi^{\prime}_{k}\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}}$. It follows that
(8.37) $Y^{+(m^{\prime}_{k})}_{k}(\boldsymbol{s})\geq\bigl{(}\chi^{\prime}_{k}\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}}\bigr{)}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}.$
By Lemma˜8.5, for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $k\in\operatorname{K_{small}}$, we have $\operatorname{dist}(a(t_{k})u(\boldsymbol{s})x_{0},x_{0})\ll\log\log t_{k}$, meaning that
(8.38) $a(t_{k})u(\boldsymbol{s})\in g_{k}\operatorname{SL}_{2}(\mathbb{Z})\text{ for some $g_{k}\in G$ satisfying }\lVert g_{k}\rVert\leq(\log t_{k})^{O(1)}.$
In particular, $a(t_{k})u(\boldsymbol{s})x_{0}\in\operatorname{supp}\chi^{\prime}_{k}$ for $k$ sufficiently large, whence the truncation $\chi^{\prime}_{k}$ plays no role. Moreover, as $k\in\operatorname{K_{small}}$, we have that $P_{k}$ is included in the $t_{k}^{O(\gamma_{1})}$-neighborhood of the origin. Combined with (8.38), this yields that every point $(p,q)\in g_{k}^{-1}P_{k}\cap\mathbb{Z}^{2}$ satisfies $\operatorname{GCD}(p,q)\leq t_{k}^{\gamma_{2}}\leq m^{\prime}_{k}$ (provided $\gamma_{1}\lll\gamma_{2}$). It follows that
(8.39) $\bigl{(}\chi^{\prime}_{k}\widetilde{{\mathbbm{1}}}^{(m^{\prime}_{k})}_{P_{k}}\bigr{)}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}=\widetilde{{\mathbbm{1}}}_{P_{k}}\bigl{(}a(t_{k})u(\boldsymbol{s})x_{0}\bigr{)}=\mathscr{S}^{+}_{k}(\boldsymbol{s}).$
Equations (8.37) and (8.39) together finish the proof. ∎
The combination of Lemmas 8.23, 8.24 yields for $\xi$-almost every $\boldsymbol{s}\in\mathbb{R}$, for large enough $n$,
$\sum_{k\in\operatorname{K_{small}}(n)}\mathscr{S}^{+}_{k}(\boldsymbol{s})\leq(\tau-1+11\varepsilon)\sum_{k\in\operatorname{K_{small}}(n)}\psi(\tau^{k})\tau^{k}.$
This concludes the proof of the asymptotic upper bound in ˜8.3 in the case $d=1$.